A Complete Bibliography of Publications in Significance (Oxford, UK)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 1.08

Title word cross-reference

$34.95 [1236]. $49.95 [1236]. $5m [921]. $99.95 [1236]. _{2} [1680]. g [1247]. h [1247]. p [309, 355, 821, 1619, 1007, 1471]. t [280, 281].

-and- [1247]. -test [281]. -tests [280]. -value [355, 1619, 1007, 1471]. -values [821].

100th [1714]. 19 [1592, 1364, 1367, 1425, 1451, 1380, 1458, 1395, 1383, 1466, 1382, 1363, 1359].

2012 [564]. 2020 [1330, 1086]. 2023 [1750]. 21st [129]. 21st-century [129]. 224pp [1236]. 256pp [1236]. 288pp [1479].

423pp [1236].

7 [1479].

9 [1236]. 978 [1479, 1236]. 978-0-19-872906-8 [1236]. 978-0-19-878784-6 [1236]. 978-1-4987-8905-9 [1236]. 978-1-5013-5400-7 [1479].

A&E [1033]. Abbott [1227]. academia [1419]. academic [1561]. accent [1005]. access [228, 1015]. Accidental [8]. according [1447]. accounting [234]. achievement [468]. act [1587, 822, 264]. action [1179]. activity [1233]. actors [983]. Acts [32]. Actuarial [115]. actuary [1289]. addiction [1286]. address [76]. Addressing [1553]. ADHD [965]. administration [80]. administrative [1293]. advantage [1507]. adventurous [969]. affect [1129]. Afghanistan [381]. afraid [745, 700]. Africa [1097]. After [1395, 515, 643, 1382]. AGA [468]. again [1040]. against [1200, 769, 1052]. Age [421, 1041, 563, 998, 1352]. Agency [414]. ages [1306, 1632]. aggression [941]. aging [390]. Agricultural [9]. agriculture [640]. AI [1648, 1126]. aid [715, 667]. aim [478]. aims [1184]. Air [1170]. aircraft [275]. airport [1043]. alarm [1454]. album [192, 1734]. Alcohol [291, 653, 1478]. alert [1414]. Alexander [363]. Alexandria [1117]. algorithm [1519, 1518, 998, 1407]. algorithmic [1303]. Algorithms [1162, 1071, 1539, 1244]. Ali [1640]. Alison [1599]. alive [1651]. Alley [1473]. Alley-oop [1473]. Alliance [1527]. allies [1254]. allocation [1055]. alone [379, 1094, 1649]. Alphonse [731]. Already [310, 549]. alt [1686]. always [1716, 1476]. Amanda [599]. amendments [1181]. America [948]. American [1292, 884, 858]. Americans [722]. amid [1003, 485]. analogues [1003]. Analysing [588, 653]. Analysis [1327, 414, 739, 556, 110, 1362, 1194, 889, 873, 714]. Analytica [1125]. analytics [1132, 931, 1670, 1473, 1513]. Anar [1721, 1166]. anarchy [656]. ANASE [275]. Andreas [1144, 1677]. Andrew [1236]. Andy [1236]. angel [1319]. Angrist [1547]. animal [220, 375]. animals [454, 1733]. Anna [1623]. annoyance [275]. Annoying [261]. anomaly [1590]. answer [537, 1599]. answered [1623, 1657, 1640, 1673, 1692, 1583, 1727, 1517, 1566, 1710]. Answering [1547]. anti [390]. anti-aging [390]. antidepressants [287, 594]. anxiety [1578]. any [304]. anyone [320]. Apollo [1518]. Appealing [89]. applications [1232]. applied [1476]. appointments [1215]. Apprentice [804]. Arab [538]. Archaeology [638]. archers [1703]. archives [13]. arcsinh [1225]. area [1751, 93]. area-wide [93]. aren’t [1094]. Argentina [1390]. argument [1174, 1544]. arithmetic [1182]. armed [622]. arms [446]. arrive [513]. arrives [1513]. Ars [660]. art [1574, 317, 1670, 239, 1171, 997]. Artificial [1204, 1025, 1265]. artistic [732]. artwork [1384]. ASA [1007]. Ask [1449, 1129, 1110, 1248, 674, 1612, 1676]. asks [537]. Assessing [518, 610, 903]. Association [1292]. assumption [417]. astronomy [582]. Asylum [98]. ate [1016]. athlete [348]. Athletics [320]. attend [571]. attends [199]. attitudes [740]. attuned [1636]. audio [1376]. August [655]. Authority [357, 1106]. authors [1399]. autopsy [888]. available [1586].

Avalanche [356]. aviation [431]. avoid [465, 836]. Award [1378]. awarded [1486].

Babel [306]. Babies [211, 653, 886]. Baboon [1645]. baby [1023, 644, 771]. back [1480, 562, 1105]. backtest [1554]. Bacon [278]. bad [1718, 1328, 478]. badgers [259]. bag [1571]. Ball [402, 462]. Ballet [1715]. ballot [276]. balls [1571]. ban [821]. bang [411]. bank [1191, 1020]. banknotes [81]. banks [288, 565]. bar [1661, 221]. Barabba [1172]. Barefoot [456]. Barents [930]. bargains [1328]. Barrow [413]. base [904]. Baseball [1654]. based [181, 984, 1303]. basketball [1221, 1473]. BASP [821]. Bastardy [51]. Bats [1508, 214]. Battle [1381, 1680]. Bay [221]. Bayes [1426, 1416, 632, 204, 419]. Bayesian [927, 1112, 185, 27, 1647, 831, 873]. BBC [262]. be [1346, 536, 1586, 1023, 500, 1153, 1474, 201, 725, 1204, 1548, 1185, 1753, 1520, 202, 418, 804, 1019, 1034, 1061, 1107, 1599, 435, 789, 759, 1414]. Beard [657]. Beating [703]. beauties [1356]. Beautiful [909, 1665, 584, 390, 1132, 474, 507]. beauty [1636, 568, 1701]. Beckerman [1236]. become [707, 743]. Becoming [249]. been [1361, 1476, 1529]. Beer [184]. bees [652]. before [340, 1213, 202, 985]. befuddlement [295]. behaviour [1324]. behavioural [400, 668, 1194]. behaviours [1628]. Behind [975, 1617, 1343]. Beijing [320]. being [1341, 1612, 151]. beings [337]. believe [321, 134]. bells [1313]. Belt [236, 296]. benefits [287, 726]. Benford [912, 234]. Bergen [1111]. Berkson [1380]. Bernoulli [1151]. Bertillon [731]. Best [765, 785, 798, 812, 827, 840, 850, 865, 348, 670, 983, 1540, 1110, 1719, 448, 418, 1580]. bet [158, 1530]. better [1480, 743, 549, 1382, 1736, 669]. Betting [143, 1483, 148, 41]. between [1751, 272, 1747, 42, 726]. Bevacqua [615]. Beware [600, 1496]. Beyond [464, 1518, 984, 729]. bias [1290, 1380, 1647, 1699]. Bicycle [494, 61]. BiDil(R) [181]. Big [582, 341, 1617, 770, 578, 591, 671, 1077, 1456, 1021, 1477, 1416, 1519, 745, 513, 1392, 1539, 1488, 1676, 952, 1037, 1438, 818]. bike [667]. Bill [617, 248]. bind [1266]. bingo [1549]. binomial [1701]. biodiversity [366, 178]. biofuels [498]. Bioinformatics [2]. Biologists [1236]. biology [581]. Biomedical [200, 454]. biosecurity [413]. biostatistician [342]. Biostatistics [530]. Bioterrorism [62]. Birds [178, 717, 981]. birdsong [607]. birth [374, 905]. birthday [246, 697, 963, 978]. bit [535]. bits [1735]. bitter [183]. black [181, 745, 888]. blackouts [1682]. Blame [1407, 711]. blanks [95]. Bletchley [419]. blind [227, 488]. blind-spot [488]. blood [1095, 1509]. Bloomsbury [1479]. blow [1208]. Blowin [654]. Blowing [212]. bluff [150]. Blurring [1577]. board [771]. BoardGameGeek [1291]. boards [487]. boldly [530]. Bolt [547]. bomb [1289]. bombs [381]. bone [563]. Bonnie [1692]. Book [763, 972, 1620, 1707, 1724, 1740, 1755, 1600]. Books [1541, 1396, 1521]. boosts [487]. bootstrap [458, 1205, 67]. Border [414]. borders [356, 643]. Borja [1504]. born [644]. borns [571]. Bot [1650]. both [1536]. bottles [211]. Bovine [259]. Bowen [1688]. bowl [308, 1061]. box [745, 658, 437].

Boxplots [118]. Brad [524]. Bradley [455]. brain [986, 653, 835]. brains [927, 1096]. break [1191, 1299]. breaker [1653]. breakfast [735, 340]. Breaking [1142, 1548, 679]. Breast [448, 929, 610, 470]. breastfeeding [448]. breasts [211]. Breiman [1331]. Brexit [923, 1041, 883, 1133]. bridge [421]. bridges [1508]. brighter [1287]. bring [319]. Bringing [968]. Britain [354, 228, 1133]. British [1040, 426, 1164, 256, 225]. British-led [225]. Britten [1623]. Brother [341]. brow [729]. BS [743]. bubble [1021]. Buchhorn [1517]. budget [1108]. Buffett [843]. build [1026]. Building [1480, 1752]. built [1739]. bull [730]. Bullshit [1440, 1295]. bun [1532]. buried [1057]. burning [903]. Bursting [1021]. Bus [667]. Business [40, 1512, 793, 1478, 988, 671]. businesses [515]. Busts [243]. Buttrey [1177, 1220, 1259, 1299, 1583]. buy [433, 465]. buying [323, 1387]. buzz [652].

C [952]. C. [520, 1391]. Calculate [1196]. calculator [620]. California [921]. Call [150, 730, 1753, 1015]. Calling [1703, 1440]. calls [1303, 859]. Cambridge [1125]. cameras [179, 395]. Campanometry [1313]. Campbell [1185]. Can [1749, 483, 1587, 1666, 1736, 1508, 670, 901, 1439, 547, 1754, 1497, 415, 241, 465, 290, 744, 320, 500, 578, 1682, 926, 1575, 447, 1125, 1207, 435, 673, 759, 194, 546]. Canada [450]. Cancer [808, 1329, 1053, 1239, 272, 929, 470]. Cane [1653]. Canine [1633]. cannabis [683]. cannot [1575]. capacity [1427]. Cape [1113]. Carbon [323, 177, 712]. card [152, 1664, 834]. care [434, 644, 84]. Career [1500, 65, 132, 83, 1097, 10, 1457, 1437, 28, 218, 47, 151, 115, 1754, 1645, 341, 1752]. carefully [1705]. carelessness [1053]. Cargo [1154]. Cargo-cult [1154]. Carnegie [1360]. carriage [687]. cars [924]. cartels [874]. case [1607, 180, 1627, 422, 930, 78, 1466, 1069, 1166]. casebook [45, 63, 96, 114, 166, 199, 216, 230, 245, 261, 277, 293, 309, 324, 339, 371, 386, 401, 417, 433, 449, 469, 486, 499, 517, 536, 551, 566, 584, 598, 612, 629, 642, 657, 670, 684, 701, 714, 728, 742, 762, 781, 795, 837, 847, 862, 914, 934, 953, 971, 990, 1009, 1027, 1045, 1063, 1080, 1099, 1119, 1138, 1155, 1174, 1195, 1216, 1235, 1256, 1278, 1296, 1316, 1335, 1354, 1374, 7, 22, 182, 355]. cases [1425, 77]. cash [1074]. casino [703]. casualty [872]. catch [447, 1535]. catching [464, 413, 214]. cattle [259]. Caucus [1353]. causal [1547]. causation [905]. Cause [1240, 1501, 1369]. causes [336]. Caution [771]. cautionary [223, 890]. CCTV [395]. Celebrate [628]. célèbre [1501]. cells [1406]. cellular [581]. census [450, 452, 451, 1292, 1330, 445, 1069, 1086]. CensusAtSchool [453]. centenary [502]. centre [580]. century [1150, 570, 1391, 129, 1548, 280]. certain [398, 1530, 534]. chains [422]. Challenge [688]. Challenging [1664]. Chalmers [743]. champagne [718]. champion [1252]. chance [629, 379, 999]. chances [895, 925]. Change [1759, 1357, 1443, 608, 1628, 59, 226, 398, 321, 1395, 712, 399, 944, 951]. changed [1190]. Changes [1411]. changing [1705, 1495, 844]. chaos [581]. Chapman [1236]. Charles [374]. charting [279]. Charts [1311, 30, 221, 1734]. Chasing [1159]. ChatGPT [1694]. cheats [464, 1675]. check [1675, 805]. checking [1181]. Chen [1561, 1560, 1562, 1559]. Chess [1675]. chi [1537]. chi-square [1537]. Chicago [1227, 1004]. child [655, 416]. children [258, 644, 1666, 700]. Childs [1236]. chimps [700]. China [336, 1248]. chips [147]. chocolate [1661, 693]. choice [304, 353]. Cholera [170, 488]. Chomsky [585]. Chopin [514]. Christmas [776, 1310]. cite [202]. cited [202]. cities [1379, 481, 485]. Citizen [869]. citizens [1008]. citizenship [1086]. City [484, 578, 1453, 482, 486, 756, 1039]. City-wide [484]. Civic [1182]. civil [1627, 1432, 890]. civilisation [789]. claims [1459]. Claire [1688]. Clark [78]. class [683, 426, 278]. classed [1019]. classroom [1116, 1549]. claw [367]. clay [426]. Cleaning [1128]. clear [474, 1475, 1474, 507]. clearly [1439, 1418]. cliché [834]. Climate [398, 944, 608, 59, 226, 532, 321, 689, 712, 399, 985]. Climatological [218]. climb [1749]. clinical [339, 287, 516, 790]. Clinton [1035]. clock [1153]. clocks [539]. close [350, 356]. closer [1538]. Closing [879, 891, 1282, 1300]. Closing/On [879]. closures [903]. clothes [243]. Cloudy [999]. club [68]. Co [1399, 1680]. Co-authors [1399]. Coach [931]. coals [968]. Coast [667]. Cocaine [1074]. code [1586]. codes [1653]. Codex [642]. Coincidence [1436, 540]. coincidences [539]. Cold [930]. collaboration [1348]. collaborations [1592]. collection [1676]. College [1614]. Collider [383, 1380]. colonies [1164]. Colorless [585]. colour [1267]. colours [728]. come [467, 739, 713]. comes [703, 1352]. Coming [747, 356]. comma [722]. Commerce [1069]. commission [1303]. commits [921]. committee [112, 420]. commodity [1054]. common [930, 1189]. Commons [168]. communicate [1196, 1136]. communicating [278]. communication [1173]. communications [1592]. communicator [1736]. communicators [1594]. community [1353]. commuter [1324]. commuting [494]. company [56, 487]. Comparing [1478]. comparisons [1231]. compass [411]. competition [980, 1689, 951, 402, 462, 720, 628, 679]. Complacency [1271]. completeness [369]. completing [192]. complex [133, 323]. complexity [1132]. complications [672]. components [556]. comprendo [1700]. Computer [1371, 1016]. concern [73, 1240]. concerned [1034]. conclusions [845]. conference [199, 1049]. conferences [1025]. confidence [306, 669, 67]. confident [669]. confidential [95]. Confirmatory [1342]. conflict [622, 803, 892]. confounders [1633]. confuse [567]. confusion [1195, 1616, 1186]. congestion [485]. conjectandi [660]. Connecting [1096]. connection [941]. consensus [1184]. consent [1569, 418]. consequences [1294, 336]. conservation [623]. Consultant [83]. consulting [219]. consults [182]. contagion [1576]. Contagious [32]. contamination [81]. Content [751]. Contents [736, 766, 786, 799, 813, 828, 841, 851, 867, 881, 899, 919, 939, 958, 976, 995, 1014, 1032, 1050, 1068, 1085, 1104, 1124, 1143, 1161, 1180, 1202, 1223, 1241, 1262, 1284, 1302, 1321, 1339, 1358, 1377, 1463, 1484, 1505, 1525]. contest [801, 1515]. continued [821]. continuous [248]. contradiction [473]. contributions [1561]. Control [30, 341, 501]. controlled [32]. controversial [1390]. conversation [1359, 1608, 1579]. convictions [1459]. convincing [1536]. cooked [735]. Cooking [1171]. corona [1577]. coronavirus [1365]. correct [319]. Correcting [831]. correlation [905]. Cortina [1504]. Costing [623, 1294]. Costs [140]. cot [77]. cotton [711]. couch [856]. Could [750, 1509, 239]. count [1330]. counterattack [819]. Counting [712, 892, 996]. countryside [760, 482]. counts [638, 1312, 1186]. coupons [192]. courageous [1645]. course [279, 457]. Court [1607, 1168, 890]. Cousins [374]. coverage [1290, 1576, 1701]. Covid [1634, 1592, 1364, 1501, 1367, 1475, 1425, 1451, 1594, 1380, 1458, 1452, 1590, 1395, 1593, 1383, 1591, 1466, 1382, 1363, 1359, 1579]. Covid-19 [1592, 1364, 1367, 1425, 1451, 1380, 1458, 1395, 1383, 1466, 1382, 1363, 1359, 1634]. cow [258, 411]. Cox [1597, 599, 1595, 1596]. Craig [747]. Cramér [1392]. crampons [1749]. crash [325]. CRC [1236]. cream [1308]. create [1576]. creates [397]. Credit [288]. Cricket [1252]. crime [1751, 186, 724, 1227, 93, 534, 1004, 227, 565]. crime-reduction [93]. crime-scene [534]. Criminal [21, 32, 1233, 78]. criminals [1159]. crisis [1410, 340, 819, 46, 1520, 1154, 1033, 1212, 1242, 709]. Critical [611]. Cross [329, 313, 359]. Cross-over [329, 313, 359]. Crossing [359]. Crossward [721, 734, 749]. Crossword [157, 347, 362, 378, 394, 410, 425, 441, 461, 477, 493, 510, 527, 543, 559, 575, 649, 664, 678, 706, 635, 692, 764, 797, 973, 1622, 1639, 1656, 1672, 1691, 1709, 1726, 1742, 1757, 191, 300, 784, 811, 826, 839, 849, 864, 878, 896, 88, 105, 122, 174, 208, 253, 269, 284, 316, 332, 619, 1158, 1177, 1199, 1220, 1238, 1259, 1299]. crowd [823]. crowds [996, 497]. Crowdsourcing [667]. Crunch [177]. Crystal [402, 462]. cult [1154]. Cultural [753, 1273]. Cultures [1331, 585]. Cup [600, 192]. curates [687]. cure [473, 1117]. Curtains [403]. curves [1616, 1701]. Cuts [1528]. Cutting [774, 1402]. cycle [1195]. Cyclic [1238]. cyclists [707].

damned [295]. Dances [421]. dancing [449]. danger [550]. dangerous [707, 909]. dangers [1750]. D’Arcy [1131]. Dare [465]. Dark [1373]. darts [478]. Darwin [374, 385, 372, 367]. Dastidar [1692]. Data [1243, 1759, 108, 215, 1489, 1396, 1479, 1089, 1188, 1715, 754, 729, 103, 230, 1527, 1722, 1293, 1439, 1361, 133, 1749, 1569, 1574, 891, 112, 1294, 1720, 61, 175, 587, 582, 1021, 1587, 1750, 1686, 273, 1373, 770, 1470, 1513, 1533, 1552, 1349, 902, 1586, 1169, 1389, 1432, 578, 1127, 1590, 745, 1002, 1092, 1343, 9, 1213, 1054, 1145, 1670, 265, 1044, 1062, 1079, 1098, 1118, 1137, 1217, 1255, 1277, 1414, 591, 586, 1397, 1676, 768, 952, 1212, 1381, 1511, 1254, 1580, 1555, 518, 671, 967, 501, 1116, 1456, 110, 1477, 818]. Data-Driven [1479]. Databases [265]. David [1597, 1498]. Day [1113, 644, 153]. days [744]. DC [1203]. dead [1128, 966]. deadlock [679]. deadly [1274]. Deal [209, 1328]. Dealing [1186]. Dear [665, 106]. Death [1187, 112, 77, 653, 930, 725, 1088, 432, 1186, 421]. Deaths [1213, 853, 1381, 225]. debate [61, 739, 1619, 984, 24]. debris [1493]. debt [1089, 229]. decades [1252]. December [1479]. Deception [215, 853]. Deceptive [1734, 1002]. decide [288, 597]. deciding [365]. Decision [1429, 41, 1497]. decision-makers [1497]. Decision-making [1429, 41]. Decisions [1103, 112, 739, 1079, 470, 1112, 1111]. deed [430]. deep [773, 1539]. Deeper [1006]. Deepwater [444]. Deer [186]. Deer-keeper [186]. defeats [1035]. defects [905]. Defending [1169]. Deferring [806]. defines [99]. Deformations [1131]. Defying [101]. degrees [696]. delayed [201]. Deming [501]. democracy [1562, 1283]. democratic [24]. Democrats [1203]. demographic [775]. denied [644]. density [375]. Department [1086]. deportation [198]. descriptions [612]. design [220, 131]. designing [1369]. Detecting [234, 1003, 516]. detective [126]. detectives [1405]. detectors [743]. detects [960]. determinism [1431]. Deterrence [725]. dethroned [1107]. Dettori [153]. development [846]. developments [116]. Diabetes [1095]. Dicing [50]. dicitur [1389]. Dick [385]. Dickens [250, 732]. Dickinson [685]. Did [1615, 1191, 373, 1426, 727]. diet [371]. differences [466]. Different [1342, 304, 833]. differential [1494]. Digital [1405, 197, 901, 1648, 326, 1608]. digits [234]. dilution [109]. dinners [68]. direct [29]. disadvantaged [1246]. disaster [1209, 1089, 444, 965]. disclosure [739]. discoveries [1554, 1370]. Discovering [1369]. discrimination [1617]. discs [273]. disease [1345, 1388, 594]. Diseases [32]. Disgusted [245]. dishes [572]. Dishonesty [215]. disparities [1379, 1286]. disruption [305]. Distance [375, 554, 1392]. Distinguishing [1751]. distribution [1206, 1147, 1165, 1072, 1225, 1109, 979, 667, 1288, 1036, 1723, 1247]. distributions [1000, 333, 1668, 1326]. Disturbing [1116]. diversity [1215, 923]. Dmanisi [695]. DNA [415]. Do [130, 1628, 287, 515, 1287, 1427, 202, 981, 57, 469, 742, 1138, 932, 1364, 537, 659, 1684, 713, 1433, 578, 1465, 808, 134, 258, 1575, 531, 957, 996, 1162, 1269, 572, 630, 1208, 700]. dock [4]. doctor [182, 199, 309, 324, 433]. doctorates [1008]. doctoring [197]. documenting [265]. Documents [803]. Does [1683, 756, 1576, 945, 1246, 12, 388, 1251, 1415, 1465, 406, 1231, 565, 1270, 1381, 907, 1129, 1248, 929, 1074]. doesn’t [1455]. Dogger [1020]. dogs [426]. doing [670, 430]. doku [113]. dollar [42]. domestic [1634, 600]. done [352, 500, 1651, 1476]. donkey [760]. Don’t [587, 1612, 285, 478, 870, 433, 932, 1385, 1716, 1162, 1599, 339, 670]. Doomsday [1544]. doping [546]. dose [624]. double [1648, 1158]. double-edged [1648]. doubt [464]. Down [1220]. downs [1018]. downside [127]. Dr [45, 63, 96, 114, 166, 182, 199, 216, 230, 245, 261, 277, 293, 309, 324, 339, 355, 371, 386, 401, 417, 433, 449, 469, 486, 499, 536, 551, 566, 584, 598, 612, 629, 642, 657, 670, 684, 701, 714, 728, 742, 762, 795, 837, 847, 862, 914, 934, 953, 971, 990, 1009, 1027, 1045, 1063, 1080, 1099, 1119, 1138, 1155, 1174, 1195, 1216, 1235, 1256, 1278, 1296, 1316, 1335, 1354, 1374, 7, 22]. Dr. [517, 781, 809, 824, 876, 894]. Dragons [1720]. draw [650]. Dreadful [186]. dreams [486]. drink [641]. drinkers [278]. drinking [638]. Drive [1759, 490]. Driven [1479]. drivers [1039]. drives [1324]. Driving [1512, 167, 924, 853]. drought [1113]. drown [587]. Drug [81, 484, 464, 1246, 60, 594]. Drugs [466, 932, 1478, 546]. DSLA [1750]. dubious [234]. Duckworth [1252]. Duct [730]. Duke [500]. Dungeness [489]. Durbin [373]. during [1634, 1593]. dust [431]. Dutch [1732]. Dylan [1236]. Dynamics [910, 911].

Eadie [1758]. Early [695, 1514, 40, 358, 1466, 1033]. Earth [1220, 1003, 1680, 241, 622, 580, 1024, 985]. earthquake [643, 609]. earthquakes [960, 926]. earths [639]. easier [861]. East [1217]. eat [681, 242]. eBay [465]. Ebola [769, 857]. economic [305, 1753]. economy [901, 1470, 1513, 1533, 1552, 883]. edge [772]. edged [1648]. Edited [1479]. Edith [1227]. edition [1236]. Editorial [16, 33, 52, 70, 103, 86, 120, 155, 252, 283, 315, 331, 346, 361, 377, 393, 409, 424, 440, 460, 476, 492, 619, 558, 574, 590, 604, 635, 692, 721, 734, 749, 752, 767, 787, 1631, 1662, 1679, 1698, 1713, 1731, 1746, 173, 190, 206, 222, 238, 267, 842, 136, 1644, 852, 800, 814, 829, 866, 880, 898, 918, 1408, 1430, 1450, 1469, 1491, 1551, 1611, 1572, 1589, 509, 526, 542, 648, 663, 677, 705, 299]. Editorial/ [509, 526, 542, 648, 663, 677, 705, 299]. Editorial/Letters/Crossward [721, 734, 749]. Editorial/Letters/Crossword [635, 692]. Education [1305, 1475, 416]. educational [39]. Edward [419]. effect [1360, 718, 1576, 1004]. Effective [1759, 1369, 1396]. effectiveness [61, 436]. effects [872, 624, 1591, 594, 19, 859, 1553]. effort [214]. effortlessly [547]. Efron [458, 455]. eggs [925]. Egypt [538]. Einstein [468]. Eleanor [1738]. Election [949, 128, 1615, 1040, 1403, 1446, 1129, 815, 1127, 1515, 403, 90]. elections [25, 984, 24]. electoral [928]. Elephants [244]. Eleven [122]. Elicitation [100]. Elizabeth [1498]. Elmer [633]. else [498]. emails [1134]. Emanuele [1097]. emergence [326]. Emily [685]. Eminent [358]. emissions [1110]. Emperor [562]. employment [726]. end [1039, 747, 1417, 1431, 1164, 1015, 1224]. ending [1144]. energy [608]. engagement [999]. Engineering [513]. England [978, 1234]. English [227, 616]. Enhancing [942]. enough [344, 214, 457, 1052, 1037]. entrance [1614]. entries [980]. environment [531]. epidemiologists [1508, 1364]. Epidemiology [1393, 1577, 484, 66, 488]. equal [430]. equity [1648]. era [696, 902]. Erebus [727]. ERNIE [151]. error [1267, 815, 1747]. errors [818]. ERS [1203]. escort [95, 140]. Estate [758]. estimate [567]. estimates [1663, 444]. Estimating [1088, 225, 375, 1591, 468, 497]. Estimation [214]. Ethics [1173, 498, 112]. Ethnic [354, 447]. ETS [177]. etymology [1570]. EU [177]. eugenics [374, 1378, 1422]. Europe [178]. European [1607]. Eurovision [1695]. Evaluating [93]. evaluation [165]. Eveline [505]. Even [717]. events [1142, 1070]. ever [537, 348, 1700]. Everest [1749]. everybody [1075]. everything [777, 498]. Evidence [984, 1222, 390, 904, 641, 181, 1332, 1229, 929, 1230, 1228, 470].

Evidence-based [984, 181]. evidence-informed [904]. evolution [695, 384, 372]. evolutionary [741]. evolve [373]. evolving [532]. exactly [641, 640]. exaggeration [1553]. example [1718]. exams [1614]. exceptional [738]. Excess [1452, 225]. Exercise [261]. exit [1040]. exoplanets [1003]. Expect [885, 1138]. expectancy [1017, 1187]. expectations [887]. expected [759]. expedition [727]. experiment [32, 109, 1654]. Experimental [220, 131]. experimentation [48]. experiments [1369]. expert [1558, 731]. explained [1319]. explanatory [49]. exploratory [1342]. Exploring [42, 581, 1004]. exponential [1368, 1072]. Export [598]. Exposing [448]. extra [1427, 1069]. extraordinary [759]. Extreme [1000, 4, 480, 1189, 1070]. Eyam [947, 602]. eyes [732, 162, 1264].

F. [1498]. face [292, 1495]. Facebook [588, 1125]. faces [1362]. facets [1274]. Facing [1750]. fact [11, 354, 1181]. fact-checking [1181]. FactCheck [395]. factors [204]. factory [357]. facts [1163, 235, 242, 1052]. fail [805, 1575]. failings [889]. failure [167, 938]. fair [37]. fairer [435]. fairground [43]. Fairness [998]. faith [1753]. fake [197, 1403, 1052]. fallacy [1043, 703, 831]. fallen [1017]. Fallon [274]. False [1163, 1554, 986, 1454]. fame [1409, 168, 468]. familiar [249, 1320]. family [1159, 416, 586, 1326]. Famous [457]. fans [1507, 1487]. far [560, 1024]. farewell [1499]. farm [1155]. farmer [711]. farmers [306]. Farr [170, 84]. fast [1208]. faster [547, 320, 1382]. Fatal [1747]. faulty [1163]. favoured [655]. Fawcett [1738]. Fear [431, 496, 550]. feat [1550]. feather [981]. February [1236]. feed [1022]. Feeding [481]. feel [1251]. fell [1706]. feloniously [186]. female [1645]. Ferguson [1004]. fever [1369]. fi [913]. Fields [1123]. fight [1114, 769, 1052, 1203]. Fighting [26, 817, 1388]. figure [1390]. Filling [95]. Film [782, 593]. finance [1554]. Financial [148, 519, 110, 28]. Finding [818, 1338, 925]. Fingerprints [534, 928]. finish [1716]. finite [422]. fire [1683, 337, 903]. fired [804]. firings [931]. First [1433, 536, 555, 407, 571, 1716, 233, 614, 1275]. first-borns [571]. fish [242]. Fisher [182, 355, 45, 63, 96, 114, 166, 199, 216, 230, 245, 261, 277, 293, 309, 324, 339, 371, 386, 401, 433, 449, 469, 486, 499, 517, 536, 551, 566, 584, 598, 612, 629, 642, 657, 670, 684, 701, 714, 728, 742, 762, 781, 795, 837, 847, 862, 809, 824, 914, 934, 953, 971, 876, 894, 990, 1009, 1027, 1045, 1063, 1080, 1099, 1119, 1138, 1155, 1174, 1195, 1216, 1235, 1256, 1278, 1296, 1316, 1335, 1354, 1374, 613, 7, 22, 457, 1210, 620, 1378, 506]. Fisher*s [417]. Fishing [673]. Fishy [242]. Fit [1748, 1537, 243]. Fitbit [855, 856]. fits [1702]. Five [1092, 1108, 1136, 1471, 1173]. fix [874]. fixing [874, 501]. flame [560]. flavour [338]. Flawed [889, 641]. Fletcher [570]. Flint [1002]. flip [831]. flock [981]. flooding [1453]. floods [1189]. Florence [1346, 1341, 1349, 328, 1348, 1550]. floresiensis [384]. flowerings [368]. flowers [1721]. flu [1076, 406, 1275]. fly [307, 566]. flying [1289]. focus [1234]. fog [1020]. follow [717]. followers [1248]. Food [1696, 48, 242, 399, 1212, 481]. Football [931, 650, 1664, 145, 227, 192, 942, 91, 656, 858]. force [1384]. Forecast [815, 1309]. forecasting [117, 116]. foreign [1700]. Forensic [83, 1242, 1232, 1229, 234, 1231, 1230, 1228, 1234]. forensics [889]. forests [1427]. forking [1635]. Form [1131]. formed [1527]. formula [533]. forthwith [1319]. fortune [655]. forward [1105]. fossil [369]. fought [1164]. foul [37]. foundations [758, 1026]. founders [1499]. four [552]. four-minute [552]. Fourth [758, 1489]. framework [1557, 1231]. France [625]. Frances [1060]. Francis [374, 502, 1249, 503, 504]. Frankenstein [596]. Franklin [727]. Franscini [1739]. Fraud [215, 516, 928, 817]. fraught [550, 550]. freaks [730]. frequency [948, 333]. Friend [1540]. Friends [1254, 457, 588]. front [1683]. Fuel [846]. Fukushima [496]. Full [1635, 687, 735]. Fundamental [434]. funding [1242]. funds [1069]. furiously [585]. Future [303, 907, 480, 163, 1552, 622, 321, 779, 1204, 591, 887, 1694, 143]. FY19 [1108].

G8 [164]. Galton [511, 502, 1370, 119, 374, 1249, 503, 504]. gambler [703]. gamblers [144]. Game [791, 43, 1132, 1695, 941, 344, 1665, 1706, 1192]. games [964]. gamma [1206, 1288]. gap [891]. gaps [1276]. gas [400]. gates [1271]. Gavin [897]. GDP [404, 1107]. Gender [1314, 1215, 1648, 1276, 1178, 891, 549, 1647, 1207, 1580]. gender-neutral [549]. general [1040, 815, 403, 90]. generalisations [1310]. Generation [1149, 855, 856]. Genetic [414]. genetics [220]. genome [436]. geographic [333]. geography [1214]. geohazards [1573]. geometry [384]. George [658, 437, 700]. Georgiou [1607, 1144, 1677]. geostatistics [640]. German [1303]. Gerrymandering [1172]. get [1039, 1364, 235, 457, 1476]. gets [198, 162]. Getting [794, 874, 822, 200, 931, 177, 560, 1236]. ghettos [354]. Ghosh [1692]. Ghosh-Dastidar [1692]. giants [684]. Gibbon [385]. Giorgi [1097]. giving [672]. glass [718, 203]. gleam [1370]. global [1508, 1093, 319]. GM [711]. go [1754, 1008, 1071, 530, 1717, 382, 957]. goal [248]. God [632]. goes [922]. gold [108, 42, 548]. golden [1151]. goldfish [308]. Goldstein [1333]. good [182, 1091, 1704, 1718, 430, 988, 1330, 418]. Goodhart [1185]. goodness [1537]. goods [1347]. Gordon [1687]. got [808, 444, 92, 29, 759]. Goujeers [1158, 1199, 1238]. governance [326, 754]. government [1293, 1419]. Graciela [615]. grades [945, 1466]. graduate [1649]. Graham [601]. Grand [844]. graphic [1370]. Graphical [1384, 1044, 1062, 1079, 1098, 1118, 1137, 1217, 1255, 1277, 1414, 133]. Graphics [1056, 507, 1236, 1236]. graphs [522]. Graunt [1388]. grave [1057]. Gravitational [888, 1454]. Great [698, 887, 1754, 602, 1056]. greatest [1550]. Greek [1504]. Green [387, 585]. greenhouse [400]. greenwood [351]. Grey [367, 1406]. groaning [775]. ground [1142]. ground-breaking [1142]. group [1366, 1238]. groups [611]. growing [73]. Growth [1131, 1368, 211, 741, 846]. guessing [473]. guest [679].

GUESTimation [679]. guide [1376, 605, 1329]. guides [1236]. guillemots [930]. Gun [724, 921]. guns [1004]. Gwen [1758].

habitable [682]. habits [638]. Hadron [383]. Haigh [1499]. hair [889]. Haiti [643]. Hal [471]. half [570]. Hall [1236, 688, 897]. Hall/CRC [1236]. halves [1695]. hand [1152]. handed [130]. handful [431]. Hans [389, 997]. happen [44, 886]. happened [1661, 683]. happens [370]. hard [996]. harm [1433]. harms [1478]. Harvard [571]. Harvey [1333]. hasty [1310]. having [430, 886]. haystacks [1454]. hazards [521, 1596]. Head [263, 229, 61]. Health [730, 194, 1508, 1660, 1648, 681, 1627, 434, 605, 84]. healthcare [817]. Hector [1236]. heels [229]. height [636]. Helicopters [381]. hell [1650]. helmet [61]. Help [1037, 1439, 1749, 1497, 744, 1682, 1025, 1418]. helped [951]. helps [367]. Henry [250, 465]. heptathlon [556]. her [1134, 342]. heraldic [446]. Herb [1394]. here [957, 427]. heritage [228]. Heroes [1283]. Heroic [947]. heroine [342]. Hey [644]. high [466, 1466]. Higher [1510]. highlights [1097]. highly [1476]. Hill [1483]. hips [243]. hire [433]. hired [804]. Historical [893, 673]. history [1292, 388, 351, 1735, 1470, 1513, 1533, 1552, 1573, 1013, 633]. Hitchcock [593]. HIV [1114]. hobbits [384]. hole [888, 1255]. Hollywood [1178, 983]. Holmes [731]. Holt [231]. home [1507, 870, 98]. homelessness [1246]. homeopathic [109]. Homo [384]. honeycombed [256]. Hood [672]. hope [994]. Horizon [109]. Horses [306]. Hot [1729]. Hotter [321, 1729]. hottest [1729]. House [5, 168]. household [1314]. households [400]. houses [203]. housing [116]. Huff [757]. Human [1607, 256, 292, 695, 241, 337, 384, 857, 1474, 871, 1548]. hundred [660]. Hungry [967, 981]. Hurricane [515, 1088]. hurricanes [519, 322]. hydrant [1683].

Iain [743]. Ian [1524, 1687]. ice [370, 1308]. iceberg [738]. Icelandic [431, 696]. idea [1456, 1477, 1416, 1519, 1392, 1210, 1539, 854, 1438]. ideas [585, 871]. identity [79, 180]. if [1462, 1754, 1418, 478, 886, 759]. ignore [605]. ignored [725]. ignoring [605]. II [226, 177, 1513]. III [1533]. illiteracy [1750]. I’m [668]. imagery [446]. imaging [986]. Imbens [1547]. Immaterials [675]. immeasurable [503]. impact [1367, 1498, 1493, 903]. impacts [1488]. impeding [789]. imperfect [1327]. implants [610]. importance [1341, 151]. important [416, 1579]. impossible [340]. impress [780]. improve [1508, 1287, 1753]. improvement [248]. Improving [48, 1325]. incarceration [1286]. inclusion [1254]. inclusive [1353]. income [710]. Inconsistent [473]. Increasing [531, 1330]. independence [739, 624, 1105]. independent [1579]. Index [94]. India [1245]. Indian [711]. Indicators [24, 178]. indices [5]. individuals [400, 854]. Industrial [1489, 40]. industry [1419, 624, 639, 161]. inequality [910, 1178, 710, 911, 1495, 1392]. Inference [1534, 1043, 67]. inferences [614]. Inflation [1390, 1115]. influence [1078]. influenced [739]. information [95, 474, 579, 507, 595]. informed [1569, 904]. infusions [614]. ingredients [1468]. Initially [1063]. inquiry [239]. inspectorate [357]. intelligence [1265, 1204, 1025]. interaction [19]. interested [1476]. Interesting [82]. interface [835]. internal [622]. international [466, 1019, 627, 1486, 628, 626]. interpretations [1044, 1062, 1079, 1098, 1118, 1137, 1217, 1255, 1277, 1414]. Interstellar [661]. intervals [1272, 67]. Interview [752, 767, 1688, 1687, 1721, 787, 1754, 842, 868, 852, 800, 814, 829, 405, 1599]. interviewer [1267]. intro [1084]. Introducing [1668]. Introduction [422, 1236, 1044]. Intuition [401]. invader [413]. Invalid [646]. invasion [225]. inversion [646]. Investigating [1360, 1345, 272]. investigation [1215]. invisible [1618]. IOM [681]. IQs [343]. Iraq [432, 225]. iris [1555]. ISBN [1479, 1236]. Island [413]. islands [279]. isolated [778]. isotopic [414]. Israney [553]. issue [1714, 1233]. issues [80, 820, 62]. Italian [1367]. Italy [1592]. IV [1552]. I’ve [1476, 1136]. Ivory [244, 667].

Jane [342]. January [1236]. jargon [1474]. Jedi [674]. Jim [373]. Job [612]. jobs [397, 1204, 1295, 1207]. Jockey [274]. John [1499, 747, 1388, 1514]. join [1203, 671]. Joined [726]. Joseph [570]. Journalism [1577]. journey [108, 580]. journeys [667]. joy [1511]. joys [219]. Judge [1160]. Julian [1504]. jump [554]. Jumping [845]. jury [76, 1699, 1160]. Just [1735, 1721, 1612]. Justice [735, 80, 32, 1135, 21, 726, 1086].

Katerina [1504]. Katie [1517]. Katrina [515]. Katrinas [322]. Keep [390, 296, 1651, 1203, 449]. keeper [186]. Keeping [755, 435, 1256]. keeps [1667]. Kenyan [760]. key [1330, 1079]. Kieren [274]. kill [550, 924]. killed [533]. kind [862]. King [716]. Kingdom [570]. Kingdome [407]. kings [1533]. Kingsley [700]. knife [313]. knights [674]. know [1557, 1558, 1364, 1556, 1112, 457, 1226, 1417, 668, 1108, 1385, 1599, 1208, 1118]. know-how [1557, 1558, 1556]. knowledge [1221]. Krebs [686].

L [1498]. labelling [323]. Laconic [633]. lady [328, 1538]. Laird [1486]. lamp [328]. Lance [625]. Land [1465]. Landscapes [351]. language [1087, 446, 1700, 700]. Laplace [746, 1165]. L’Aquila [609]. Large [383]. last [536, 602, 451]. later [294]. latest [342]. Latin [113]. Latine [1389]. launch [1689]. law [724, 912, 234, 1036, 1185, 1717, 1181, 729, 274, 1514]. Lawnmowers [1090]. Laws [236, 296, 1409, 616]. lead [945, 1246, 727]. leaders [164]. leadership [1468]. leads [1554]. league [1626, 290, 942, 931, 227]. leaks [1361]. learn [742, 258, 429, 585, 1125]. learned [788]. learning [795, 1232, 175, 1593, 585, 1539, 1666]. least [1094]. leave [1455, 1733]. leaves [473]. lectures [567]. led [225]. Left [807, 130]. left-handed [130]. legacy [564, 1249, 1559, 3, 546]. legal [683, 1144]. legs [1626]. Leighton [1544]. lens [1383]. Leo [1331]. less [837, 681, 1498]. Lessons [1681, 1013, 1273, 340, 857]. Let [1346, 1733]. Letters [14, 34, 53, 71, 87, 104, 121, 137, 156, 172, 189, 207, 268, 509, 526, 542, 635, 692, 648, 663, 677, 705, 721, 734, 749, 973, 1398, 1420, 1441, 1460, 1481, 1502, 1522, 1542, 1564, 1581, 1601, 1621, 1638, 1655, 1671, 1690, 1708, 1725, 1741, 1756, 299, 252, 283, 315, 331, 346, 361, 377, 393, 409, 424, 440, 460, 476, 492, 619, 558, 574, 590, 604, 764, 784, 797, 811, 826, 839, 849, 864, 878, 896, 915, 936, 955, 1336]. Letters/ [252, 283, 315, 331, 346, 361, 377, 393, 409, 424, 440, 460, 476, 492, 558, 574, 590, 604, 764, 784, 797, 811, 826, 839, 849, 864, 878, 896, 915, 936, 955, 1336]. Letters/Puzzle [1398]. level [49, 1233, 685]. Lewis [311, 1252]. LGBT [1254]. liable [1071]. libel [616]. Library [1117]. licks [693]. LiDAR [351]. lie [1718]. Lies [295, 1734]. Life [379, 112, 777, 772, 794, 1017, 1187, 1060, 421, 179, 1350, 551, 1155]. life-savers [179]. light [1346, 324, 1719, 675]. lights [1667, 1287, 247]. like [985]. likelihood [625]. likely [1039, 550, 571]. Lilla [617]. limit [383]. limits [1221, 181, 1228]. Line [169, 1062]. lines [1577]. link [272]. linkage [1477]. links [1378, 726]. Lipsey [327]. listen [1612]. listener [605]. Listening [595, 429]. listing [365]. lists [679]. literacy [989, 1334]. Literary [732]. literature [1383]. little [517, 1735, 1406, 535, 247, 1734, 472]. live [241, 257, 969, 499]. lives [1294, 750, 653]. Living [1332, 578, 386, 689]. location [1074]. lockdowns [1591]. LoCurto [617]. log [979]. log-normal [979]. logarithmically [386, 630]. London [707, 175, 464, 564, 257, 488, 335, 546, 903]. long [499, 554, 344, 1019, 518, 789]. long-term [1019, 518]. longevity [1548]. look [1538, 1386, 1418]. Looking [308, 1721]. Lord [327, 686, 217, 1022]. lose [600, 1447]. lost [144, 727, 1700, 126, 1099]. lot [1255]. lottery [1684, 535, 149]. love [1706]. low [624]. low-dose [624]. lower [1510, 945]. Luck [943, 808, 1192]. lucky [457]. Lunatics [84]. lunch [430].

Macaulay [1164]. MacFeely [1579]. Machine [1232, 513, 1666, 1264]. machines [1204, 788]. made [517, 1498, 1586]. Madness [1221]. magistri [642]. mail [29]. major [962]. make [1251, 1194, 1736, 470]. makers [1497]. makes [1475, 1167, 1069]. Making [569, 474, 933, 1618, 1070, 398, 1429, 41, 514, 507, 319]. Malawi [165]. males [336]. Malthusian [1148]. man [617, 731, 1022]. manage [640]. Managing [11]. Manchester [1248]. many [756, 816, 1364, 241, 1571, 1008, 1078, 693, 336, 257, 258, 322, 1575, 966, 395, 1004, 497, 1454]. maps [570, 933, 256]. marathon [856, 550]. march [312, 1221]. Marcus [1566]. Maria [1088]. marijuana [922]. marine [673]. Mario [1504]. Marital [305]. mark [1167, 1402, 553]. markets [41]. Markov [514, 422]. marriage [650]. Martin [633]. Martyn [1483]. mass [1576]. masses [611]. massive [1536]. Mastrodomenico [1673]. match [1231, 382]. matches [659]. matching [1147]. Math [970]. mathematician [325]. mathematics [1524, 1754]. Maths [1578, 426]. Matrixx [506]. matter [12]. mattered [497]. matters [1314, 99, 1214, 1041]. may [1153, 418, 789, 1208, 1459]. Mayhew [250]. mazurkas [514]. McGhee [1657]. McKay [1688]. me [566, 629, 262, 307, 1393, 754]. mean [1415, 1428, 1231, 1073, 1118, 504, 1270]. meaning [385]. measure [794, 901, 710, 1290, 731, 1612, 1184, 989, 23]. measured [1660]. Measurement [1267, 99]. measures [1229, 93]. Measuring [1751, 39, 1304, 1042, 1632, 275, 985, 466, 503]. medals [907]. media [1290, 1576, 168]. median [1276]. median-sized [1276]. Medical [1569, 470, 65]. Medicine [181, 163]. medieval [421]. meet [457]. Meeting [129]. meets [1419, 1467]. megacities [480]. megalithic [363]. melts [370]. members [1490]. memes [661]. Men [549, 442]. Mendel [372]. mention [285]. Mentoring [1649]. mere [1239, 1384]. merger [888]. Mervyn [405]. Meshimbayeva [1721, 1166]. meta [1362, 1194]. meta-analysis [1362, 1194]. metalog [1668]. method [1455]. methodological [1410]. methods [96, 64]. Mexico [1312]. Miami [513]. Michelle [1657]. microscopic [889]. Microscopium [1570]. mid [162]. mid-2007 [162]. MIDAS [108]. middle [278, 1632]. might [1023, 1417, 1189]. Migrants [713, 1019]. migration [98, 1056]. mile [552]. Military [1275]. milk [304, 614]. millennials [1059]. Millicent [1738]. million [528, 1248]. millionaire [620]. millions [511]. Minard [312]. Mind [1087, 61, 1145, 185]. minds [309, 1311]. minimum [641]. Mining [1291, 108]. minute [552]. minutes [168]. Misfortune [1053]. Misgivings [672]. mislead [1415, 1516]. Misleading [78, 417]. misrepresented [119]. missing [945]. missionary [1652]. mistake [947, 770, 1077]. mistaken [180]. mistakes [1488]. mobilise [1587]. mobility [1367]. Moby [385]. model [321, 56, 672, 1703, 754, 367, 1596, 437, 759]. modeling [249]. Modelling [1592, 872, 859, 745]. models [133, 689, 1310]. Modern [816, 1351, 794, 901, 1187, 1130, 1614, 239, 754]. moments [698]. money [715, 200, 1077]. Moneyball [524]. Monitoring [165, 112, 131]. monkeys [528, 685]. month [1039]. Monty [688]. Moo [258]. Moo-cow [258]. mood [583]. Moon [566]. moons [1635]. Moore [465]. moral [570]. Mortality [84, 1452]. mosaic [256]. Moscow [312]. Moser [217]. most [978, 1094, 1023, 416, 778, 1476]. mother [1652]. mothers [1645]. moulds [1653]. move [1645]. movement [690]. movements [753]. mover [1253]. movie [524]. movies [601]. Moving [549]. much [681, 320, 579, 1058, 242, 319, 565, 472, 922, 1529]. multi [925]. multi-yolk [925]. multilevel [249]. Multiple [473, 1070]. multiplicity [682]. mumble [469]. Mummy [258]. municipal [484]. murder [732]. murders [335]. murky [793, 1002]. Murrell [1236]. Music [1681, 607, 514, 1092]. musical [1409]. must [1475]. My [1637, 239, 457, 1016, 1295, 588, 1037, 150, 1754, 1509]. mystery [732, 126]. myth [1664, 354]. myths [442, 564, 448].

N [1498]. nail [789]. name [333, 1685]. names [1023]. Nan [1486]. Napoleon [312]. Narrative [1759]. Nate [702]. Nathan [631]. nation [1739, 583, 1022]. National [132, 149, 243, 965, 931, 25, 194]. nationwide [395]. natural [1654]. nature [383]. navs [750]. nearly [433]. necessarily [645]. need [743, 1365, 1229, 1651, 1026, 1599]. needing [501]. needle [1454]. needs [1480, 1686, 1395, 129, 1242, 1244]. negative [636]. Neighbourhood [116, 416]. neighbourhoods [483]. neighbours [515]. Neither [1465]. Neonicotinoids [652]. network [960, 1348]. networks [696, 1112, 1612, 1647]. neutral [549]. never [742, 1417]. Newcastle [968]. News [1, 20, 38, 107, 125, 141, 159, 176, 193, 210, 224, 240, 255, 271, 286, 302, 318, 334, 349, 364, 380, 396, 412, 428, 443, 463, 479, 495, 512, 529, 545, 561, 577, 592, 606, 621, 637, 651, 666, 680, 694, 752, 767, 708, 723, 737, 1401, 1424, 1445, 1464, 1485, 1506, 1526, 1546, 1568, 1585, 1606, 1625, 1642, 1659, 1674, 1693, 1711, 787, 1728, 1744, 842, 852, 800, 814, 829, 882, 900, 920, 940, 959, 977, 1263, 1285, 1322, 1340, 1052, 1142, 868, 1479]. News/ [868]. News/Interview/Editorial [752, 767, 787, 842, 852, 800, 814, 829]. next [1509, 597, 790]. NFL [1537]. Nguyen [1479]. nice [1061, 420, 1352]. night [1211]. Nightingale [1346, 1349, 328, 1348, 1434, 1550, 1338]. Nightingales [1351]. Nine [805]. No [279, 1093, 1507, 1700, 1626, 537, 1588, 1433, 1144, 209]. Nobel [442, 1547]. nocere [1433]. noise [275, 1003, 581, 1454]. non [1008, 1534, 1433]. non-random [1534]. non-US [1008]. Nonhlanhla [1705]. normal [1225, 979, 1531, 1723]. North [722, 727]. North-West [727]. Norwegian [24]. Notebook [1165]. Novel [1056, 342]. novels [732]. November [1236]. Nowcasting [583]. nuclear [624, 489, 161]. Nudge [1323]. Nuijten [1743]. number [322, 403]. Numbers [1479, 1706, 1178, 317, 505, 6, 1718, 337, 1684, 1388, 1472, 531, 1717, 496, 774, 342]. nurse [182]. nutrition [399]. Nylon [1259].

obesity [406, 722]. obey [388]. objective [1229]. objects [1350]. obscurity [1409]. observational [501]. obtaining [418]. ocean [773, 1058]. odds [1200, 553, 1043]. odds-setter [553]. offences [1751]. offenders [21]. offending [726]. Office [98]. Official [25, 1646, 279, 82, 1015, 47]. Oh [1588]. oil [444, 319]. Old [871]. Olympian [553]. Olympic [907, 464, 556, 1472, 548]. Olympic-style [1472]. Olympics [555, 175, 320, 549]. One [1451, 1603, 382, 1105, 1499, 537, 467, 722, 1078, 945, 1007, 1372, 1604]. one-sided [1372, 1604]. online [703, 228, 143]. only [744, 162]. oop [1473]. open [980, 986]. Opinion [232, 1479]. opinions [128]. Opioids [1286]. opportunities [1386]. opportunity [301, 110]. optimist [1608]. option [750]. Oracle [843]. orang [843]. orang-utan [843]. orangutans [429]. organic [304]. organisation [1651]. origins [1456, 1227]. Oscar [1211, 260]. other [1478, 682]. ought [1185, 759]. Our [532, 1705, 927, 323, 1499, 741, 1661, 822, 99, 1204, 1629]. outbreak [1345]. Outcault [617]. outcome [802]. outcomes [1129]. outcry [610]. outliers [1496, 1091]. outperform [546]. overcome [1578]. Overcoming [546]. overconfidence [668]. overfitting [1554]. overlooked [1434]. oversell [1347]. Owen [1236]. owns [729]. Oxford [1236]. Oxygen [1749].

P [658]. Pachedu [1167]. painted [239]. painting [675]. Palma [710]. Pandemic [1593, 1367, 1363]. Pandemics [1368, 421, 1076]. Panic [294, 970, 1034]. Paper [1528, 1194]. paperback [1479]. parachutes [1508]. paradox [1380, 836]. ParalympicsGB [1051]. parameter [1250]. Pareto [717]. park [1683, 419]. part [869, 1112, 823, 1579, 1513, 1533, 1552]. participation [564]. particularly [1636]. party [776]. Passage [727]. Passionate [1448]. past [907, 887, 673]. paths [1635]. patients [473]. pattern [335]. patterns [923]. Paul [1236]. Pawprints [426]. pay [1683, 1276, 565]. Peace [622]. peaks [1127, 1404]. Pearson [421, 1250]. pedantry [862]. pedestal [442]. peer [1383, 873]. peer-reviewed [1383]. Penalty [1663, 124, 653, 725, 1042]. pension [806]. pensions [303, 46]. People [563, 433, 469, 756, 1364, 241, 1587, 257, 134, 1246, 636, 966, 1753, 1476, 892]. perceive [630]. percentages [1465]. perfect [399]. perform [1615]. performance [140, 487]. performing [997]. personal [1215, 1095, 110]. personalities [277]. personality [1016]. Perspective [893]. perspectives [732]. persuasion [1098]. Petchey [1236]. Peter [897]. Petri [572]. pharmaceutical [466, 10]. phase [177]. phenology [368]. philanthropy [672]. Philip [539]. philosophies [64]. phish [1535]. physics [1436]. Piano [552, 670]. Pick [1308]. Pictish [446]. pictures [593, 944]. Picturing [593]. piece [1326]. piecewise [1072]. Pietro [1532]. Pill [272]. pills [183]. Piscatori [642]. pitfalls [912]. Pitt [524]. place [1751, 1646]. plague [947]. plagues [421]. plan [1239, 1025, 1357]. Plane [489]. planets [682]. planning [319]. plant [489]. Plastic [1058]. play [1684]. player [670]. players [1684, 91]. Playfair [1057]. Playing [535, 768, 1183, 869]. plea [103, 1328]. please [805]. pleasure [1462, 1006]. pledge [1629]. plots [49, 169]. plugged [1361]. PMs [1504]. pockets [1006]. Poetry [505]. point [567, 1179, 560]. points [1151, 1210, 1250, 951]. poison [727]. poker [1192]. Police [1244, 1379]. policy [904]. Political [1483, 801, 1411, 1059]. Politick [407]. politics [1312]. poll [1040, 232]. Polling [128]. Polls [1403, 90, 1615, 1446, 1035]. polluting [482]. Polyconic [617]. Pompeii [638]. ponders [199]. Pooling [1716]. pop [1681]. poppycock [247]. Pops [693]. popular [978, 1023]. popularity [801, 1515]. population [1735, 531, 367, 116]. populations [241, 673]. positives [986]. possibilistically [270]. Post [1064, 1049]. Post-conference [1049]. Post-truth [1064]. poster [627]. postmortem [1446]. potato [856]. poverty [303, 1696, 1042, 1184]. Power [1036, 1428, 1348, 247, 489]. Power-law [1036]. practice [1508, 30, 1173, 329, 418]. practices [1580]. praise [96, 1737]. pre [1015]. pre-release [1015]. Precision [640]. predict [907, 946, 926]. predictable [802, 1076, 335]. Predicting [260, 403, 322]. prediction [887, 938]. predictions [321, 1637]. predictive [745, 1310]. predictors [108]. pregnancy [1018]. Premier [227]. premises [792].

Prepare [1493]. prepared [1474]. present [887]. President [122, 596, 597]. presidential [1403, 1001]. Press [1236]. pressure [761, 1529]. pressure/a [761]. prevent [1682, 1310]. Preventing [1239, 516]. prevention [1532]. price [5, 874, 641, 343, 466, 656, 1426, 632]. price-fixing [874]. prices [466]. pricing [641]. Primary [728]. Primum [1433]. Principal [556]. principle [522]. prior [31]. Prisoner [140, 95]. Pritchard [1599]. Privacy [273, 418, 1574, 1494, 754]. private [352, 84, 1459]. privatization [213]. Prize [88, 105, 122, 174, 208, 253, 269, 284, 316, 332, 1158, 1177, 1199, 1220, 1238, 1259, 1299, 1486, 54, 138, 191, 300, 1547]. prizes [442]. probabilistic [270]. Probability [1530, 1549, 1510, 1571, 59, 226, 76, 861, 1514]. probable [534]. problem [195, 688, 246, 697, 963, 962, 1276, 1005]. problems [1320]. process [48, 501]. Product [1702, 390]. profession [471, 884]. professionalise [1397]. professor [1754, 1333, 363]. profiles [415]. profiling [447]. profit [1462]. Profligate [94]. programme [1518]. programming [1488]. programs [433]. progress [391, 789]. projection [140]. projections [1735]. Promiscuous [860]. promises [912]. promising [1239]. Promote [1194]. properly [1569]. proportion [435]. proportional [1596]. prosecutor [831]. prosper [499]. prostate [1637]. Prosthetic [1152]. Protecting [164]. protection [1294]. proven [941]. Psychology [821, 1717]. Public [1479, 84, 1215, 743, 1704, 1576, 1719, 740, 418, 1459]. publication [860, 127]. publications [1597, 1064]. Publish [201]. published [729]. Puerto [1089, 1105]. puff [757]. pull [627, 490]. pull-out [627]. Puncture [1588]. Pundits [669]. punishment [227]. purpose [199, 1748]. Purposes [64]. pursuit [1532, 1135]. put [862]. Putin [1127]. Putney [395]. Putting [1704, 442]. Puzzle [15, 35, 72, 915, 936, 955, 1281, 1336, 1398, 1421, 1442, 1461, 1482, 1503, 1523, 1543, 1565, 1582, 1602, 1697, 1610, 1630, 1643, 1678, 1712, 1730, 1745, 54, 138, 1530, 1549, 1510, 1571, 1375]. Pye [505].

Q&A [1758, 1743, 1172, 1234]. Quality [1234, 1170, 9, 97]. Quantifying [366]. quantity [1074]. quantum [1436]. queens [1533]. question [79, 1509, 1086]. Questions [1623, 1657, 1640, 1673, 1692, 1583, 1727, 1517, 1566, 1710, 1737, 537, 674, 1619, 1547, 1255, 1599]. quick [1446, 1312]. quite [854, 759]. Quomodo [1389].

R [520, 1391, 1236, 1236, 1149, 1473]. R. [613, 1378]. Race [1537, 181]. Racial [1379, 1699]. Racing [274]. racket [906]. Radar [539]. radiation [624]. radio [605]. RAE [261]. Raghib [1640]. Rail [213]. raining [1252]. raise [406]. Ramanujan [522]. random [1534, 667, 1210, 1055]. Randomisation [438, 1575]. Randomised [855, 494, 790]. Randomising [1114]. Rank [1023, 1703]. Ranking [91]. Rao [520, 1391, 1392]. rapid [59, 226]. rare [415, 639]. rate [1454]. rated [832]. rates [1187]. Rathbone [1738]. ratio [871]. rats [756]. RCTs [1553]. re [1402]. re-mark [1402]. react [337]. Read [146, 202]. Reader [688]. Readers [821, 427]. Reading [1435, 783]. ready [1599]. Real [175, 66, 1694, 1634, 197, 960, 1403, 1347, 1476, 353]. Real-time [66]. realistic [273]. reality [319]. Really [261, 756, 304, 924, 290, 1248, 727, 319, 1579, 94]. reasonable [464]. Recapturing [746]. Recession [310]. reclassified [683]. recommendations [1173, 929]. Recommended [783]. record [369, 1548]. record-breaking [1548]. recordings [1005]. records [549, 673]. recreational [1478]. Red [365, 1095, 1664, 834, 247, 367]. redistricting [1172]. reducing [400]. reduction [93]. ref [227]. refereeing [701]. References [1598]. referendum [1404, 1041]. reflect [1459]. Reflections [149, 77, 1049]. Reflexive [1199]. reform [1273, 434]. regarding [95]. registered [1273]. Registration [264]. regression [1078, 1428, 504, 1438]. regs [1244]. regularisation [1438]. regulation [1303]. Regulator [1234]. relationship [42]. relative [1290]. release [1015]. relevant [211]. Reliability [1230, 1702]. reliable [1573]. relic [741]. religious [446]. remarkable [1060]. Remembering [1560, 1595, 897]. reminded [1579]. Remote [1593]. Renewable [608]. Reopening [515]. Repeal [236]. repeating [1735]. replicable [1194]. Replication [1428, 1410]. report [688, 681, 1110]. reported [1590]. reports [1273]. reproducibility [1325, 819]. request [1108]. research [1508, 1569, 1410, 200, 454, 1380, 611, 1194, 1342, 921, 1242, 572, 616]. researcher [115, 1459]. residents [886]. resiliency [875]. resisted [186]. resources [1716]. respond [821]. Response [1011, 1029, 1047, 1065, 1082, 1101, 1121, 1237, 1258, 1280, 1298, 1318, 1355, 992, 1140, 1157, 1176, 1198, 1219, 1375, 25, 911, 1330]. response/Puzzle [1375]. restaurants [805]. result [1260, 37, 833]. results [1404]. retired [1378]. return [1201, 1531, 355]. revealed [986]. reveals [1452]. Review [1440, 1479, 782, 1010, 1028, 1046, 1081, 1139, 1156, 1175, 1197, 1218, 1755, 98, 1100, 1120, 1257, 1279, 1297, 1317]. reviewed [1383]. reviewing [873]. Reviews [17, 36, 55, 69, 85, 102, 123, 139, 154, 171, 188, 205, 237, 251, 266, 282, 298, 314, 330, 345, 360, 376, 392, 408, 423, 439, 459, 475, 491, 508, 523, 541, 557, 573, 589, 603, 618, 634, 647, 662, 676, 691, 704, 719, 733, 748, 796, 810, 825, 838, 848, 863, 877, 916, 935, 954, 763, 1620, 1707, 1724, 1740, 972, 991, 1600]. revisited [1127, 297]. Revisiting [947]. revitalise [483]. revolution [1273, 1489]. Revolutionising [1660]. rhetoric [1239, 1411]. rich [110]. Richard [716, 632]. Richthofen [468]. Rico [1089, 1105]. Ridge [1438]. rigging [984]. right [837, 368, 198, 177, 1534, 668, 669]. rights [890, 1188, 1607]. Rigorous [1310]. RIP [404]. rise [1115, 1245]. Risedronate [262]. Risk [1558, 1557, 738, 1556, 875, 273, 1268, 365, 406, 609, 1666, 278, 1183, 1303, 518, 161]. risk-based [1303]. risks [610]. risky [1512, 1478]. road [1498, 872, 1287, 1051, 1719]. roads [707, 656]. Rob [1673]. Robbing [565]. Robert [1754]. Robin [672]. role [1267, 745, 24]. roll [1520]. Romano [426]. roots [387]. Rosling [389]. roulette [1191]. row [625]. Royal [1738]. RPI [94]. RSS [1260]. rugby [951]. rule [1416, 717, 858]. rules [388]. run [547, 320]. runner [856]. Running [317, 1200]. runs [395]. Russia [1747, 1404, 1452]. Russian [1127]. Ryanair [1055].

S [1269, 309]. sacked [931]. sacrifice [947]. safe [307]. safely [530]. Safer [950]. safest [750]. Safety [6, 337, 179, 233, 213, 1287, 1719]. sagas [696]. Sagittarius [644]. said [1754]. Sailing [350]. sake [1174]. salaries [1078]. Sally [78]. Salt [681]. Sam [1177, 1220, 1259, 1299, 1583]. same [554, 1162]. sample [1037]. samples [1534]. Sampling [80, 214, 375]. sanctions [432]. sands [388]. sandwiches [278]. Santos [1754]. SARS [66]. sat [750]. satisfied [961]. Sautoy [1566]. save [1293, 744, 1111, 750, 1666]. savers [179]. saves [433]. Saving [1294, 653]. say [1381]. says [1371]. scale [1634, 1696, 1158]. scandal [1732, 500, 853, 610, 1125]. scanner [1492]. scares [196]. Scatterplots [187, 297]. scene [534]. scenes [975]. Scepticism [1563, 472]. schedule [1025]. Scholar [263]. scholastic [858]. School [290, 416, 945, 1466]. schools [166, 1501, 352]. sci [913]. sci-fi [913]. Science [215, 1382, 1234, 1527, 1091, 869, 1563, 986, 2, 1617, 160, 1325, 1389, 473, 92, 129, 624, 819, 988, 835, 740, 1171, 436, 1230, 1397, 1242, 1579, 1254, 609]. sciences [1026]. scientific [74, 1455, 1348, 1154, 616]. scientifically [450]. sclerosis [473]. Score [605, 288]. Scoring [951]. Scotland [95, 739, 140, 164]. Scott [1498]. Scottish [276]. Scrambled [925]. scrambling [1439]. screen [470]. screenings [929]. scrutiny [74, 140]. Scutari [1434]. sea [1111, 445, 930]. seagull [690]. seals [367]. search [1003, 1193, 1555]. season [908]. Seat [296, 236]. seating [1055]. Second [908, 747]. secondary [1501]. secret [1399, 1422, 1443, 1462]. Secretary [1069]. Securing [792]. security [399, 481]. Seduced [544]. see [636, 754, 1162, 1418]. seeds [1130]. Seeing [1664, 254]. seem [759]. sees [324]. selection [962, 1699]. self [924, 1330]. self-driving [924]. self-response [1330]. semen [730]. Senate [1615]. sense [1070]. separation [696]. sequence [1449]. series [117]. Serious [964]. serve [1153, 946]. Service [264, 1432, 194]. Services [28, 110]. set [1527, 791, 1555]. sets [1116]. setter [553]. Seven [556, 625]. seventh [289]. severe [1309, 1113]. Sex [832, 1676, 871, 1580]. Sex-rated [832]. Sexes [135]. sexy [471]. shadow [1035]. shadows [1243]. shaker [1253]. Shakespeare [528]. shall [536, 969]. shallow [387]. shape [653]. Shaping [739]. shared [1292]. sharing [1569, 902, 754]. Sharp [183]. Sheconomics [487]. sheet [370]. Sherlock [731]. shibboleth [1614]. shiny [1487]. Shipman [3]. shock [1487, 1073]. shoot [670]. shootings [1576]. shootouts [1663]. shoots [387]. shop [353]. short [1366, 844, 1212, 518, 1212]. short-changing [844]. short-term [518]. shots [1703]. Should [1019, 296, 681, 76, 1586, 242, 725, 1753, 470, 321, 671]. shoulders [684]. shouldn’t [44]. show [1144]. Shrinking [485]. shy [182]. sickest [391]. Side [594, 1216]. Side-effects [594]. sided [1372, 1604]. sidelines [1458]. sift [1454]. Sigma [301]. sign [644, 1144]. signals [1152]. significance [569, 1272, 287, 685, 1059, 12, 1193, 1224, 506, 1629, 627, 720, 628, 525, 1613]. significant [569]. silence [429]. silly [537]. Silver [702]. simple [96].

Simply [1119]. Simpson [836, 419]. since [225]. sincerely [202]. Singing [674]. single [1194, 1553]. sings [595]. sinh [1225]. sinh-arcsinh [1225]. Sir [1597, 665, 743, 1595, 374]. Siracusano [506]. Six [340, 1306, 696, 301]. Size [99, 576]. sized [1276]. sizes [497]. Sizing [243]. skill [1192]. skills [1700]. Skin [1267]. skulls [695]. sky [1721]. Slams [844]. slavery [816, 1164, 521]. Sleeping [1356]. sleeves [1520]. slender [1536]. slowdown [1093]. slowly [1705]. Small [1488, 1717, 353, 584]. smart [672]. smartly [1492]. smartphone [960]. smashes [659]. Smashing [906]. smile [1131]. Smoke [162, 922]. smugglers [563]. soaring [343]. Soccer [1132, 1487]. Social [1646, 1612, 1135, 1026, 696]. societies [358]. society [579, 1304, 256, 1738]. soggy [1532]. solved [1520]. Somalia [709]. some [1509, 1328, 820, 352, 672]. someone [924]. Something [1031]. Sound [1474, 429]. source [1555, 1233]. Soviet [624]. sown [1130]. Space [1493, 1680, 530]. spake [1319]. spam [26]. Spanish [1275]. spawned [395]. Speaking [1411]. species [1645, 365]. spectre [1380]. speech [1001, 1005]. speeches [1211]. speed [485]. spell [339]. spending [1719]. spill [444]. spoiled [276]. Spoilt [304]. Sporting [551]. sports [564]. spot [1663, 488]. spread [148, 66]. spring [1126, 538]. Springer [54, 88, 105, 122, 138, 174, 191, 208, 253, 269, 284, 300, 316]. square [1537]. squares [113]. Squeezed [1177]. SSA [1490, 1669]. stack [531]. stadiums [483]. stained [203]. stair [685]. Stakes [147]. stand [1626]. standard [211]. standards [1527, 39]. Stander [1504]. star [1681, 126]. starling [289]. stars [1721]. start [1426, 427]. Started [1236]. starting [1179]. Stat [1260]. state [806, 341, 1307]. statement [1007, 1471]. States [434]. staticians [754]. station [903]. statistic [395]. Statistical [1292, 1272, 928, 1738, 507, 1488, 67, 569, 1501, 1096, 250, 739, 732, 1173, 1468, 1043, 62, 758, 1590, 1395, 168, 407, 1059, 1703, 580, 358, 129, 779, 888, 585, 819, 1413, 239, 228, 3, 126, 989, 1224, 1580, 287, 889, 1334, 506]. statistically [788, 534]. Statistician [1129, 1110, 1248, 1652, 657, 1399, 1422, 1443, 1462, 310, 112, 420, 776, 44, 1449, 1739, 1038, 1417, 778, 620, 969, 1164, 239, 524, 1160, 437, 65, 25, 8, 83, 10, 232, 28, 218, 47]. Statisticians [720, 628, 203, 744, 1520, 1448, 373, 517, 537, 301, 341, 456, 611, 457, 1015, 671, 1636, 745]. Statistics [1462, 292, 75, 454, 1627, 185, 163, 1412, 1406, 611, 890, 1312, 1716, 1135, 1106, 1168, 161, 195, 279, 1480, 82, 74, 184, 166, 699, 821, 913, 869, 40, 593, 1749, 276, 1563, 830, 1754, 388, 317, 563, 301, 514, 25, 570, 158, 415, 1702, 89, 254, 135, 381, 1570, 1475, 2, 884, 1253, 26, 698, 761, 98, 1130, 168, 1667, 1682, 1227, 24, 1265, 1516, 92, 872, 1752, 1275, 385, 567, 725, 78, 1329, 319, 1323, 1295, 1236, 1734, 1706, 1753, 1436, 965, 1717, 1356, 1434, 1171, 445, 1467, 436, 832, 1154, 854, 952, 997, 1022, 1015]. statistics [1123, 1172, 1493, 1579, 1136, 1254, 925, 639, 1535, 951, 350, 817, 342, 987, 1313, 576, 627, 1486, 1646, 132, 628, 248, 465, 1353, 264, 357, 643, 1418, 626]. Statisticum [1570]. statophobia [1117]. stats [791, 544, 535, 1681, 1736, 1105, 1181]. Status [58, 683]. stealth [179]. steer [1474]. Stefano [1739]. stellar [1003]. step [170, 1105]. steps [1105]. Stereotypically [847]. Stern [1252]. Steve [1579]. Stewart [1524]. sticker [192, 1487]. sticks [1308]. Still [1200, 1651, 795]. Stone [405]. Stonehenge [285]. Stones [446]. stood [1529]. Stop [1534]. stopped [886]. stopping [858]. stops [1379]. stories [1343, 605, 1413]. storm [399]. story [1500, 1093, 65, 132, 83, 1617, 235, 10, 1457, 1437, 28, 218, 47, 151, 363, 126, 115]. storyteller [310, 997]. Storytelling [1759, 1396]. strange [854]. strategies [381]. street [1287, 533]. streets [395]. Strength [761, 1472, 1721]. Stress [780]. Strictly [987]. strikes [562, 489]. striking [322]. stripped [142]. Strong [1459, 697]. stronger [1382]. Structure [133, 1386]. Student [506]. students [1649, 567, 1019]. studies [1593, 501]. study [1095, 1466]. Studying [1005]. stuff [1476]. stupid [1737]. style [514, 1472]. subjective [1612]. subtle [313]. success [1051]. suffragists [1738]. sugar [1095]. suicides [711]. suit [548]. summit [164]. Sun [1024]. Super [1061]. superhuman [442]. supermarkets [353]. supply [1366]. support [929]. Supporting [1353, 983]. Supreme [890]. surname [923]. surnames [333]. Surprise [1040]. surveillance [164]. Survey [1646, 407, 243]. surveys [1314, 1293, 1628, 674]. Survival [391, 1329]. Survive [194]. Suspect [1404]. suspicious [1449]. sustainability [241]. sweat [729]. Sweet [1734]. swimmer [548]. Swimming [1467, 1716, 548]. swine [406]. sword [1648]. Symbol [446]. symmetry [101]. Synchronicity [1193]. syndrome [908, 58]. syndromic [164]. synoptic [195]. Synthesis [97]. Synthetic [1432]. Syrian [803]. system [1627, 1191, 1395, 1033, 1529]. systematic [131]. systems [1303, 435].

Tables [522, 1626, 290, 942]. Tagging [198]. tails [710]. Taiwan [1562]. Take [1268, 124, 629]. taken [1204]. takes [1607]. tale [223, 890, 1002, 247]. tales [329, 1067]. Talk [18, 1344, 1722, 1031, 1694]. Talking [166]. Tanzania [1181]. tape [730]. Targeted [1315]. tasting [1538]. tax [179]. taxation [291]. TB [259]. Tea [614, 1538]. Teach [158]. Teaching [913, 1537, 1315, 1097]. Team [1167]. teenagers [1190]. telescopes [1617]. television [730]. tell [197, 465, 290, 1207, 673]. Telling [1413]. template [407]. Ten [403]. tennis [659, 844, 1153]. term [1019, 518]. Terror [727]. terrorism [1453, 1274, 1385]. terrorist [447]. terrorists [1090]. test [1537, 1372, 1604, 281]. Testing [383, 1733, 484, 1366]. Tests [1366, 1501, 1365, 220, 1372, 1604, 280, 789]. text [1686]. Thanos [1148]. their [199, 391, 1591, 367]. them [296, 236, 1294]. thematic [570]. theorem [632, 1151]. theory [488]. There [1185, 1346, 579, 1534, 1058, 319, 497, 353, 1520]. thin [1536]. thing [568]. Things [1112, 739, 340, 134, 1575, 1108, 1136, 435, 759]. think [1439, 1365, 415, 540, 1226, 788, 854, 1189]. Thinking [270, 1705, 1163, 1370, 779]. third [722, 1236]. Thom [363]. Thompson [1131]. those [247]. thoughts [27]. threaten [1181]. Three [1468, 49, 660, 614]. three-level [49]. thrived [1646]. Thus [1319]. Tick [1261]. ticket [1039]. ticklish [1005]. ties [1266]. Tigray [1627]. Tim [231]. Time [1344, 117, 1397, 1577, 1443, 368, 388, 177, 66, 960, 108, 494, 1411, 1153, 228, 1034, 1107, 1301, 1381, 1299, 1277]. Time-series [117]. Timeline [699]. timely [23]. times [82, 1190, 1733, 969]. timing [741]. Timothy [1483]. Titan [1148]. Titanic [738, 1209, 1201]. Toby [311]. TOC [1400, 1423, 1444, 1545, 1567, 1584, 1605, 1624, 1641, 1658]. today [673]. together [822, 981]. told [1146]. toll [1452, 1088]. tombs [638]. tomorrow [481]. tongues [183]. Too [336, 472, 323, 170, 641, 466, 1004]. took [1721]. tool [484, 281]. Tools [762]. tooth [741, 367]. Tootsie [693]. Top [274]. tortoise [1532]. Tour [625, 1384]. tower [306]. Town [1113]. towns [1114]. toxicity [220]. Tracking [1359, 391]. trade [762]. Trading [192]. Tragedy [994, 830, 857, 1559]. tragic [947]. trail [339]. transform [323]. transforming [1092]. translation [1099]. transmission [1591]. transparency [998]. transparent [1119]. travel [18]. travels [950]. Treasury [248]. tree [1159, 351]. trees [1721, 712, 306, 1427]. trial [730, 75, 494, 274, 1233, 313, 609, 21, 506]. Trials [1038, 855, 329, 359, 78, 1305, 516, 790]. Triangle [49]. tribulations [1305]. tribute [1346]. trick [669]. Trigger [1116]. trillion [1427]. trimmings [735]. triple [478]. trouble [230]. troubles [1144]. troubling [1249]. trucks [1128]. true [1452, 703]. Trump [1001, 1035, 1108]. Trumpists [1615]. Trust [1106, 1344, 57, 415]. Truth [432, 247, 1665, 1064, 1098]. try [478, 870, 1366]. Trying [490]. tuberculosis [1681]. Tukey [1247]. Tunbridge [245]. Tunnel [513]. turned [997]. Turning [1151, 1210, 1250]. turnout [1129]. TV [277, 1736]. twist [913]. Two [1252, 1331, 485, 556, 1695, 1362, 585, 1372, 1604, 1326, 395, 1105, 1133]. two-piece [1326]. two-sided [1372, 1604]. Two-speed [485]. types [1137]. Tzioli [1504].

UEFA [650]. Ugandan [1681]. ugly [1665]. UK [1106, 279, 414, 816, 1215, 25, 149, 815, 333, 451, 1041, 1591, 1042, 51, 1015, 886, 243, 1646]. Ukraine [1747]. Unanswered [1619]. uncertain [689, 1307]. Uncertainties [689, 398]. uncertainty [1376, 133, 1636, 595, 1112, 1429, 1111, 1023, 689, 1306, 1222]. Uncovering [95]. Underdispersion [1590]. underrepresented [884]. understand [1700]. Understanding [1494, 740, 709, 133, 1231, 681, 66, 803, 726]. unemployment [23]. unexpected [329, 885, 168]. unfinished [732]. ungreen [482]. unimagined [1696]. unintended [1294]. unit [641]. United [570, 434, 1248]. universe [777]. University [1236, 1626, 500]. unknown [413, 50]. unlikely [540]. unnecessarily [789]. unpredictability [356]. Unravelling [1132]. unreasonable [436]. unreliably [275]. unsustainable [452]. until [162]. upon [199]. Upright [1137]. ups [1018]. Urban [481, 483, 485]. urbanisation [485]. USA [1495, 1286, 1170, 322]. Usain [547]. use [1480, 160, 1246, 1244]. used [1078, 1022]. Using [1454, 484, 1534, 1703, 854, 67, 1612, 470]. utan [843].

v [506]. Vaccine [196]. valid [450]. validation [1310]. validity [1230]. value [355, 200, 1000, 1619, 1007, 1471, 1106]. values [821, 4]. variables [49]. Varian [471]. variance [714, 1150, 1288]. variation [963, 1005]. Variations [570]. vary [1074]. Vaughan [1544]. Vegas [142]. veritas [111]. versus [1090, 287, 1647, 1185, 1342]. very [275, 254, 1732, 833, 565]. victimisation [1751]. victims [816]. Victorian [732, 1349]. Victorians [358]. Video [941]. view [1558, 325, 1680, 248, 1458, 668]. Viking [696]. village [433]. villagers [429]. Vince [1172]. vino [111]. violence [1634, 600, 941, 921]. violent [1004]. Violet [1653]. virginica [1555]. virtuosos [552]. virus [905]. visible [1618]. vision [228]. visitation [602]. Visual [1370]. visualisation [1349]. visualisations [1587]. Visualising [690, 1209, 19]. Visuals [1759]. vital [745, 1275]. viz [1686]. volcano [431]. Volkswagen [853]. vote [1465, 966, 883, 597]. voter [1129]. voters [1245]. voting [923, 1695, 435]. vulnerable [1453].

wages [343]. waists [243]. waiting [344]. wake [1753]. wake-up [1753]. Wales [576, 978, 1234]. walk [1216]. Walking [1062]. Wall [533]. Want [1194, 1310, 202, 1599, 789]. wants [1086]. War [1609, 895, 381, 1627, 1747, 1020, 803, 1717, 1022, 1275]. warming [1093]. warning [1033, 1116]. Warren [843]. Was [1001, 1540, 276, 683, 457, 985, 997, 1529, 738]. Washington [1203]. wasps [68]. waste [715, 1153]. wastewater [484]. watch [949, 116, 755]. Water [1260, 1002]. waves [888, 1454]. way [1660, 235, 1110, 1162]. ways [1092]. wealth [1495]. weapons [576]. weather [37, 1532, 872, 1070]. web [879, 765, 785, 798, 812, 827, 840, 850, 865, 917, 937, 956, 974, 993, 1012, 1030, 1048, 1083, 1159, 1178, 1102, 1122, 1141, 1207, 1066, 1337]. wedding [679]. wedge [400]. week [945]. weekly [343]. Weibull [1109]. weigh [760]. Weighing [406, 1020]. weight [722, 494, 1447]. weightlifting [1472]. weird [134]. Weisberg [1394]. welcome [1490, 1531]. well [162, 1612, 1447, 1447]. well-being [1612]. wellbeing [305]. Wells [245]. Wen [1561, 1560, 1562, 1559]. Wen-chen [1561, 1560, 1562, 1559]. Were [924, 1418, 555, 1130, 985, 497]. West [727]. Western [789]. Whale [160]. wheat [343]. Where [1130, 1075, 957, 276, 1646, 1417, 1304, 713, 538, 478]. Which [1684, 862, 365, 416, 759]. whirl [701]. whisky [338]. white [181]. whiter [1287]. Who [57, 397, 1540, 145, 1146, 1747, 1057, 198, 1319, 1739, 1071, 1164, 1022, 211, 745, 1304, 700, 729]. whom [457]. wide [484, 93]. WiFi [675]. Wild [152]. Wilde [1216]. Wiley [332, 1158, 1177, 1199, 1220, 1238, 1259, 1299]. Will [1293, 513, 1548, 1700, 862, 1364, 323, 747, 636, 1204, 597]. Williams [1544]. willing [1418]. Wills [601]. win [600, 659, 554]. wind [212, 654, 1208]. wine [1095]. winner [883]. winners [260, 1716, 546]. Winning [1515, 1747, 1061]. wins [145]. Winter [1067, 1033]. withdrawal [943]. within [581, 78]. Without [643, 1524, 761]. woman [1078]. Women [1629, 1613, 1646, 884, 1253, 1649, 549, 844, 487, 1245, 1353, 1472, 1348]. women’s [549]. wonderful [1211]. Wood [1060]. woods [254]. Woodward [539]. Woolton [1022]. WORD [1269]. Wordlers [1609]. words [258]. work [932, 287, 1204, 447, 1476]. Working [820, 1128]. works [21]. World [1735, 1275, 1524, 747, 270, 722, 1425, 744, 99, 761, 1226, 689, 778, 1476, 1162, 1418, 1327, 633, 600, 192, 1479]. worlds [682, 1342]. worms [371]. worse [1615]. worthwhile [806]. would [1061, 886]. wraps [1256]. wretched [31]. write [1435, 720]. Writers [679]. Writing [720, 628, 525, 980, 1689]. wrong [845, 444, 1071, 982, 759]. wrongs [1188].

xkcd [661]. xvi [1479].

Yardley [342]. Yau [631]. Year [627, 628, 626, 738, 1451, 1007, 1301, 1320, 1260]. years [1504, 40, 947, 613, 660, 457, 1131, 1471, 1133]. Yellow [1369]. Yemen [1212]. Yende [1705]. Yende-Zuma [1705]. yolk [925]. York [756, 1039]. young [130, 469, 390, 254, 1246, 21, 720, 628, 969, 679]. yours [847]. You’ve [29, 759].

Zachary [1164]. Zero [1113]. Zika [905]. Zimbabwe [1167]. Zipf [700]. Zombie [1128]. zone [892]. zoo [1412]. Zuma [1705].

References

Anonymous:2004:Na

[1] Anonymous. News. Significance (Oxford, England), 1(1):2–5, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gilks:2004:BNS

[2] Wally Gilks. Bioinformatics: new science–new statistics? Significance (Oxford, England), 1(1):7–9, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spiegelhalter:2004:SSL

[3] David Spiegelhalter and Nicky Best. Shipman’s statistical legacy. Significance (Oxford, England), 1(1):10–12, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Heffernan:2004:EVD

[4] Janet Heffernan and Jonathan Tawn. Extreme values in the dock. Significance (Oxford, England), 1(1):13–17, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Briscoe:2004:HPI

[5] Simon Briscoe. House price indices. Significance (Oxford, England), 1 (1):18–20, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dale:2004:SN

[6] Angela Dale. Safety in numbers. Significance (Oxford, England), 1(1): 21–25, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fisher:2004:DFCa

[7] Fisher. Dr Fisher’s casebook. Significance (Oxford, England), 1(1):26, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cox:2004:AS

[8] David Cox. The Accidental Statistician. Significance (Oxford, England), 1(1):27–29, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Neil:2004:ADQ

[9] Stuart Neil. Agricultural data quality. Significance (Oxford, England), 1 (1):30–32, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Grieve:2004:CSP

[10] Andy Grieve. Career story: Pharmaceutical Statistician. Significance (Oxford, England), 1(1):33–35, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Caulcutt:2004:MF

[11] Roland Caulcutt. Managing by fact. Significance (Oxford, England), 1 (1):36–38, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2004:DSM

[12] R. Allan Reese. Does significance matter? Significance (Oxford, England), 1(1):39–40, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:A

[13] Anonymous. From the archives. Significance (Oxford, England), 1(1): 41, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:La

[14] Anonymous. Letters. Significance (Oxford, England), 1(1):42–43, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Pa

[15] Anonymous. Puzzle. Significance (Oxford, England), 1(1):43, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Ea

[16] Anonymous. Editorial. Significance (Oxford, England), 1(1):42, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Ra

[17] Anonymous. Reviews. Significance (Oxford, England), 1(1):44–45, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lievesley:2004:TT

[18] Denise Lievesley. Talk and travel. Significance (Oxford, England), 1(1): 46–47, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Woodward:2004:VIE

[19] Phil Woodward and Mark Whitlock. Visualising interaction effects. Significance (Oxford, England), 1(1):48, March 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Nb

[20] Anonymous. News. Significance (Oxford, England), 1(2):50–53, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Sheila:2004:CJT

[21] Sheila, M. Bird, Clive, and B. Fairweather. Criminal justice on trial: What works for young offenders. Significance (Oxford, England), 1(2): 55–58, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fisher:2004:DFCb

[22] Fisher. Dr Fisher’s casebook. Significance (Oxford, England), 1(2):59, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Thomas:2004:TMU

[23] Ray Thomas. A timely measure of unemployment. Significance (Oxford, England), 1(2):60–63, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Longva:2004:IDD

[24] Svein Longva. Indicators for democratic debate: the role of statistics in Norwegian elections. Significance (Oxford, England), 1(2):64–66, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cook:2004:OSU

[25] Len Cook. Official statistics and UK elections: a response from the UK’s National Statistician. Significance (Oxford, England), 1(2):67–68, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Goodman:2004:FSS

[26] Joshua Goodman and David Heckerman. Fighting spam with statistics. Significance (Oxford, England), 1(2):69–72, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lindley:2004:BT

[27] Dennis Lindley. Bayesian thoughts. Significance (Oxford, England), 1 (2):73–75, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mohamed:2004:CSF

[28] Omar Mohamed. Career story: Financial Services Statistician. Significance (Oxford, England), 1(2):76–77, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mohamed:2004:YGD

[29] Omar Mohamed. You’ve got direct mail. Significance (Oxford, England), 1(2):78–80, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Caulcutt:2004:CCP

[30] Roland Caulcutt. Control charts in practice. Significance (Oxford, England), 1(2):81–84, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lindley:2004:WP

[31] Dennis Lindley. That wretched prior. Significance (Oxford, England), 1 (2):85–87, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:CDA

[32] Anonymous. The Contagious Diseases Acts: a controlled experiment in criminal justice. Significance (Oxford, England), 1(2):88–89, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Eb

[33] Anonymous. Editorial. Significance (Oxford, England), 1(2):90, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Lb

[34] Anonymous. Letters. Significance (Oxford, England), 1(2):90–91, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Pb

[35] Anonymous. Puzzle. Significance (Oxford, England), 1(2):91, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Rb

[36] Anonymous. Reviews. Significance (Oxford, England), 1(2):92–93, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Duckworth:2004:FRF

[37] Frank Duckworth. A fair result in foul weather. Significance (Oxford, England), 1(2):94–96, June 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Nc

[38] Anonymous. News. Significance (Oxford, England), 1(3):98–101, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Goldstein:2004:MES

[39] Harvey Goldstein. Measuring educational standards. Significance (Oxford, England), 1(3):103–105, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Barnard:2004:BIS

[40] George Barnard. Business and industrial statistics: the early years. Significance (Oxford, England), 1(3):106–108, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Williams:2004:DMB

[41] Leighton Vaughan Williams. Decision-making in betting markets. Significance (Oxford, England), 1(3):109–112, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mills:2004:ERB

[42] Terence Mills. Exploring the relationship between gold and the dollar. Significance (Oxford, England), 1(3):113–115, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Atkinson:2004:FG

[43] Richard Atkinson and John Haigh. A fairground game. Significance (Oxford, England), 1(3):116–117, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Clough:2004:ISH

[44] Helen Clough and Patrick Brown. It shouldn’t happen to a statistician. Significance (Oxford, England), 1(3):118–120, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:DFCa

[45] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 1(3): 121, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Pomery:2004:PC

[46] Michael Pomery. The pensions crisis. Significance (Oxford, England), 1 (3):122–124, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rice:2004:CSO

[47] Melissa Rice. Career story: Official Statistician. Significance (Oxford, England), 1(3):125–126, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gilmour:2004:IFP

[48] Steven Gilmour. Improving a food process by experimentation. Significance (Oxford, England), 1(3):127–129, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Alexander:2004:TPT

[49] Neal Alexander. Triangle plots of three-level explanatory variables. Significance (Oxford, England), 1(3):130–131, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

OHagan:2004:DU

[50] Tony O’Hagan. Dicing with the unknown. Significance (Oxford, England), 1(3):132–133, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spiegelhalter:2004:BU

[51] David Spiegelhalter. Bastardy in the UK. Significance (Oxford, England), 1(3):134–135, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Ec

[52] Anonymous. Editorial. Significance (Oxford, England), 1(3):136, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Lc

[53] Anonymous. Letters. Significance (Oxford, England), 1(3):136–138, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:SPP

[54] Anonymous. Springer prize puzzle. Significance (Oxford, England), 1(3): 138, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Rc

[55] Anonymous. Reviews. Significance (Oxford, England), 1(3):139–141, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lauritzen:2004:MC

[56] Steffen Lauritzen. A model company. Significance (Oxford, England), 1(3):142–144, September 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:WDY

[57] Anonymous. Who do you trust? Significance (Oxford, England), 1(4): 146–149, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Marmot:2004:SS

[58] Michael Marmot. Status syndrome. Significance (Oxford, England), 1 (4):150–154, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Challenor:2004:PRC

[59] Peter Challenor. The probability of rapid climate change. Significance (Oxford, England), 1(4):155–158, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Senn:2004:WDD

[60] Stephen Senn. When is a drug not a drug? Significance (Oxford, England), 1(4):159–161, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cook:2004:MYH

[61] Adrian Cook. Mind your head: the data and debate on bicycle helmet effectiveness. Significance (Oxford, England), 1(4):162–163, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Halloran:2004:BSI

[62] Elizabeth Halloran and Ira Longini. Bioterrorism: the statistical issues. Significance (Oxford, England), 1(4):164–168, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:DFCb

[63] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 1(4): 169, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Armitage:2004:PMP

[64] Peter Armitage. Purposes, methods, philosophies. Significance (Oxford, England), 1(4):170–173, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Coleman:2004:CSM

[65] Ruth Coleman. Career story: Medical Statistician. Significance (Oxford, England), 1(4):174–175, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Donnelly:2004:RTE

[66] Christl Donnelly and Azra Ghani. Real-time epidemiology: Understanding the spread of SARS. Significance (Oxford, England), 1(4):176–179, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wood:2004:SIU

[67] Michael Wood. Statistical inference using bootstrap confidence intervals. Significance (Oxford, England), 1(4):180–182, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spiegelhalter:2004:WCD

[68] David Spiegelhalter. On wasps and club dinners. Significance (Oxford, England), 1(4):183–184, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Rd

[69] Anonymous. Reviews. Significance (Oxford, England), 1(4):185–187, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Ed

[70] Anonymous. Editorial. Significance (Oxford, England), 1(4):188, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Ld

[71] Anonymous. Letters. Significance (Oxford, England), 1(4):188–189, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2004:Pc

[72] Anonymous. Puzzle. Significance (Oxford, England), 1(4):189, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jeffers:2004:GC

[73] John Jeffers. A growing concern. Significance (Oxford, England), 1(4): 190–192, December 2004. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:MSS

[74] Anonymous. More scrutiny for scientific statistics. Significance (Oxford, England), 2(1):2–5, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dawid:2005:ST

[75] Philip Dawid. Statistics on trial. Significance (Oxford, England), 2(1): 6–8, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gardner-Medwin:2005:WPS

[76] Tony Gardner-Medwin. What probability should a jury address? Significance (Oxford, England), 2(1):9–12, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hill:2005:RCD

[77] Ray Hill. Reflections on the cot death cases. Significance (Oxford, England), 2(1):13–16, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nobles:2005:MSW

[78] Richard Nobles and David Schiff. Misleading statistics within criminal trials: The Sally Clark case. Significance (Oxford, England), 2(1):17–19, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Balding:2005:QI

[79] David Balding. A question of identity. Significance (Oxford, England), 2 (1):20–23, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Aitken:2005:SIA

[80] Colin Aitken. Sampling issues in the administration of justice. Significance (Oxford, England), 2(1):24–27, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Laycock:2005:DCB

[81] Patrick Laycock. Drug contamination of banknotes. Significance (Oxford, England), 2(1):28–29, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:ITO

[82] Anonymous. Interesting times for official statistics. Significance (Oxford, England), 2(1):30–33, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Evett:2005:CSC

[83] Ian Evett. Career story: Consultant Forensic Statistician. Significance (Oxford, England), 2(1):34–37, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spiegelhalter:2005:PPH

[84] David Spiegelhalter. Public or private health care?: Farr on “the mortality of lunatics”. Significance (Oxford, England), 2(1):38–39, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:Ra

[85] Anonymous. Reviews. Significance (Oxford, England), 2(1):40–43, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:Ea

[86] Anonymous. Editorial. Significance (Oxford, England), 2(1):44, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:La

[87] Anonymous. Letters. Significance (Oxford, England), 2(1):44–45, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:SPCa

[88] Anonymous. Springer Prize Crossword. Significance (Oxford, England), 2(1):45, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Donnelly:2005:AS

[89] Peter Donnelly. Appealing statistics. Significance (Oxford, England), 2 (1):46–48, March 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:GEP

[90] Anonymous. The 2005 General Election polls. Significance (Oxford, England), 2(2):50–53, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McHale:2005:RFP

[91] Ian McHale and Phil Scarf. Ranking football players. Significance (Oxford, England), 2(2):54–57, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mahon:2005:HSG

[92] Basil Mahon. How science got statistics. Significance (Oxford, England), 2(2):58–61, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Marchant:2005:EAW

[93] Paul Marchant. Evaluating area-wide crime-reduction measures. Significance (Oxford, England), 2(2):62–65, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Whelan:2005:RRP

[94] Shane Whelan. RPI: a Really Profligate Index? Significance (Oxford, England), 2(2):66–68, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bird:2005:FBU

[95] Sheila Bird. Filling in the blanks: Uncovering confidential information regarding Scotland’s prisoner escort. Significance (Oxford, England), 2 (2):69–73, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:DFCa

[96] Anonymous. Dr Fisher’s casebook: In praise of simple methods. Significance (Oxford, England), 2(2):74, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nelder:2005:SQ

[97] John Nelder. Synthesis and quality. Significance (Oxford, England), 2 (2):75–77, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Langham:2005:AMR

[98] Alison Langham. Asylum and migration: a review of Home Office statistics. Significance (Oxford, England), 2(2):78–80, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hand:2005:SMH

[99] David J. Hand. Size matters — how measurement defines our world. Significance (Oxford, England), 2(2):81–83, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

OHagan:2005:E

[100] Tony O’Hagan. Elicitation. Significance (Oxford, England), 2(2):84–86, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Venn:2005:DS

[101] John Venn. Defying symmetry. Significance (Oxford, England), 2(2): 87–88, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:Rb

[102] Anonymous. Reviews. Significance (Oxford, England), 2(2):89–91, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:EPD

[103] Anonymous. Editorial: a plea for data. Significance (Oxford, England), 2(2):92, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:Lb

[104] Anonymous. Letters. Significance (Oxford, England), 2(2):92–93, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:SPCb

[105] Anonymous. Springer Prize Crossword. Significance (Oxford, England), 2(2):93, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Blastland:2005:LDM

[106] Michael Blastland. ‘Dear More or Less… ’. Significance (Oxford, England), 2(2):94–96, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:Na

[107] Anonymous. News. Significance (Oxford, England), 2(3):98–101, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gibbens:2005:DGM

[108] Richard Gibbens and Wiebke Werft. Data gold mining: MIDAS and journey time predictors. Significance (Oxford, England), 2(3):102–105, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bland:2005:HHD

[109] Martin Bland. The Horizon homeopathic dilution experiment. Significance (Oxford, England), 2(3):106–109, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hand:2005:ROD

[110] David J. Hand. A rich opportunity: Data analysis in personal financial services. Significance (Oxford, England), 2(3):110–113, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hin:2005:VV

[111] Lin Y. Hin. In vino veritas. Significance (Oxford, England), 2(3):114– 115, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Campbell:2005:LDD

[112] Mike Campbell. Life and death decisions: a statistician on a data monitoring and ethics committee. Significance (Oxford, England), 2(3):116– 118, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Freeman:2005:LSD

[113] Geoff Freeman. Latin squares and su doku. Significance (Oxford, England), 2(3):119–122, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:DFCb

[114] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 2 (3):123–124, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Whelan:2005:CSA

[115] Shane Whelan. Career story: Actuarial researcher. Significance (Oxford, England), 2(3):125–127, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Garcia:2005:NWP

[116] Paul Garcia. Neighbourhood watch: Population forecasting for new housing developments. Significance (Oxford, England), 2(3):128–130, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Chatfield:2005:TSF

[117] Chris Chatfield. Time-series forecasting. Significance (Oxford, England), 2(3):131–133, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2005:B

[118] R. Allan Reese. Boxplots. Significance (Oxford, England), 2(3):134– 135, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gilchrist:2005:GM

[119] Warren Gilchrist. Galton misrepresented. Significance (Oxford, England), 2(3):136–137, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:Eb

[120] Anonymous. Editorial. Significance (Oxford, England), 2(3):138, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:Lc

[121] Anonymous. Letters. Significance (Oxford, England), 2(3):138–139, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:SPCc

[122] Anonymous. Springer Prize Crossword — The President’s Eleven. Significance (Oxford, England), 2(3):139, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:Rc

[123] Anonymous. Reviews. Significance (Oxford, England), 2(3):140–142, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Eastaway:2005:HTP

[124] Rob Eastaway. How to take a penalty. Significance (Oxford, England), 2(3):143–144, September 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:Nb

[125] Anonymous. News. Significance (Oxford, England), 2(4):146–148, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spiegelhalter:2005:MLS

[126] David Spiegelhalter. The mystery of the lost star: a statistical detective story. Significance (Oxford, England), 2(4):150–153, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Copas:2005:DP

[127] John Copas. The downside of publication. Significance (Oxford, England), 2(4):154–157, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Baines:2005:POE

[128] Paul Baines. Polling opinions in the 2005 election. Significance (Oxford, England), 2(4):158–161, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mardia:2005:MSN

[129] Kanti Mardia and Walter Gilks. Meeting the statistical needs of 21stcentury science. Significance (Oxford, England), 2(4):162–165, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Altman:2005:DLH

[130] Doug Altman and Martin Bland. Do the left-handed die young? Significance (Oxford, England), 2(4):166–170, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Finney:2005:EDS

[131] David Finney. Experimental design and systematic monitoring. Significance (Oxford, England), 2(4):171–173, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dunnell:2005:CSI

[132] Karen Dunnell. Career story: Inside National Statistics. Significance (Oxford, England), 2(4):174–176, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Best:2005:SUG

[133] Nicky Best and Peter Green. Structure and uncertainty: Graphical models for understanding complex data. Significance (Oxford, England), 2 (4):177–181, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Matthews:2005:WDP

[134] Robert Matthews. Why do people believe weird things? Significance (Oxford, England), 2(4):182–184, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Edwards:2005:SS

[135] A. W. F. Edwards. Sexes and statistics. Significance (Oxford, England), 2(4):185–186, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Joyce:2005:E

[136] Helen Joyce and Julian Champkin. Editorial. Significance (Oxford, England), 2(4):187, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:Ld

[137] Anonymous. Letters. Significance (Oxford, England), 2(4):187–188, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:SPP

[138] Anonymous. Springer prize puzzle. Significance (Oxford, England), 2(4): 188, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2005:Rd

[139] Anonymous. Reviews. Significance (Oxford, England), 2(4):189–190, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Eyre:2005:PES

[140] David Eyre and Sheila Bird. Prisoner escort in Scotland: Costs, projection and performance under scrutiny. Significance (Oxford, England), 2(4):191–192, December 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:Na

[141] Anonymous. News. Significance (Oxford, England), 3(1):2–4, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Eadington:2006:VS

[142] William Eadington. Vegas stripped. Significance (Oxford, England), 3 (1):5–9, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Williams:2006:BFO

[143] Professor Leighton Vaughan Williams. Betting the future online. Significance (Oxford, England), 3(1):10–12, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Griffiths:2006:LG

[144] Dr Mark Griffiths. The lost gamblers. Significance (Oxford, England), 3 (1):13–15, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Goddard:2006:WWF

[145] John Goddard. Who wins the football? Significance (Oxford, England), 3(1):16–19, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Forrest:2006:RAA

[146] Dr David Forrest and Dr Robert Simmons. Read all about it? Significance (Oxford, England), 3(1):20–21, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Johnson:2006:SC

[147] Dr Oliver Johnson. Stakes and chips. Significance (Oxford, England), 3 (1):22–25, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Buik:2006:FSB

[148] David Buik. Financial spread betting. Significance (Oxford, England), 3 (1):26–27, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Haigh:2006:RUN

[149] Dr. John Haigh. Reflections on the UK National Lottery. Significance (Oxford, England), 3(1):28–29, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Woodward:2006:CMB

[150] Phil Woodward. Call my bluff. Significance (Oxford, England), 3(1): 30–32, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Shirley:2006:CSI

[151] Stephanie Shirley. Career story: the importance of being ERNIE. Significance (Oxford, England), 3(1):33–36, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Diaconis:2006:WC

[152] Persi Diaconis. Wild card. Significance (Oxford, England), 3(1):37–40, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Docherty:2006:DD

[153] Jenny Docherty. Dettori Day. Significance (Oxford, England), 3(1): 41–42, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:Ra

[154] Anonymous. Reviews. Significance (Oxford, England), 3(1):43, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:E

[155] Anonymous. Editorial. Significance (Oxford, England), 3(1):44, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:La

[156] Anonymous. Letters. Significance (Oxford, England), 3(1):44–45, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:C

[157] Anonymous. Crossword. Significance (Oxford, England), 3(1):45, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Croucher:2006:TSY

[158] John S. Croucher. Teach statistics? You bet! Significance (Oxford, England), 3(1):46–48, March 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:Nb

[159] Anonymous. News. Significance (Oxford, England), 3(2):50–53, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hammond:2006:WSH

[160] Philip Hammond. Whale science — and how (not) to use it. Significance (Oxford, England), 3(2):54–58, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wilson:2006:SRN

[161] Granville Tunnicliffe Wilson. Statistics and risk in the nuclear industry. Significance (Oxford, England), 3(2):59–62, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hackshaw:2006:SGY

[162] Allan Hackshaw. Smoke gets in your eyes — well, only until mid-2007. Significance (Oxford, England), 3(2):63–65, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harbron:2006:SMF

[163] Chris Harbron. Statistics and the medicine of the future. Significance (Oxford, England), 3(2):66–68, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Robertson:2006:PLS

[164] Chris Robertson. Protecting the leaders — syndromic surveillance for the G8 summit in Scotland. Significance (Oxford, England), 3(2):69–72, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Barahono:2006:MEM

[165] Carlos Barahono and Ian Wilson. Monitoring and evaluation in Malawi. Significance (Oxford, England), 3(2):73–75, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:DFCa

[166] Anonymous. Dr Fisher’s casebook: Talking statistics to schools. Significance (Oxford, England), 3(2):76, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Davis:2006:DF

[167] Tim Davis and Julian Champkin. Driving out failure. Significance (Oxford, England), 3(2):77–80, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lawrance:2006:MSF

[168] Tony Lawrance. 15 minutes of statistical fame: Unexpected statistics in the House of Commons and in the media. Significance (Oxford, England), 3(2):81–84, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2006:LP

[169] R. Allan Reese. Line plots. Significance (Oxford, England), 3(2):85–87, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Chetwynd:2006:CST

[170] Amanda Chetwynd and Hilary Short. Cholera — a step too Farr. Significance (Oxford, England), 3(2):88–90, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:Rb

[171] Anonymous. Reviews. Significance (Oxford, England), 3(2):91, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:Lb

[172] Anonymous. Letters. Significance (Oxford, England), 3(2):92–93, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2006:Ea

[173] Julian Champkin. Editorial. Significance (Oxford, England), 3(2):92, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:SPCa

[174] Anonymous. Springer Prize Crossword. Significance (Oxford, England), 3(2):93, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Davies:2006:RDR

[175] Neville Davies. Real data, real learning and the London Olympics. Significance (Oxford, England), 3(2):94–96, June 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:Nc

[176] Anonymous. News. Significance (Oxford, England), 3(3):98–101, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Clough:2006:CTC

[177] Kirsty Clough. Crunch time on carbon: Getting it right in EU ETS phase II. Significance (Oxford, England), 3(3):102–105, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gregory:2006:BBI

[178] Dr Richard Gregory. Birds as biodiversity indicators for Europe. Significance (Oxford, England), 3(3):106–110, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mountain:2006:SCS

[179] Dr Linda Mountain. Safety cameras: stealth tax or life-savers? Significance (Oxford, England), 3(3):111–113, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Davies:2006:CMI

[180] Simon Davies. A case of mistaken identity. Significance (Oxford, England), 3(3):114–117, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ellison:2006:MBW

[181] George Ellison. Medicine in black and white: BiDil(R): race and the limits of evidence-based medicine. Significance (Oxford, England), 3(3): 118–121, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:DFCb

[182] Anonymous. Dr Fisher’s Casebook: a shy nurse consults the good doctor. Significance (Oxford, England), 3(3):122, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Senn:2006:STB

[183] Stephen Senn. Sharp tongues and bitter pills. Significance (Oxford, England), 3(3):123–125, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:BS

[184] Anonymous. Beer and statistics. Significance (Oxford, England), 3(3): 126–129, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Griffiths:2006:SBM

[185] Thomas L. Griffiths. Statistics and the Bayesian mind. Significance (Oxford, England), 3(3):130–133, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fanshawe:2006:DCD

[186] Tom Fanshawe. Dreadful crime! Deer-keeper feloniously resisted! Significance (Oxford, England), 3(3):134–135, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2006:S

[187] R. Allan Reese. Scatterplots. Significance (Oxford, England), 3(3):136– 138, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:Rc

[188] Anonymous. Reviews. Significance (Oxford, England), 3(3):139, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:Lc

[189] Anonymous. Letters. Significance (Oxford, England), 3(3):140, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2006:Eb

[190] Julian Champkin. Editorial. Significance (Oxford, England), 3(3):140, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:SPCb

[191] Anonymous. Springer prize crossword. Significance (Oxford, England), 3(3):141, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hayes:2006:TCC

[192] Kevin Hayes and Ailish Hannigan. Trading coupons: completing the World Cup football sticker album. Significance (Oxford, England), 3 (3):142–144, September 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:Nd

[193] Anonymous. News. Significance (Oxford, England), 3(4):146–149, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Edward:2006:NHS

[194] Nigel Edward. The National Health Service — Can it Survive? Significance (Oxford, England), 3(4):150–152, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Abakuks:2006:SPS

[195] Andris Abakuks. The synoptic problem and statistics. Significance (Oxford, England), 3(4):153–157, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Andrews:2006:VS

[196] Nick Andrews. Vaccine scares. Significance (Oxford, England), 3(4):158– 161, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Farid:2006:DDH

[197] Dr Hany Farid. Digital doctoring: how to tell the real from the fake. Significance (Oxford, England), 3(4):162–166, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bird:2006:TDW

[198] Sheila Bird. Tagging and deportation: who gets it right? Significance (Oxford, England), 3(4):167–170, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:DFCc

[199] Anonymous. Dr Fisher’s casebook: the doctor attends a conference and ponders upon their purpose. Significance (Oxford, England), 3(4):171, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Chalmers:2006:BRW

[200] Iain Chalmers. Biomedical research: Are we getting value for money? Significance (Oxford, England), 3(4):172–175, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Morrison:2006:PD

[201] S. J. Morrison. Publish and be delayed. Significance (Oxford, England), 3(4):176–178, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Simkin:2006:DYS

[202] Mikhail Simkin and Vwany Roychowdhury. Do you sincerely want to be cited? Or: read before you cite. Significance (Oxford, England), 3 (4):179–181, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Edwards:2006:SSG

[203] Professor A. W. F. Edwards. Statisticians in stained glass houses. Significance (Oxford, England), 3(4):182–183, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

OHagan:2006:BF

[204] Tony O’Hagan. Bayes factors. Significance (Oxford, England), 3(4):184– 186, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:Rd

[205] Anonymous. Reviews. Significance (Oxford, England), 3(4):187–188, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2006:Ec

[206] Julian Champkin. Editorial. Significance (Oxford, England), 3(4):189, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:Ld

[207] Anonymous. Letters. Significance (Oxford, England), 3(4):189–190, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2006:SPCc

[208] Anonymous. Springer Prize Crossword. Significance (Oxford, England), 3(4):190, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wolstenholme:2006:DND

[209] Linda Wolstenholme and John Haigh. Deal or no deal? Significance (Oxford, England), 3(4):191–192, December 2006. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:Na

[210] Anonymous. News. Significance (Oxford, England), 4(1):2–5, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cole:2007:BBB

[211] Tim Cole. Babies, bottles, breasts: is the WHO growth standard relevant? Significance (Oxford, England), 4(1):6–10, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Genton:2007:BW

[212] Marc Genton and Amanda Hering. Blowing in the wind. Significance (Oxford, England), 4(1):11–14, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Evans:2007:RSR

[213] Andrew Evans. Rail safety and rail privatization. Significance (Oxford, England), 4(1):15–18, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ke:2007:ESE

[214] Weiming Ke and Brandon T. Bales. Estimation of sampling effort for catching enough bats. Significance (Oxford, England), 4(1):19–21, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hand:2007:DDD

[215] David Hand. Deception and dishonesty with data: Fraud in science. Significance (Oxford, England), 4(1):22–25, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:DFCa

[216] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 4 (1):26, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Moser:2007:LM

[217] Claus Moser and Julian Champkin. Lord Moser. Significance (Oxford, England), 4(1):27–30, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Perry:2007:CSC

[218] Matthew Perry and Julian Champkin. Career story: Climatological Statistician. Significance (Oxford, England), 4(1):31–32, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Chatfield:2007:JC

[219] Chris Chatfield. The joys of consulting. Significance (Oxford, England), 4 (1):33–36, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Festing:2007:EDG

[220] Michael Festing. Experimental design, genetics and animal toxicity tests. Significance (Oxford, England), 4(1):37–40, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2007:BBC

[221] Allan Reese. Bay! Bar charts. Significance (Oxford, England), 4(1): 41–44, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2007:Ea

[222] Julian Champkin. Editorial. Significance (Oxford, England), 4(1):45–46, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Edwards:2007:CT

[223] Anthony Edwards. A cautionary tale. Significance (Oxford, England), 4 (1):47–48, March 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:Nb

[224] Anonymous. News. Significance (Oxford, England), 4(2):50–53, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Zeger:2007:EED

[225] Scott L. Zeger. Estimating excess deaths in Iraq since the US–British-led invasion. Significance (Oxford, England), 4(2):54–59, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Challenor:2007:PRC

[226] Peter Challenor. The probability of rapid climate change: II. Significance (Oxford, England), 4(2):60–62, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Goddard:2007:RBC

[227] John Goddard. Is the ref blind? Crime and punishment in English Premier League football. Significance (Oxford, England), 4(2):63–66, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Southall:2007:VBT

[228] Humphrey Southall. A vision of Britain through time: online access to “statistical heritage”. Significance (Oxford, England), 4(2):67–70, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tapp:2007:HHD

[229] Chris Tapp. Head over heels in debt? Significance (Oxford, England), 4 (2):71–73, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:DFCb

[230] Anonymous. Dr Fisher’s casebook The trouble with data. Significance (Oxford, England), 4(2):74, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Holt:2007:TH

[231] Tim Holt and Julian Champkin. Tim Holt. Significance (Oxford, England), 4(2):75–76, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kemp:2007:OPS

[232] Bess Kemp. Opinion Poll Statistician. Significance (Oxford, England), 4 (2):77–78, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Senn:2007:SF

[233] Stephen Senn. Safety first? Significance (Oxford, England), 4(2):79–80, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kumar:2007:DDD

[234] Kuldeep Kumar and Sukanto Bhattacharya. Detecting the dubious digits: Benford’s law in forensic accounting. Significance (Oxford, England), 4(2):81–83, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Goldacre:2007:WFG

[235] Ben Goldacre. When the facts get in the way of a story. Significance (Oxford, England), 4(2):84–85, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:SBL

[236] Anonymous. Seat belt laws — Repeal them? Significance (Oxford, England), 4(2):86–89, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:Ra

[237] Anonymous. Reviews. Significance (Oxford, England), 4(2):90, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2007:Eb

[238] Julian Champkin. Editorial. Significance (Oxford, England), 4(2):91–92, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Simkin:2007:MSC

[239] Mikhail Simkin. My statistician could have painted that! A statistical inquiry into modern art. Significance (Oxford, England), 4(2):93–96, June 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:Nc

[240] Anonymous. News. Significance (Oxford, England), 4(3):98–101, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Desvaux:2007:SHP

[241] Martin Desvaux. The sustainability of human populations: How many people can live on Earth? Significance (Oxford, England), 4(3):102– 107, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mortimer:2007:FFF

[242] David Mortimer. Fishy food facts: how much fish should we eat? Significance (Oxford, England), 4(3):108–112, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Treleaven:2007:HFY

[243] Philip Treleaven. How to fit into your clothes: Busts, waists, hips and the UK National Sizing Survey. Significance (Oxford, England), 4(3):113– 117, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Burn:2007:EI

[244] Bob Burn. Elephants and ivory. Significance (Oxford, England), 4(3): 118–122, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:DFCc

[245] Anonymous. Dr Fisher’s casebook: Disgusted of Tunbridge Wells. Significance (Oxford, England), 4(3):123, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Borja:2007:BP

[246] Mario Cortina Borja and John Haigh. The birthday problem. Significance (Oxford, England), 4(3):124–127, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Parris:2007:TAL

[247] Matthew Parris. The truth about those little red lights: a tale of power and poppycock. Significance (Oxford, England), 4(3):128–129, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Eagle:2007:SBT

[248] Angela Eagle. The Statistics Bill: the Treasury’s view: the goal of continuous improvement. Significance (Oxford, England), 4(3):130–132, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Goldstein:2007:BFM

[249] Harvey Goldstein. Becoming familiar with multilevel modeling. Significance (Oxford, England), 4(3):133–135, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2007:HMS

[250] Julian Champkin. Henry Mayhew: the statistical Dickens. Significance (Oxford, England), 4(3):136–138, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:Rb

[251] Anonymous. Reviews. Significance (Oxford, England), 4(3):139, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:LE

[252] Anonymous. Letters/ editorial. Significance (Oxford, England), 4(3): 140–141, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:SPCa

[253] Anonymous. Springer Prize Crossword. Significance (Oxford, England), 4(3):141, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

DuFeu:2007:SWS

[254] Chris Du Feu. Seeing the woods — statistics for the very young. Significance (Oxford, England), 4(3):142–144, September 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:Nd

[255] Anonymous. News. Significance (Oxford, England), 4(4):146–149, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Thomas:2007:HMM

[256] Bethan Thomas and Danny Dorling. Human mosaic: maps of honeycombed British society. Significance (Oxford, England), 4(4):150–154, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hollis:2007:HMP

[257] John Hollis. How many people live in London? Significance (Oxford, England), 4(4):155–158, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McMurray:2007:MCM

[258] Bob McMurray. Moo-cow! Mummy! More! How do children learn so many words? Significance (Oxford, England), 4(4):159–163, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Donnelly:2007:BTC

[259] Christl Donnelly and David Cox. Bovine TB in cattle and badgers. Significance (Oxford, England), 4(4):164–167, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Pardoe:2007:POW

[260] Iain Pardoe. Predicting Oscar winners. Significance (Oxford, England), 4(4):168–173, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:DFCd

[261] Anonymous. Dr Fisher’s casebook: RAE (really annoying exercise). Significance (Oxford, England), 4(4):174, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bland:2007:RBM

[262] Martin Bland. Risedronate, the BBC and me. Significance (Oxford, England), 4(4):175–178, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Briscoe:2007:HS

[263] Simon Briscoe. Head scholar. Significance (Oxford, England), 4(4):179– 181, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Holt:2007:SRS

[264] Tim Holt. The Statistics and Registration Service Act. Significance (Oxford, England), 4(4):182–183, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2007:DDD

[265] Allan Reese. Databases and documenting data. Significance (Oxford, England), 4(4):184–186, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:Rc

[266] Anonymous. Reviews. Significance (Oxford, England), 4(4):187, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2007:Ec

[267] Julian Champkin. Editorial. Significance (Oxford, England), 4(4):188, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:L

[268] Anonymous. Letters. Significance (Oxford, England), 4(4):188–189, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2007:SPCb

[269] Anonymous. Springer Prize Crossword. Significance (Oxford, England), 4(4):189, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Clarke:2007:TPP

[270] Lee Clarke. Thinking possibilistically in a probabilistic world. Significance (Oxford, England), 4(4):190–192, December 2007. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:Na

[271] Anonymous. News. Significance (Oxford, England), 5(1):2–5, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hannaford:2008:ILB

[272] Phil Hannaford. Investigating the link between the Pill and cancer. Significance (Oxford, England), 5(1):6–10, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hand:2008:PDD

[273] David Hand. Privacy, data discs and realistic risk. Significance (Oxford, England), 5(1):11–14, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Haigh:2008:RLT

[274] John Haigh. Racing, the Law and the Top Jockey: the Kieren Fallon trial. Significance (Oxford, England), 5(1):15–17, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Brooker:2008:AMA

[275] Peter Brooker. ANASE: measuring aircraft noise annoyance very unreliably. Significance (Oxford, England), 5(1):18–24, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bird:2008:SSB

[276] Sheila Bird. The Scottish spoiled ballot: where was statistics? Significance (Oxford, England), 5(1):25–28, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:DFCa

[277] Anonymous. Dr Fisher’s casebook: TV personalities. Significance (Oxford, England), 5(1):29, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spiegelhalter:2008:BSM

[278] David Spiegelhalter and Hauke Riesch. Bacon sandwiches and middle — class drinkers: the risk of communicating risk. Significance (Oxford, England), 5(1):30–33, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Alldritt:2008:NIC

[279] Richard Alldritt and Richard Laux. No islands: charting a course for UK official statistics. Significance (Oxford, England), 5(1):34–36, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Senn:2008:CT

[280] Stephen Senn. A century of t-tests. Significance (Oxford, England), 5(1): 37–39, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Senn:2008:TT

[281] Stephen Senn. The t-test tool. Significance (Oxford, England), 5(1): 40–41, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:Ra

[282] Anonymous. Reviews. Significance (Oxford, England), 5(1):42–43, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:LEa

[283] Anonymous. Letters/ editorial. Significance (Oxford, England), 5(1): 44–45, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:SPCa

[284] Anonymous. Springer Prize Crossword. Significance (Oxford, England), 5(1):45, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jones:2008:DMS

[285] Chris Jones. Don’t mention Stonehenge! Significance (Oxford, England), 5(1):46–48, March 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:Nb

[286] Anonymous. News. Significance (Oxford, England), 5(2):50–53, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Johnson:2008:DAW

[287] Blair T. Johnson and Irving Kirsch. Do antidepressants work? Statistical significance versus clinical benefits. Significance (Oxford, England), 5(2): 54–58, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Edelman:2008:CHB

[288] David Edelman. Credit this: how the banks decide your credit score. Significance (Oxford, England), 5(2):59–61, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cavagna:2008:SS

[289] Andrea Cavagna and Irene Giardina. The seventh starling. Significance (Oxford, England), 5(2):62–66, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Goldstein:2008:SLT

[290] Harvey Goldstein and George Leckie. School league tables: what can they really tell us? Significance (Oxford, England), 5(2):67–69, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ludbrook:2008:AT

[291] Anne Ludbrook. Alcohol and taxation. Significance (Oxford, England), 5(2):70–73, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bowman:2008:SHF

[292] Adrian Bowman. Statistics with a human face. Significance (Oxford, England), 5(2):74–77, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:DFCb

[293] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 5 (2):78, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Briscoe:2008:PNL

[294] Simon Briscoe. Panic now! Or later? Significance (Oxford, England), 5 (2):79–81, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Giles:2008:LDL

[295] Chris Giles. Lies, damned lies and befuddlement. Significance (Oxford, England), 5(2):82–83, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Allsop:2008:SBL

[296] Richard Allsop, Oliver Carsten, Andrew Evans, and Robert Gifford. Seat belt laws: why we should keep them. Significance (Oxford, England), 5 (2):84–86, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2008:SR

[297] Allan Reese. Scatterplots revisited. Significance (Oxford, England), 5 (2):87–89, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:Rb

[298] Anonymous. Reviews. Significance (Oxford, England), 5(2):90–91, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2008:EL

[299] Julian Champkin, Mick Roberts, Ian Flindell, Paul LeMasurier, and Stephen Senn. Editorial/ letters. Significance (Oxford, England), 5(2): 92–93, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:SPCb

[300] Anonymous. Springer prize crossword. Significance (Oxford, England), 5(2):93, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Coleman:2008:SSO

[301] Shirley Coleman. Six Sigma: an opportunity for statistics and for statisticians. Significance (Oxford, England), 5(2):94–96, June 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:Nc

[302] Anonymous. News. Significance (Oxford, England), 5(3):98–99, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Carroll:2008:FPF

[303] Patrick Carroll. Future pensions: future poverty? Significance (Oxford, England), 5(3):100–103, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Butler:2008:SCO

[304] Gillian Butler. Spoilt for choice — is organic milk really any different? Significance (Oxford, England), 5(3):104–107, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Aassve:2008:MDE

[305] Arnstein Aassve, Gianni Betti, Stefano Mazzuco, and Letizia Mencarini. Marital disruption and economic wellbeing. Significance (Oxford, England), 5(3):108–111, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nicholls:2008:HFT

[306] Geoff Nicholls. Horses or farmers? The tower of Babel and confidence in trees. Significance (Oxford, England), 5(3):112–117, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Brooker:2008:ISM

[307] Peter Brooker. Is it safe for me to fly? Significance (Oxford, England), 5(3):118–121, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dorling:2008:LOG

[308] Danny Dorling. Looking from outside the goldfish bowl. Significance (Oxford, England), 5(3):122–125, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:DFCc

[309] Anonymous. Dr Fisher’s casebook: The Doctor minds his p s. Significance (Oxford, England), 5(3):126, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Blastland:2008:SSR

[310] Michael Blastland. The storyteller and the statistician or Recession? Already? Significance (Oxford, England), 5(3):127–129, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2008:TL

[311] Julian Champkin. Toby Lewis. Significance (Oxford, England), 5(3):130– 132, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Joyce:2008:MNM

[312] Helen Joyce. Minard and Napoleon’s march on Moscow. Significance (Oxford, England), 5(3):133–134, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jones:2008:CTS

[313] Byron Jones. The cross-over trial: a subtle knife. Significance (Oxford, England), 5(3):135–137, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:Rc

[314] Anonymous. Reviews. Significance (Oxford, England), 5(3):138, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:LEb

[315] Anonymous. Letters/ editorial. Significance (Oxford, England), 5(3): 139–140, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:SPCc

[316] Anonymous. Springer Prize Crossword. Significance (Oxford, England), 5(3):140, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2008:RNS

[317] Julian Champkin. Running the numbers: statistics as art. Significance (Oxford, England), 5(3):141–144, September 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:Nd

[318] Anonymous. News. Significance (Oxford, England), 5(4):146–148, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Pike:2008:HMO

[319] Richard Pike. How much oil is really there? Making correct statistics bring reality to global planning. Significance (Oxford, England), 5(4): 149–152, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hilbe:2008:ABO

[320] Joseph Hilbe. Athletics at the Beijing Olympics: how much faster can anyone run? Significance (Oxford, England), 5(4):153–158, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Knutti:2008:HSW

[321] Reto Knutti. Hotter or not? Should we believe model predictions of future climate change? Significance (Oxford, England), 5(4):159–162, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

OHagan:2008:HMK

[322] Tony O’Hagan, Bob Ward, and Katie Coughlin. How many Katrinas? Predicting the number of hurricanes striking the USA. Significance (Oxford, England), 5(4):163–167, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Boardman:2008:CLT

[323] Brenda Boardman. Carbon labelling: too complex or will it transform our buying? Significance (Oxford, England), 5(4):168–171, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:DFCd

[324] Anonymous. Dr Fisher’s casebook: The Doctor sees the light. Significance (Oxford, England), 5(4):172, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bingham:2008:CMV

[325] Nick Bingham. The crash of 2008: a mathematician’s view. Significance (Oxford, England), 5(4):173–175, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Macintosh:2008:EDG

[326] Ann Macintosh. The emergence of digital governance. Significance (Oxford, England), 5(4):176–178, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2008:LL

[327] Julian Champkin. Lord Lipsey. Significance (Oxford, England), 5(4): 179–180, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Joyce:2008:FNL

[328] Helen Joyce. Florence Nightingale: a lady with more than a lamp. Significance (Oxford, England), 5(4):181–182, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jones:2008:CTP

[329] Byron Jones and Scott Haughie. Cross-over trials in practice: tales of the unexpected. Significance (Oxford, England), 5(4):183–184, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:Rd

[330] Anonymous. Reviews. Significance (Oxford, England), 5(4):185–186, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:LEc

[331] Anonymous. Letters/ editorial. Significance (Oxford, England), 5(4): 187–188, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2008:WPC

[332] Anonymous. Wiley Prize Crossword. Significance (Oxford, England), 5 (4):188, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McElduff:2008:WNF

[333] Fiona McElduff, Pablo Mateos, Angie Wade, and Mario Cortina Borja. What’s in a name? The frequency and geographic distributions of UK surnames. Significance (Oxford, England), 5(4):189–192, December 2008. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:Na

[334] Anonymous. News. Significance (Oxford, England), 6(1):2–4, March 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spiegelhalter:2009:LMP

[335] David Spiegelhalter and Arthur Barnett. London murders: a predictable pattern? Significance (Oxford, England), 6(1):5–8, March 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hesketh:2009:TMM

[336] Thérèse Hesketh. Too many males in China: the causes and the consequences. Significance (Oxford, England), 6(1):9–13, March 2009. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gwynne:2009:SNH

[337] Steven Gwynne. Safety in numbers: How human beings react to fire. Significance (Oxford, England), 6(1):14–19, March 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wishart:2009:FW

[338] David Wishart. The flavour of whisky. Significance (Oxford, England), 6 (1):20–26, March 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:DFCa

[339] Anonymous. Dr Fisher’s casebook: Don’t spell it “clinical trail”. Significance (Oxford, England), 6(1):27, March 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Janeway:2009:SIT

[340] William Janeway. Six impossible things before breakfast: Lessons from the crisis. Significance (Oxford, England), 6(1):28–31, March 2009. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Goddard:2009:BBS

[341] Ivor Goddard. Big Brother and state control — a new career for statisticians? Significance (Oxford, England), 6(1):32–34, March 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Yardley:2009:NSN

[342] Jane Yardley. A novel in statistics: Numbers and Jane Yardley: why the heroine of her latest novel is a biostatistician. Significance (Oxford, England), 6(1):35–37, March 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Joyce:2009:WWP

[343] Helen Joyce. Of weekly wages and the price of wheat — and soaring IQs. Significance (Oxford, England), 6(1):38–39, March 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jones:2009:WGH

[344] Byron Jones. The waiting game: how long is long enough? Significance (Oxford, England), 6(1):40–41, March 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:Ra

[345] Anonymous. Reviews. Significance (Oxford, England), 6(1):42, March 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:LEa

[346] Anonymous. Letters/ editorial. Significance (Oxford, England), 6(1): 43–44, March 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:Ca

[347] Anonymous. Crossword. Significance (Oxford, England), 6(1):44, March 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hilbe:2009:BEA

[348] Joseph Hilbe. Best ever athlete? Significance (Oxford, England), 6(1): 45–48, March 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:Nb

[349] Anonymous. News. Significance (Oxford, England), 6(2):50–51, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wolstenholme:2009:SCS

[350] Linda Wolstenholme. Sailing close to the statistics. Significance (Oxford, England), 6(2):52–57, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Crow:2009:LLH

[351] Peter Crow. Landscapes and LiDAR: history under the greenwood tree. Significance (Oxford, England), 6(2):58–62, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Green:2009:WPS

[352] Francis Green, Stephen Machin, Richard Murphy, and Yu Zhu. What have private schools done for (some of) us? Significance (Oxford, England), 6(2):63–67, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wheeler:2009:SSS

[353] Frederick Wheeler. The supermarkets and the small shop: is there a real choice? Significance (Oxford, England), 6(2):68–77, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Simpson:2009:EGB

[354] Ludi Simpson and Nissa Finney. Ethnic ghettos in Britain: a fact or a myth? Significance (Oxford, England), 6(2):72–75, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:DFCb

[355] Anonymous. Dr Fisher’s Casebook: Return of the p-value. Significance (Oxford, England), 6(2):76–77, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ramos:2009:ACN

[356] Osvanny Ramos. Avalanche! Coming! Now! Close to the borders of unpredictability. Significance (Oxford, England), 6(2):78–81, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Maclean:2009:SAA

[357] Ian Maclean. Is the Statistics Authority another factory inspectorate? Significance (Oxford, England), 6(2):82–85, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Magnello:2009:EVE

[358] Eileen Magnello. Eminent Victorians and early statistical societies. Significance (Oxford, England), 6(2):86–88, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jones:2009:CCT

[359] Byron Jones and Scott Haughie. Crossing over to cross-over trials. Significance (Oxford, England), 6(2):89–90, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:Rb

[360] Anonymous. Reviews. Significance (Oxford, England), 6(2):91, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:LEb

[361] Anonymous. Letters/ editorial. Significance (Oxford, England), 6(2): 92–93, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:Cb

[362] Anonymous. Crossword. Significance (Oxford, England), 6(2):93, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Sixsmith:2009:MSP

[363] Edmund Sixsmith. The megalithic story of Professor Alexander Thom. Significance (Oxford, England), 6(2):94–96, June 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:Nc

[364] Anonymous. News. Significance (Oxford, England), 6(3):98–99, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Keith:2009:RLD

[365] David Keith. Red listing: deciding which species are at risk. Significance (Oxford, England), 6(3):100–104, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Buckland:2009:QB

[366] Steve Buckland. Quantifying biodiversity. Significance (Oxford, England), 6(3):105–107, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Thomas:2009:GSR

[367] Len Thomas. Grey seals red in tooth and claw: how Darwin helps model their population. Significance (Oxford, England), 6(3):108–112, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Bell:2009:WTR

[368] Sandra Bell. When the time is right: phenology and flowerings. Significance (Oxford, England), 6(3):113–116, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Benton:2009:CFR

[369] Michael Benton. The completeness of the fossil record. Significance (Oxford, England), 6(3):117–121, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Bamber:2009:WHW

[370] Jonathan Bamber. What happens when an ice sheet melts? Significance (Oxford, England), 6(3):122–125, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Anonymous:2009:DFCc

[371] Anonymous. Dr Fisher’s casebook: a diet of worms. Significance (Oxford, England), 6(3):126, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Reese:2009:DME

[372] R. Allan Reese. Darwin, Mendel and the evolution of evolution. Significance (Oxford, England), 6(3):127–129, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Anonymous:2009:JDH

[373] Anonymous. Jim Durbin: How did statisticians evolve? Significance (Oxford, England), 6(3):130–131, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Gillham:2009:CCD

[374] Nicholas W. Gillham. Cousins: Charles Darwin, Sir Francis Galton and the birth of eugenics. Significance (Oxford, England), 6(3):132– 135, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Marques:2009:DSE

[375] Tiago Marques. Distance sampling: estimating animal density. Significance (Oxford, England), 6(3):136–137, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Anonymous:2009:Rc

[376] Anonymous. Reviews. Significance (Oxford, England), 6(3):138–139, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Anonymous:2009:LEc

[377] Anonymous. Letters/ editorial. Significance (Oxford, England), 6(3): 140–141, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Anonymous:2009:Cc

[378] Anonymous. Crossword. Significance (Oxford, England), 6(3):141, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Burchell:2009:LWC

[379] Mark Burchell. Life: what is the chance that we are alone? Significance (Oxford, England), 6(3):142–144, September 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). Darwin 200th Anniversary.

Anonymous:2009:Nd

[380] Anonymous. News. Significance (Oxford, England), 6(4):146–147, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fairweather:2009:ASW

[381] Clive Fairweather. Afghanistan: statistics of war: Helicopters, bombs and strategies. Significance (Oxford, England), 6(4):148–150, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spiegelhalter:2009:OMG

[382] David Spiegelhalter and Yin-Lam Ng. One match to go! Significance (Oxford, England), 6(4):151–153, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cowan:2009:TNL

[383] Glen Cowan. Testing nature to the limit: the Large Hadron Collider. Significance (Oxford, England), 6(4):154–158, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jungers:2009:GHH

[384] William Jungers and Karen Baab. The geometry of hobbits: Homo floresiensis and human evolution. Significance (Oxford, England), 6(4): 159–164, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Montemurro:2009:SMD

[385] Marcelo Montemurro and Damián Zanette. The statistics of meaning: Darwin, Gibbon and Moby Dick. Significance (Oxford, England), 6(4): 165–169, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:DFCd

[386] Anonymous. Dr Fisher’s casebook: Living logarithmically. Significance (Oxford, England), 6(4):170–171, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Blunt:2009:GSL

[387] Gordon Blunt. Green shoots … or shallow roots? Significance (Oxford, England), 6(4):172–175, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Buchanan:2009:STD

[388] Mark Buchanan. The sands of time: does history obey the rules of statistics? Significance (Oxford, England), 6(4):176–178, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rosling:2009:HR

[389] Hans Rosling. Hans Rosling. Significance (Oxford, England), 6(4):179– 181, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bland:2009:KYB

[390] Martin Bland. Keep young and beautiful: evidence for an “anti-aging” product? Significance (Oxford, England), 6(4):182–183, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jones:2009:SST

[391] Byron Jones. Survival of the sickest: tracking their progress. Significance (Oxford, England), 6(4):184–186, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:Rd

[392] Anonymous. Reviews. Significance (Oxford, England), 6(4):187–188, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:LEd

[393] Anonymous. Letters/ editorial. Significance (Oxford, England), 6(4): 189–190, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2009:Cd

[394] Anonymous. Crossword. Significance (Oxford, England), 6(4):190, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarleton:2009:FHM

[395] Alice Tarleton. FactCheck: how many CCTV cameras?: How two streets in Putney spawned a nationwide statistic that runs and runs. Significance (Oxford, England), 6(4):191–192, December 2009. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:Na

[396] Anonymous. News. Significance (Oxford, England), 7(1):2–4, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anyadike-Danes:2010:WCJ

[397] Michael Anyadike-Danes, Mark Hart, and Karen Bonner. Who creates the jobs? Significance (Oxford, England), 7(1):5–8, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Chandler:2010:CCM

[398] Richard Chandler, Jonathan Rougier, and Mat Collins. Climate change: making certain what the uncertainties are. Significance (Oxford, England), 7(1):9–12, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nelson:2010:PSF

[399] Gerald C. Nelson. The perfect storm: food security and nutrition under climate change. Significance (Oxford, England), 7(1):13–16, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gilligan:2010:BWR

[400] Jonathan Gilligan, Thomas Dietz, Gerald T. Gardner, Paul C. Stern, and Michael P. Vandenbergh. The behavioural wedge: reducing greenhouse gas by individuals and households. Significance (Oxford, England), 7(1): 17–20, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:DFCa

[401] Anonymous. Dr Fisher’s casebook: Intuition. Significance (Oxford, England), 7(1):21, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:CBCa

[402] Anonymous. The 2010 Crystal Ball Competition. Significance (Oxford, England), 7(1):22–23, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Moon:2010:CNT

[403] Nick Moon. Curtains at Number Ten: predicting the general election. Significance (Oxford, England), 7(1):24–26, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Marks:2010:GR

[404] Nic Marks. GDP RIP (1933–2010). Significance (Oxford, England), 7(1): 27–30, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Stone:2010:MS

[405] Mervyn Stone. Mervyn Stone [interview]. Significance (Oxford, England), 7(1):31–33, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mamuneas:2010:WDO

[406] Diamanto Mamuneas and Maria Viskaduraki. Weighing in: does obesity raise the risk from swine flu? Significance (Oxford, England), 7(1): 34–35, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lewin:2010:PSK

[407] Chris Lewin. The Politick Survey of a Kingdome: the first statistical template? Significance (Oxford, England), 7(1):36–39, March 2010. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:Ra

[408] Anonymous. Reviews. Significance (Oxford, England), 7(1):40–42, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:LEa

[409] Anonymous. Letters/ editorial. Significance (Oxford, England), 7(1): 43–44, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:Ca

[410] Anonymous. Crossword. Significance (Oxford, England), 7(1):44, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wong:2010:CCB

[411] Yan Wong. “Cow, compass, bang”. Significance (Oxford, England), 7(1): 45–48, March 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:Nb

[412] Anonymous. News. Significance (Oxford, England), 7(2):50–52, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jarrad:2010:BIB

[413] Frith Jarrad, Peter Whittle, Susan Barrett, and Kerrie Mengersen. Barrow Island’s biosecurity: catching the unknown invader. Significance (Oxford, England), 7(2):53–57, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Balding:2010:GIA

[414] David Balding, Michael Weale, Michael Richards, and Mark Thomas. Genetic and isotopic analysis and the UK Border Agency. Significance (Oxford, England), 7(2):58–61, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Curran:2010:DPR

[415] James Curran. Are DNA profiles as rare as we think? Or can we trust DNA statistics? Significance (Oxford, England), 7(2):62–66, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Leckie:2010:SFN

[416] George Leckie, Rebecca Pillinger, Jennifer Jenkins, and Jon Rasbash. School, family, neighbourhood: which is most important to a child’s education? Significance (Oxford, England), 7(2):67–70, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:DFCb

[417] Anonymous. Dr Fisher*s casebook: a misleading assumption. Significance (Oxford, England), 7(2):71, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Stanley:2010:PPG

[418] Fiona Stanley. Privacy or public good? Why not obtaining consent may be best practice. Significance (Oxford, England), 7(2):72–75, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Simpson:2010:ESB

[419] Edward Simpson. Edward Simpson: Bayes at Bletchley Park. Significance (Oxford, England), 7(2):76–80, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Campbell:2010:SNC

[420] Mike Campbell. A statistician on a NICE committee. Significance (Oxford, England), 7(2):81–84, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hanley:2010:AMP

[421] James Hanley and Elizabeth Turner. Age in medieval plagues and pandemics: Dances of Death or Pearson’s bridge of life? Significance (Oxford, England), 7(2):85–87, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Haigh:2010:IMC

[422] John Haigh. Introduction to Markov chains — the finite case. Significance (Oxford, England), 7(2):88–89, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:Rb

[423] Anonymous. Reviews. Significance (Oxford, England), 7(2):90, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:LEb

[424] Anonymous. Letters/ editorial. Significance (Oxford, England), 7(2): 91–92, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:Cb

[425] Anonymous. Crossword. Significance (Oxford, England), 7(2):92, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Musto:2010:PCR

[426] Garrod Musto and Leslie Cram. Pawprints in clay: Romano–British dogs in a maths class. Significance (Oxford, England), 7(2):93–96, June 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wasserstein:2010:NRS

[427] Ron Wasserstein and Martin Dougherty. New readers start here. Significance (Oxford, England), 7(3):98, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:Nc

[428] Anonymous. News. Significance (Oxford, England), 7(3):99–100, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mengersen:2010:SSL

[429] Kerrie Mengersen. The sound of silence: listening to the villagers to learn about orangutans. Significance (Oxford, England), 7(3):101–106, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lai:2010:DGD

[430] Aaron Lai. Is doing a good deed equal to having a good lunch? Significance (Oxford, England), 7(3):107–111, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Brooker:2010:FHD

[431] Peter Brooker. Fear in a handful of dust: aviation and the Icelandic volcano. Significance (Oxford, England), 7(3):112–115, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spagat:2010:TDI

[432] Michael Spagat. Truth and death in Iraq under sanctions. Significance (Oxford, England), 7(3):116–120, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:DFCc

[433] Anonymous. Dr Fisher’s casebook: The Doctor (nearly) saves the village (or: don’t buy programs, hire people). Significance (Oxford, England), 7 (3):121, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gruber:2010:FHC

[434] Jonathan Gruber and Jasjeet Singh Sekhon. Fundamental health care reform for the United States. Significance (Oxford, England), 7(3):122– 127, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Thompson:2010:KTP

[435] Gavin Thompson. Keeping things in proportion: how can voting systems be fairer? Significance (Oxford, England), 7(3):128–132, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Speed:2010:UES

[436] Terry Speed. The unreasonable effectiveness of statistics in genome science. Significance (Oxford, England), 7(3):133, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wasserstein:2010:GBM

[437] Ron Wasserstein. George Box: a model statistician. Significance (Oxford, England), 7(3):134–135, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hoare:2010:RWW

[438] Zoë Hoare. Randomisation: what, why and how? Significance (Oxford, England), 7(3):136–138, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:Rc

[439] Anonymous. Reviews. Significance (Oxford, England), 7(3):139, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:LEc

[440] Anonymous. Letters/ editorial. Significance (Oxford, England), 7(3): 140–141, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:Cc

[441] Anonymous. Crossword. Significance (Oxford, England), 7(3):141, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dorling:2010:PMP

[442] Danny Dorling. Putting men on a pedestal: Nobel prizes as superhuman myths? Significance (Oxford, England), 7(3):142–144, September 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:Nd

[443] Anonymous. News. Significance (Oxford, England), 7(4):146–148, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

MacDonald:2010:DDH

[444] Ian MacDonald. Deepwater disaster: how the oil spill estimates got it wrong. Significance (Oxford, England), 7(4):149–154, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Solow:2010:SCS

[445] Andrew R. Solow and Derek Tittensor. The sea, the Census and statistics. Significance (Oxford, England), 7(4):155–158, December 2010. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lee:2010:PSS

[446] Rob Lee, Philip Jonathan, and Pauline Ziman. Pictish Symbol Stones: religious imagery, heraldic arms or a language? Significance (Oxford, England), 7(4):159–163, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Press:2010:CTC

[447] William Press. To catch a terrorist: can ethnic profiling work? Significance (Oxford, England), 7(4):164–167, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rumbelow:2010:BBE

[448] Helen Rumbelow. Breast is best? Exposing the myths of breastfeeding. Significance (Oxford, England), 7(4):168–170, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:DFCd

[449] Anonymous. Dr Fisher’s casebook: Keep dancing! Significance (Oxford, England), 7(4):171, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Baskerville:2010:CCS

[450] Jon Baskerville, Jack Miles, Robert Groves, Neville Davies, and Martha Aliaga. Canada census 2011: Not scientifically valid? Significance (Oxford, England), 7(4):172, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Miles:2010:UCL

[451] Jack Miles. UK census 2011 — the last? Significance (Oxford, England), 7(4):173–174, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Groves:2010:UCU

[452] Robert Groves. US census: unsustainable? Significance (Oxford, England), 7(4):174–175, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Davies:2010:C

[453] Neville Davies, Kate Richards, Martha Aliaga, and Rebecca Nichols. CensusAtSchool. Significance (Oxford, England), 7(4):175, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Festing:2010:SAB

[454] Michael Festing. Statistics and animals in biomedical research. Significance (Oxford, England), 7(4):176–177, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2010:BE

[455] Julian Champkin. Bradley Efron. Significance (Oxford, England), 7(4): 178–181, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hewson:2010:BS

[456] Paul Hewson, Aiste Kalinauskaite, Julian Stander, and Yuzhi Cai. Barefoot statisticians. Significance (Oxford, England), 7(4):182–184, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Keyfitz:2010:FFF

[457] Nathan Keyfitz. Fisher and friends: (or: Famous statisticians whom in the course of my 96 years I was lucky enough to meet and get to know). Significance (Oxford, England), 7(4):185, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Boos:2010:EB

[458] Dennis Boos and Leonard Stefanski. Efron’s bootstrap. Significance (Oxford, England), 7(4):186–188, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:Rd

[459] Anonymous. Reviews. Significance (Oxford, England), 7(4):189, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:LEd

[460] Anonymous. Letters/ editorial. Significance (Oxford, England), 7(4): 190–191, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:Cd

[461] Anonymous. Crossword. Significance (Oxford, England), 7(4):191, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2010:CBCb

[462] Anonymous. The 2010 Crystal Ball Competition. Significance (Oxford, England), 7(4):192, December 2010. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:Na

[463] Anonymous. News. Significance (Oxford, England), 8(1):2–4, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Erotokritou-Mulligan:2011:BRD

[464] Ioulietta Erotokritou-Mulligan, Peter Sönksen, and Richard Holt. Beyond reasonable doubt: catching the drug cheats at the London Olympic. Significance (Oxford, England), 8(1):5–9, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gastwirth:2011:DYB

[465] Joseph Gastwirth and Wesley Johnson. Dare you buy a Henry Moore on eBay?: Statistics can tell you what to avoid. Significance (Oxford, England), 8(1):10–14, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kanavos:2011:DUP

[466] Panos Kanavos and Sotiris Vandoros. Drugs US: are prices too high?: Measuring international pharmaceutical price differences. Significance (Oxford, England), 8(1):15–18, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cassimally:2011:WCO

[467] Khalil A. Cassimally. We come from one. Significance (Oxford, England), 8(1):19–21, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Simkin:2011:REA

[468] Mikhail Simkin and Vwani Roychowdhury. Von Richthofen, Einstein and the AGA: Estimating achievement from fame. Significance (Oxford, England), 8(1):22–26, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:DFCa

[469] Anonymous. Dr Fisher’s casebook: Why do young people mumble? Significance (Oxford, England), 8(1):27, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wainer:2011:HSW

[470] Howard Wainer. How should we screen for breast cancer?: Using evidence to make medical decisions. Significance (Oxford, England), 8(1):

28–30, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2011:HVS

[471] Julian Champkin. Hal Varian and the sexy profession. Significance (Oxford, England), 8(1):31–34, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Speed:2011:TMT

[472] Terry Speed. Too much or too little scepticism? Significance (Oxford, England), 8(1):35, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kovalchik:2011:MSC

[473] Stephanie Kovalchik. Multiple sclerosis: a cure or a contradiction?: Inconsistent science leaves patients guessing. Significance (Oxford, England), 8(1):36–38, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2011:MIB

[474] Julian Champkin. Making information beautiful — and clear. Significance (Oxford, England), 8(1):39–41, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:Ra

[475] Anonymous. Reviews. Significance (Oxford, England), 8(1):42, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:LEa

[476] Anonymous. Letters/ editorial. Significance (Oxford, England), 8(1): 43–45, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:Ca

[477] Anonymous. Crossword. Significance (Oxford, England), 8(1):45, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tibshirani:2011:DTT

[478] Ryan Tibshirani. Don’t try for the triple 20: Where to aim if you are bad at darts. Significance (Oxford, England), 8(1):46–48, March 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:Nb

[479] Anonymous. News. Significance (Oxford, England), 8(2):50–52, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Canton:2011:EFM

[480] James Canton. The extreme future of megacities. Significance (Oxford, England), 8(2):53–56, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Teng:2011:UFS

[481] Paul Teng, Margarita Escaler, and Mely Caballero-Anthony. Urban food security: Feeding tomorrow’s cities. Significance (Oxford, England), 8 (2):57–60, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Smaje:2011:UCP

[482] Chris Smaje. The ungreen city — or the polluting countryside? Significance (Oxford, England), 8(2):61–64, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Coates:2011:CNS

[483] Dennis Coates and Brad Humphreys. Can new stadiums revitalise urban neighbourhoods? Significance (Oxford, England), 8(2):65–69, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Banta-Green:2011:CWD

[484] Caleb Banta-Green and Jennifer Field. City-wide drug testing using municipal wastewater: a new tool for drug epidemiology. Significance (Oxford, England), 8(2):70–74, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wu:2011:TSU

[485] Tong Wu. Two-speed urbanisation: Shrinking cities amid urban congestion. Significance (Oxford, England), 8(2):75–78, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:DFCb

[486] Anonymous. Dr Fisher’s casebook: City dreams. Significance (Oxford, England), 8(2):79, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Pine:2011:SWM

[487] Karen Pine. Sheconomics: Why more women on boards boosts company performance. Significance (Oxford, England), 8(2):80–81, June 2011. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lai:2011:LCB

[488] Aaron Lai. London cholera and the blind-spot of an epidemiology theory. Significance (Oxford, England), 8(2):82–85, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Pitfield:2011:DPS

[489] David Pitfield. Dungeness: Plane strikes nuclear power plant? Significance (Oxford, England), 8(2):86–88, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Griffiths:2011:TPD

[490] Martin Griffiths. Trying to pull out of the drive. Significance (Oxford, England), 8(2):89–91, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:Rb

[491] Anonymous. Reviews. Significance (Oxford, England), 8(2):92, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:LEb

[492] Anonymous. Letters/ editorial. Significance (Oxford, England), 8(2): 93–94, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:Cb

[493] Anonymous. Crossword. Significance (Oxford, England), 8(2):94, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Groves:2011:BWC

[494] Jeremy Groves. Bicycle weight and commuting time: a randomised trial. Significance (Oxford, England), 8(2):95–96, June 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:Nc

[495] Anonymous. News. Significance (Oxford, England), 8(3):98–99, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spiegelhalter:2011:FNF

[496] David Spiegelhalter. Fear and numbers in Fukushima. Significance (Oxford, England), 8(3):100–103, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Watson:2011:HMW

[497] Ray Watson and Paul Yip. How many were there when it mattered?: Estimating the sizes of crowds. Significance (Oxford, England), 8(3):104– 107, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mortimer:2011:EBL

[498] Nigel Mortimer. Ethics for biofuels … and everything else. Significance (Oxford, England), 8(3):108–111, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:DFCc

[499] Anonymous. Dr Fisher’s casebook: Live long and prosper. Significance (Oxford, England), 8(3):112, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ince:2011:DUS

[500] Darrel Ince. The Duke University scandal — what can be done? Significance (Oxford, England), 8(3):113–115, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Young:2011:DDO

[501] S. Stanley Young and Alan Karr. Deming, data and observational studies: a process out of control and needing fixing. Significance (Oxford, England), 8(3):116–120, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2011:FGC

[502] Julian Champkin. Francis Galton centenary. Significance (Oxford, England), 8(3):121, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Maier:2011:FGM

[503] Dan Maier. Francis Galton: Measuring the immeasurable. Significance (Oxford, England), 8(3):122–123, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Senn:2011:FGR

[504] Stephen Senn. Francis Galton and regression to the mean. Significance (Oxford, England), 8(3):124–126, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2011:EPP

[505] Julian Champkin. Eveline Pye: Poetry in numbers. Significance (Oxford, England), 8(3):127–130, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ziliak:2011:MVS

[506] Stephen T. Ziliak. Matrixx v. Siracusano and Student v. Fisher: Statistical significance on trial. Significance (Oxford, England), 8(3):131–134, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Niemi:2011:SGM

[507] Jarad Niemi and Andrew Gelman. Statistical graphics: Making information clear — and beautiful. Significance (Oxford, England), 8(3):135–137, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:Rc

[508] Anonymous. Reviews. Significance (Oxford, England), 8(3):138–139, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:ELa

[509] Anonymous. Editorial/ letters. Significance (Oxford, England), 8(3):140– 141, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:Cc

[510] Anonymous. Crossword. Significance (Oxford, England), 8(3):141, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bottone:2011:GM

[511] Michele Bottone. Galton’s millions. Significance (Oxford, England), 8 (3):142–144, September 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:Nd

[512] Anonymous. News. Significance (Oxford, England), 8(4):146–147, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Palmer:2011:WBM

[513] David Palmer. Will the big machine arrive?: Engineering the Miami Tunnel. Significance (Oxford, England), 8(4):148–153, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Collins:2011:CMM

[514] Tom Collins, Robin Laney, Alistair Willis, and Paul H. Garthwaite. Chopin, mazurkas and Markov: Making music in style with statistics. Significance (Oxford, England), 8(4):154–159, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

LeSage:2011:DWN

[515] James LeSage, R. Kelley Pace, Richard Campanella, Nina Lam, and Xingjian Liu. Do what the neighbours do: Reopening businesses after Hurricane Katrina. Significance (Oxford, England), 8(4):160–163, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Weir:2011:FCT

[516] Christopher Weir and Gordon Murray. Fraud in clinical trials: Detecting it and preventing it. Significance (Oxford, England), 8(4):164–168, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:DFCd

[517] Anonymous. Dr. Fisher’s casebook: What are little statisticians made of? Significance (Oxford, England), 8(4):169, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wainer:2011:ALT

[518] Howard Wainer and Lawrence Hubert. Assessing long-term risk with short-term data. Significance (Oxford, England), 8(4):170–171, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Liechty:2011:FH

[519] John C. Liechty and Richard Foster. Financial hurricanes. Significance (Oxford, England), 8(4):172–174, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2011:CRR

[520] Julian Champkin. C. R. Rao. Significance (Oxford, England), 8(4):175– 178, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Smith:2011:HS

[521] Simon Smith and Martin Forster. The hazards of slavery. Significance (Oxford, England), 8(4):179–182, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gelman:2011:TGR

[522] Andrew Gelman. Tables as graphs: the Ramanujan principle. Significance (Oxford, England), 8(4):183, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:Rd

[523] Anonymous. Reviews. Significance (Oxford, England), 8(4):184, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Stefani:2011:MBP

[524] Ray Stefani and Jim Albert. Moneyball: Brad Pitt, the statistician and the movie. Significance (Oxford, England), 8(4):185–186, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gray:2011:WS

[525] Laura Gray. Writing with Significance! Significance (Oxford, England), 8(4):186, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:ELb

[526] Anonymous. Editorial/ letters. Significance (Oxford, England), 8(4):187– 189, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2011:Cd

[527] Anonymous. Crossword. Significance (Oxford, England), 8(4):189, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anderson:2011:MMS

[528] Jesse Anderson. A million monkeys and Shakespeare. Significance (Oxford, England), 8(4):190–192, December 2011. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Na

[529] Anonymous. News. Significance (Oxford, England), 9(1):2–3, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ploutz-Snyder:2012:BSG

[530] Robert Ploutz-Snyder. To boldly and safely go: Biostatistics in space. Significance (Oxford, England), 9(1):4–7, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Scott:2012:IPE

[531] Marian Scott. Increasing population and the environment: How do the numbers stack up? Significance (Oxford, England), 9(1):8–12, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hawkins:2012:OEC

[532] Ed Hawkins. Our evolving climate. Significance (Oxford, England), 9(1): 13–15, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Salmon:2012:FKW

[533] Felix Salmon. The formula that killed Wall Street. Significance (Oxford, England), 9(1):16–20, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Neumann:2012:FCS

[534] Cedric Neumann. Fingerprints at the crime-scene: Statistically certain, or probable? Significance (Oxford, England), 9(1):21–25, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McHale:2012:PLL

[535] Ian McHale and Rose D. Baker. Playing the lottery with a little bit of stats know-how. … Significance (Oxford, England), 9(1):25–28, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:DFCa

[536] Anonymous. Dr Fisher’s casebook: And the first shall be last. Significance (Oxford, England), 9(1):29, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Briggs:2012:WDS

[537] Matt Briggs. Why do statisticians answer silly questions that no one ever asks? Significance (Oxford, England), 9(1):30–31, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Khatab:2012:ASW

[538] Khaled Khatab. The Arab Spring: Where is Egypt now? Significance (Oxford, England), 9(1):32–34, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Woodward:2012:PWP

[539] Philip Woodward. Philip Woodward on Philip Woodward: Radar, clocks and coincidences. Significance (Oxford, England), 9(1):35–39, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jones:2012:WCU

[540] Byron Jones and Robb Muirhead. What a coincidence! It’s not as unlikely as you think. Significance (Oxford, England), 9(1):40–42, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Ra

[541] Anonymous. Reviews. Significance (Oxford, England), 9(1):43, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:EL

[542] Anonymous. Editorial/ letters. Significance (Oxford, England), 9(1):44– 45, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Ca

[543] Anonymous. Crossword. Significance (Oxford, England), 9(1):45, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harkness:2012:SS

[544] Timandra Harkness. Seduced by stats? Significance (Oxford, England), 9(1):46–48, February 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Nb

[545] Anonymous. News. Significance (Oxford, England), 9(2):2–3, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Stefani:2012:ODL

[546] Ray Stefani. Overcoming the doping legacy: Can London’s winners outperform the drugs of 1988? Significance (Oxford, England), 9(2):4–8, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Barrow:2012:HUB

[547] John D. Barrow. How Usain Bolt can run faster — effortlessly. Significance (Oxford, England), 9(2):9–12, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Stefani:2012:OSG

[548] Ray Stefani. Olympic swimming gold: the suit or the swimmer in the suit? Significance (Oxford, England), 9(2):13–17, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kovalchik:2012:MRW

[549] Stephanie Kovalchik. Men’s records and women’s: are the women better already?: Moving towards a gender-neutral Olympics. Significance (Oxford, England), 9(2):18–23, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cheah:2012:MLK

[550] Benjamin Cheah. Is the Marathon likely to kill you?: Fraught with danger or fraught with fear? Significance (Oxford, England), 9(2):24–26, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:DFCb

[551] Anonymous. Dr Fisher’s casebook: Sporting Life. Significance (Oxford, England), 9(2):27, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wainer:2012:PVF

[552] Howard Wainer. Piano virtuosos and the four-minute mile. Significance (Oxford, England), 9(2):28–29, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2012:MIO

[553] Julian Champkin. Mark Israney: Olympian odds-setter. Significance (Oxford, England), 9(2):30–33, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Haigh:2012:LJH

[554] John Haigh. The long jump: How to jump the same distance and win. Significance (Oxford, England), 9(2):34–36, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Christesen:2012:WWF

[555] Paul Christesen. When were the first Olympics? Significance (Oxford, England), 9(2):37–39, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fanshawe:2012:STP

[556] Tom Fanshawe. Seven into two: Principal components analysis and the Olympic heptathlon. Significance (Oxford, England), 9(2):40–42, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Rb

[557] Anonymous. Reviews. Significance (Oxford, England), 9(2):43, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:LEa

[558] Anonymous. Letters/ editorial. Significance (Oxford, England), 9(2): 44–45, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Cb

[559] Anonymous. Crossword. Significance (Oxford, England), 9(2):45, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Davies:2012:HFF

[560] Neville Davies, John Marriott, Kate Richards, and Lauren Adams. How far to the flame? Getting to the Point in 2012. Significance (Oxford, England), 9(2):46–48, April 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Nc

[561] Anonymous. News. Significance (Oxford, England), 9(3):2–3, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fretwell:2012:ESB

[562] Peter Fretwell. The Emperor strikes back. Significance (Oxford, England), 9(3):4–7, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cole:2012:PSS

[563] Tim Cole. People smugglers, statistics and bone age. Significance (Oxford, England), 9(3):8–12, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Green:2012:LSP

[564] Ken Green. London 2012 and sports participation: the myths of legacy. Significance (Oxford, England), 9(3):13–16, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reilly:2012:RBC

[565] Barry Reilly, Neil Rickman, and Robert Witt. Robbing banks: Crime does pay — but not very much. Significance (Oxford, England), 9(3): 17–21, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:DFCc

[566] Anonymous. Dr Fisher’s casebook: Fly me to the Moon. Significance (Oxford, England), 9(3):22, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Morris:2012:WPP

[567] Danielle Morris. What’s the point of a point estimate?: Why statistics lectures confuse students. Significance (Oxford, England), 9(3):23–25, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

OHagan:2012:TB

[568] Tony O’Hagan. A thing of beauty. Significance (Oxford, England), 9 (3):26–28, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Baker:2012:MSS

[569] Leanne F. Baker and Joseph F. Mudge. Making statistical significance more significant. Significance (Oxford, England), 9(3):29–30, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cook:2012:CHM

[570] Robert Cook and Howard Wainer. A century and a half of moral statistics in the United Kingdom: Variations on Joseph Fletcher’s thematic maps. Significance (Oxford, England), 9(3):31–36, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Millner:2012:FBM

[571] Antony Millner and Raphael Calel. Are first-borns more likely to attend Harvard? Significance (Oxford, England), 9(3):37–39, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

VanDerwerken:2012:WPD

[572] Douglas VanDerwerken. What Petri dishes have to do with your research. Significance (Oxford, England), 9(3):40–42, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Rc

[573] Anonymous. Reviews. Significance (Oxford, England), 9(3):43–44, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:LEb

[574] Anonymous. Letters/ editorial. Significance (Oxford, England), 9(3): 45–46, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Cc

[575] Anonymous. Crossword. Significance (Oxford, England), 9(3):46, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ankers:2012:SWS

[576] John Ankers. The size of Wales: Statistics as weapons. Significance (Oxford, England), 9(3):47–48, June 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Nd

[577] Anonymous. News. Significance (Oxford, England), 9(4):2–3, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Keller:2012:BDC

[578] Sallie Ann Keller, Steven E. Koonin, and Stephanie Shipp. Big data and city living — what can it do for us? Significance (Oxford, England), 9 (4):4–7, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hilbert:2012:HMI

[579] Martin Hilbert. How much information is there in the “information society”? Significance (Oxford, England), 9(4):8–12, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ma:2012:SJC

[580] Ping Ma. A statistical journey to the centre of the Earth. Significance (Oxford, England), 9(4):13–16, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Johnston:2012:CWE

[581] Iain Johnston. The chaos within: Exploring noise in cellular biology. Significance (Oxford, England), 9(4):17–21, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Feigelson:2012:BDA

[582] Eric D. Feigelson and G. Jogesh Babu. Big data in astronomy. Significance (Oxford, England), 9(4):22–25, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lansdall-Welfare:2012:NMN

[583] Thomas Lansdall-Welfare, Vasileios Lampos, and Nello Cristianini. Nowcasting the mood of the nation. Significance (Oxford, England), 9(4): 26–28, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:DFCd

[584] Anonymous. Dr Fisher’s casebook: Small is beautiful. Significance (Oxford, England), 9(4):29, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Norvig:2012:CGI

[585] Peter Norvig. Colorless green ideas learn furiously: Chomsky and the two cultures of statistical learning. Significance (Oxford, England), 9(4): 30–33, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Roy:2012:DF

[586] Deb Roy. The data family. Significance (Oxford, England), 9(4):34– 37, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Emerson:2012:DDD

[587] John W. Emerson and Michael J. Kane. Don’t drown in the data. Significance (Oxford, England), 9(4):38–39, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Salter-Townshend:2012:AMF

[588] Michael Salter-Townshend. Analysing my Facebook friends. Significance (Oxford, England), 9(4):40–42, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Rd

[589] Anonymous. Reviews. Significance (Oxford, England), 9(4):43–44, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:LEc

[590] Anonymous. Letters/ editorial. Significance (Oxford, England), 9(4): 45–46, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rose:2012:BDF

[591] Sherri Rose. Big data and the future. Significance (Oxford, England), 9(4):47–48, August 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Ne

[592] Anonymous. News. Significance (Oxford, England), 9(5):2–4, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Baxter:2012:PPH

[593] Mike Baxter. Picturing the pictures: Hitchcock, statistics and film. Significance (Oxford, England), 9(5):5–9, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wijlaars:2012:SEA

[594] Linda Wijlaars. Side-effects in antidepressants: the drug or the disease? Significance (Oxford, England), 9(5):10–13, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Brown:2012:LUI

[595] Ethan Brown and Nick Bearman. Listening to uncertainty: Information that sings. Significance (Oxford, England), 9(5):14–17, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cochran:2012:FP

[596] James J. Cochran and David J. Curry. Frankenstein for President. Significance (Oxford, England), 9(5):18–22, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

VanDerwerken:2012:NPW

[597] Douglas VanDerwerken. The next President: Will your vote decide it? Significance (Oxford, England), 9(5):23–26, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:DFCe

[598] Anonymous. Dr Fisher’s casebook: Export or die. Significance (Oxford, England), 9(5):27, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2012:AC

[599] Julian Champkin. Amanda Cox. Significance (Oxford, England), 9(5): 28–31, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Brimicombe:2012:BWL

[600] Allan Brimicombe and Rebecca Cafe. Beware, win or lose: Domestic violence and the World Cup. Significance (Oxford, England), 9(5):32– 35, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2012:MGW

[601] Julian Champkin. 2008 movies, by Graham Wills. Significance (Oxford, England), 9(5):36–37, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fanshawe:2012:ELG

[602] Tom Fanshawe. Eyam and “the last great visitation”. Significance (Oxford, England), 9(5):38–41, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Re

[603] Anonymous. Reviews. Significance (Oxford, England), 9(5):42, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:LEd

[604] Anonymous. Letters/ editorial. Significance (Oxford, England), 9(5): 43–44, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McConway:2012:SIR

[605] Kevin McConway and David Spiegelhalter. Score and ignore: a radio listener’s guide to ignoring health stories. Significance (Oxford, England), 9(5):45–48, October 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Nf

[606] Anonymous. News. Significance (Oxford, England), 9(6):2–3, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Araya-Salas:2012:BM

[607] Marcelo Araya-Salas. Is birdsong music? Significance (Oxford, England), 9(6):4–7, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bevan:2012:REC

[608] Godfrey Bevan. Renewable energy and climate change. Significance (Oxford, England), 9(6):8–12, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Prats:2012:LES

[609] Jordi Prats. The L’Aquila earthquake: Science or risk on trial? Significance (Oxford, England), 9(6):13–16, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spiegelhalter:2012:BIS

[610] David Spiegelhalter, Stuart Knight, and Tony Sant. Breast implants: the scandal, the outcry, and assessing the risks. Significance (Oxford, England), 9(6):17–21, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kenna:2012:SSC

[611] Ralph Kenna and Bertrand Berche. Statistics of statisticians: Critical masses for research groups. Significance (Oxford, England), 9(6):22– 25, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:DFCf

[612] Anonymous. Dr Fisher’s casebook: Job descriptions. Significance (Oxford, England), 9(6):26, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Edwards:2012:RFY

[613] Professor A. E. W. Edwards and Professor Sir Walter Bodmer. R. A. Fisher — 50 years on. Significance (Oxford, England), 9(6):27–29, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Senn:2012:TTI

[614] Stephen Senn. Tea for three: Of infusions and inferences and milk in first. Significance (Oxford, England), 9(6):30–33, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Carriquiry:2012:GB

[615] Alicia Carriquiry. Graciela Bevacqua. Significance (Oxford, England), 9 (6):34–36, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wilmshurst:2012:ELL

[616] Peter Wilmshurst. English libel laws and scientific research. Significance (Oxford, England), 9(6):37–39, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2012:PML

[617] Julian Champkin. Polyconic man: by Lilla LoCurto and Bill Outcault. Significance (Oxford, England), 9(6):40–41, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:Rf

[618] Anonymous. Reviews. Significance (Oxford, England), 9(6):42–43, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2012:ELC

[619] Anonymous. Editorial, Letters, Crossword. Significance (Oxford, England), 9(6):44–45, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ross:2012:FMS

[620] Gavin Ross. Fisher and the millionaire: the statistician and the calculator. Significance (Oxford, England), 9(6):46–48, December 2012. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Na

[621] Anonymous. News. Significance (Oxford, England), 10(1):2–3, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hegre:2013:PEF

[622] Håvard Hegre. Peace on Earth?: the future of internal armed conflict. Significance (Oxford, England), 10(1):4–8, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McCarthy:2013:CC

[623] Donal McCarthy. Costing conservation. Significance (Oxford, England), 10(1):9–13, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Moller:2013:ELD

[624] Anders Pape Møller and Timothy A. Mousseau. The effects of low-dose radiation: Soviet science, the nuclear industry — and independence? Significance (Oxford, England), 10(1):14–19, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kasdan:2013:SRT

[625] David Oliver Kasdan. Seven in a row?: The Tour de France and the likelihood of Lance. Significance (Oxford, England), 10(1):20–22, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wasserstein:2013:IYS

[626] Ron Wasserstein. International Year of Statistics: It’s the International Year of Statistics! Significance (Oxford, England), 10(1):23, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:SIY

[627] Anonymous. Significance International Year of Statistics pull-out poster. Significance (Oxford, England), 10(1):24–25, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dwan:2013:YSW

[628] Kerry Dwan, Laura Gray, and Julian Champkin. Young Statisticians Writing Competition: Celebrate the International Year of Statistics in 2013 by writing with Significance! Significance (Oxford, England), 10 (1):26, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:DFCa

[629] Anonymous. Dr Fisher’s casebook: Take a chance on me. Significance (Oxford, England), 10(1):27, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Varshney:2013:WDW

[630] Lav R. Varshney and John Z. Sun. Why do we perceive logarithmically? Significance (Oxford, England), 10(1):28–31, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Grant:2013:NY

[631] Robert Grant. Nathan Yau. Significance (Oxford, England), 10(1):32– 35, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hooper:2013:RPB

[632] Martyn Hooper. Richard Price, Bayes’ theorem, and God. Significance (Oxford, England), 10(1):36–39, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:LHW

[633] Anonymous. Laconic History of the World, by Martin Elmer. Significance (Oxford, England), 10(1):40–41, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Ra

[634] Anonymous. Reviews. Significance (Oxford, England), 10(1):42–43, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:ELCa

[635] Anonymous. Editorial/letters/crossword. Significance (Oxford, England), 10(1):44–45, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Perlman:2013:WWW

[636] Patrik Perlman. When will we see people of negative height? Significance (Oxford, England), 10(1):46–48, February 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Nb

[637] Anonymous. News. Significance (Oxford, England), 10(2):2–5, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Baxter:2013:PTD

[638] Mike Baxter. Pompeii, tombs and drinking habits: Archaeology counts. Significance (Oxford, England), 10(2):6–11, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tyrer:2013:SRE

[639] Mark Tyrer and John P. Sykes. The statistics of the rare earths industry. Significance (Oxford, England), 10(2):12–16, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Oliver:2013:PAG

[640] Margaret A. Oliver. Precision agriculture and geostatistics: How to manage agriculture more exactly. Significance (Oxford, England), 10(2):17– 22, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Duffy:2013:PDT

[641] John C. Duffy. The price of a drink — too exactly? Flawed evidence for minimum unit pricing. Significance (Oxford, England), 10(2):23–27, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:DFCb

[642] Anonymous. Dr Fisher’s casebook: Codex magistri Piscatori. Significance (Oxford, England), 10(2):28, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Orelien:2013:HAE

[643] Jean Orelien, Robert Philippe, Antwone Wesner, James Ashley, Justin Fisher, and Fritz Scheuren. Haiti after the earthquake: Statistics Without Borders. Significance (Oxford, England), 10(2):29–32, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Moodie:2013:HBW

[644] Erica Moodie, Amee Manges, and James Hanley. Hey baby, what’s your sign? How children born under Sagittarius are denied day care. Significance (Oxford, England), 10(2):33–36, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dorling:2013:IN

[645] Danny Dorling. It is necessarily so. Significance (Oxford, England), 10 (2):37–39, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Senn:2013:II

[646] Stephen Senn. Invalid inversion. Significance (Oxford, England), 10(2): 40–42, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Rb

[647] Anonymous. Reviews. Significance (Oxford, England), 10(2):43–44, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:ELa

[648] Anonymous. Editorial/ letters. Significance (Oxford, England), 10(2): 45–46, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Ca

[649] Anonymous. Crossword. Significance (Oxford, England), 10(2):46, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wallace:2013:FMU

[650] Michael Wallace and John Haigh. Football and marriage — and the UEFA draw. Significance (Oxford, England), 10(2):47–48, April 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Nc

[651] Anonymous. News. Significance (Oxford, England), 10(3):2–5, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Goulson:2013:NBW

[652] Dave Goulson. Neonicotinoids and bees: What’s all the buzz? Significance (Oxford, England), 10(3):6–11, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mardia:2013:ABD

[653] Kanti Mardia, Fred Bookstein, and John Kent. Alcohol, babies and the death penalty: Saving lives by analysing the shape of the brain. Significance (Oxford, England), 10(3):12–16, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hirst:2013:BW

[654] Bill Hirst, Philip Jonathan, and David Randell. Blowin’ in the wind. Significance (Oxford, England), 10(3):17–19, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Brewer:2013:ACL

[655] Lauren M. Brewer and James J. Cochran. August’s child is … favoured by fortune. Significance (Oxford, England), 10(3):20–24, June 2013. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Skinner:2013:PAR

[656] Brian Skinner and Brad Carlin. The price of anarchy — on the roads and in football. Significance (Oxford, England), 10(3):25–30, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:DFCc

[657] Anonymous. Dr Fisher’s casebook: Beard the statistician in his den. Significance (Oxford, England), 10(3):31, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:GPB

[658] Anonymous. George E. P. Box, 1919–2013. Significance (Oxford, England), 10(3):32–34, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Budrus:2013:TDS

[659] Sarah Budrus, Susan Vander Plas, and Dianne Cook. In tennis, do smashes win matches? Significance (Oxford, England), 10(3):35–38, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Edwards:2013:ACT

[660] Professor Anthony Edwards. Ars conjectandi three hundred years on. Significance (Oxford, England), 10(3):39–41, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:IMX

[661] Anonymous. Interstellar memes on xkcd. Significance (Oxford, England), 10(3):42–43, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Rc

[662] Anonymous. Reviews. Significance (Oxford, England), 10(3):44, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:ELb

[663] Anonymous. Editorial/ letters. Significance (Oxford, England), 10(3): 45–47, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Cb

[664] Anonymous. Crossword. Significance (Oxford, England), 10(3):47, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2013:DS

[665] Julian Champkin. Dear Sir. Significance (Oxford, England), 10(3):48, June 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Nd

[666] Anonymous. News. Significance (Oxford, England), 10(4):2–3, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McInerney:2013:BBR

[667] James McInerney, Alex Rogers, and Nicholas R. Jennings. Bus, bike and random journeys: Crowdsourcing aid distribution in Ivory Coast. Significance (Oxford, England), 10(4):4–9, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mannes:2013:KIR

[668] Albert Mannes and Don Moore. I know I’m right! A behavioural view of overconfidence. Significance (Oxford, England), 10(4):10–14, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Smith:2013:PCT

[669] Ben Smith and Jadrian Wooten. Pundits: the confidence trick: Better confident than right? Significance (Oxford, England), 10(4):15–18, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:DFCd

[670] Anonymous. Dr Fisher’s casebook: Don’t shoot the piano player, he’s doing the best he can. Significance (Oxford, England), 10(4):19, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Walker:2013:BDB

[671] David Walker and Kaiser Fung. Big data and big business: Should statisticians join in? Significance (Oxford, England), 10(4):20–25, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lewis:2013:MAS

[672] Michael A. Lewis. Misgivings about smart giving: Some complications of the Robin Hood model of philanthropy. Significance (Oxford, England), 10(4):26–29, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Thurstan:2013:FPW

[673] Ruth Thurstan. Fishing up the past: What historical records can tell us about marine populations today. Significance (Oxford, England), 10(4): 30–33, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Grajales:2013:SSJ

[674] Carlos Gómez Grajales. Singing surveys and Jedi knights: How not to ask questions. Significance (Oxford, England), 10(4):34–37, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Arnall:2013:ILP

[675] Timo Arnall, Jørn Knutsen, and Einar Sneve Martinussen. Immaterials: Light painting WiFi. Significance (Oxford, England), 10(4):38–39, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Rd

[676] Anonymous. Reviews. Significance (Oxford, England), 10(4):40–41, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:ELd

[677] Anonymous. Editorial/ letters. Significance (Oxford, England), 10(4): 42–44, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Cc

[678] Anonymous. Crossword. Significance (Oxford, England), 10(4):44, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Vukcevic:2013:YWC

[679] Damjan Vukcevic. Young Writers Competition: GUESTimation: Breaking the deadlock on wedding guest lists. Significance (Oxford, England), 10(4):45–48, August 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Ne

[680] Anonymous. News. Significance (Oxford, England), 10(5):2–5, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cook:2013:SHM

[681] Nancy R. Cook. Salt: How much less should we eat for health?: Understanding the recent IOM report. Significance (Oxford, England), 10(5): 6–10, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rushby:2013:MWO

[682] Andrew Rushby. A multiplicity of worlds: Other habitable planets. Significance (Oxford, England), 10(5):11–15, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hamilton:2013:CCW

[683] Ian Hamilton. The cannabis class: What happened when the legal status of cannabis was reclassified? Significance (Oxford, England), 10(5):16– 19, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:DFCe

[684] Anonymous. Dr Fisher’s casebook: the shoulders of giants. Significance (Oxford, England), 10(5):20, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kelly:2013:EDM

[685] Mark Kelly. Emily Dickinson and monkeys on the stair Or: What is the significance of the 5% significance level? Significance (Oxford, England), 10(5):21–22, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2013:LK

[686] Julian Champkin. Lord Krebs. Significance (Oxford, England), 10(5): 23–29, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2013:CFC

[687] Julian Champkin. A carriage full of curates. Significance (Oxford, England), 10(5):30–31, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:RCR

[688] Anonymous. Readers’ Challenge report: The Monty Hall problem. Significance (Oxford, England), 10(5):32–33, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lee:2013:UCM

[689] Lindsay Lee. Uncertainties in climate models: Living with uncertainty in an uncertain world. Significance (Oxford, England), 10(5):34–39, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Demsar:2013:VMS

[690] Urška Demšar and Emiel van Loon. Visualising movement: the seagull. Significance (Oxford, England), 10(5):40–42, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Re

[691] Anonymous. Reviews. Significance (Oxford, England), 10(5):43–44, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:ELCb

[692] Anonymous. Editorial/letters/crossword. Significance (Oxford, England), 10(5):45–46, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Heid:2013:TPH

[693] Cory Heid. Tootsie Pops: How many licks to the chocolate? Significance (Oxford, England), 10(5):47–48, October 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Nf

[694] Anonymous. News. Significance (Oxford, England), 10(6):3–5, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Craze:2013:EHE

[695] Paul Craze. Early human evolution and the skulls of Dmanisi. Significance (Oxford, England), 10(6):6–11, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Carron:2013:VSS

[696] Pádraig Mac Carron and Ralph Kenna. Viking sagas: Six degrees of Icelandic separation: Social networks from the Viking era. Significance (Oxford, England), 10(6):12–17, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Borja:2013:SBP

[697] Mario Cortina Borja. The strong birthday problem. Significance (Oxford, England), 10(6):18–20, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Grajalez:2013:GMS

[698] Carlos Gómez Grajalez, Eileen Magnello, Robert Woods, and Julian Champkin. Great moments in statistics. Significance (Oxford, England), 10(6):21–28, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:TS

[699] Anonymous. Timeline of statistics. Significance (Oxford, England), 10 (6):23–26, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Yang:2013:WAG

[700] Charles Yang. Who’s afraid of George Kingsley Zipf? Or: Do children and chimps have language? Significance (Oxford, England), 10(6):29– 34, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:DFCf

[701] Anonymous. Dr Fisher’s casebook: a refereeing whirl. Significance (Oxford, England), 10(6):35, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2013:NS

[702] Julian Champkin and Alison Oliver. Nate Silver. Significance (Oxford, England), 10(6):36–39, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Peacock:2013:WGF

[703] Eloise Peacock. When the gambler’s fallacy comes true: Beating the online casino. Significance (Oxford, England), 10(6):40–42, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Rf

[704] Anonymous. Reviews. Significance (Oxford, England), 10(6):43, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:ELe

[705] Anonymous. Editorial/ letters. Significance (Oxford, England), 10(6): 44–45, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2013:Cd

[706] Anonymous. Crossword. Significance (Oxford, England), 10(6):45, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Aberdein:2013:LRB

[707] Jody Aberdein and David Spiegelhalter. Have London’s roads become more dangerous for cyclists? Significance (Oxford, England), 10(6):46– 48, December 2013. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:Na

[708] Anonymous. News. Significance (Oxford, England), 11(1):2–3, February 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Zeid:2014:UCS

[709] Alaa Abou Zeid and James J. Cochran. Understanding the crisis in Somalia. Significance (Oxford, England), 11(1):4–9, February 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cobham:2014:IAA

[710] Alex Cobham and Andy Sumner. Is inequality all about the tails?: The Palma measure of income inequality. Significance (Oxford, England), 11 (1):10–13, February 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Plewis:2014:IFS

[711] Ian Plewis. Indian farmer suicides: Is GM cotton to blame? Significance (Oxford, England), 11(1):14–18, February 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mackey:2014:CTC

[712] Brendan Mackey. Counting trees, carbon and climate change. Significance (Oxford, England), 11(1):19–23, February 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hill:2014:MWD

[713] Timothy Martyn Hill and Alan Smith. Migrants: Where do they come from? Significance (Oxford, England), 11(1):24–29, February 2014. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:DFCa

[714] Anonymous. Dr Fisher’s casebook: Analysis of variance. Significance (Oxford, England), 11(1):30, February 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Barder:2014:AWM

[715] Owen Barder. Is aid a waste of money? Significance (Oxford, England), 11(1):31–32, February 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2014:RK

[716] Julian Champkin. Richard King. Significance (Oxford, England), 11(1): 33–36, February 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rispoli:2014:EBF

[717] Fred J. Rispoli, Suhua Zeng, Tim Green, and Jennifer Higbie. Even birds follow Pareto’s 80–20 rule. Significance (Oxford, England), 11(1): 37–38, February 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2014:CGE

[718] Julian Champkin. The champagne glass effect. Significance (Oxford, England), 11(1):39–41, February 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:Ra

[719] Anonymous. Reviews. Significance (Oxford, England), 11(1):42–44, February 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:YSW

[720] Anonymous. Young Statisticians Writing Competition: Write with Significance! Significance (Oxford, England), 11(1):44, February 2014. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:ELCa

[721] Anonymous. Editorial/letters/crossward. Significance (Oxford, England), 11(1):45–46, February 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dudzic:2014:NAO

[722] Stella Dudzic. North Americans — one third of the world’s weight?: Obesity and the comma. Significance (Oxford, England), 11(1):47–48, February 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:Nb

[723] Anonymous. News. Significance (Oxford, England), 11(2):2–5, April 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gius:2014:GLC

[724] Mark Gius. Gun law and crime. Significance (Oxford, England), 11 (2):6–8, April 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nagin:2014:DDP

[725] Daniel Nagin. Deterrence and the death penalty: Why the statistics should be ignored. Significance (Oxford, England), 11(2):9–13, April 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wingfield:2014:JJU

[726] Melissa Wingfield and Paul Trenell. Joined up justice: Understanding the links between employment, benefits and offending. Significance (Oxford, England), 11(2):14–19, April 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Millar:2014:ETN

[727] Keith Millar, Adrian Bowman, and William Battersby. The Erebus, the Terror and the North-West Passage: Did lead really poison Franklin’s lost expedition? Significance (Oxford, England), 11(2):20–26, April 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:DFCb

[728] Anonymous. Dr Fisher’s casebook: Primary colours. Significance (Oxford, England), 11(2):27, April 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

vanBelle:2014:DLB

[729] Gerald van Belle and Leslie Ruiter. Data and the law: Beyond the sweat of the brow: Who owns published data? And what is data? Significance (Oxford, England), 11(2):28–31, April 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bland:2014:HFT

[730] Martin Bland. Health freaks on trial: Duct tape, bull semen and the call of television. Significance (Oxford, England), 11(2):32–35, April 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Farebrother:2014:ABM

[731] Richard Farebrother and Julian Champkin. Alphonse Bertillon and the measure of man: More expert than Sherlock Holmes. Significance (Oxford, England), 11(2):36–39, April 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

DeRose:2014:VEL

[732] Catherine DeRose, Carrie Roy, and Fred Boehm. Victorian eyes: Literary, statistical, and artistic perspectives on Victorian novels — and Dickens’s unfinished murder mystery. Significance (Oxford, England), 11 (2):40–43, April 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:Rb

[733] Anonymous. Reviews. Significance (Oxford, England), 11(2):44, April 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:ELCb

[734] Anonymous. Editorial/letters/crossward. Significance (Oxford, England), 11(2):45–46, April 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2014:JFC

[735] Julian Champkin. Justice is a full cooked breakfast with trimmings. Significance (Oxford, England), 11(2):47–48, April 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:Ca

[736] Anonymous. Contents. Significance (Oxford, England), 11(3):1, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:Nc

[737] Anonymous. News. Significance (Oxford, England), 11(3):2–5, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bigg:2014:IRT

[738] Grant Bigg and Steve Billings. The iceberg risk in the Titanic year of 1912: Was it exceptional? Significance (Oxford, England), 11(3):6–10, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cuthbert:2014:STC

[739] Jim Cuthbert and Margaret Cuthbert. Shaping things to come?: How decisions on statistical disclosure and analysis influenced the debate on Scotland’s independence. Significance (Oxford, England), 11(3):11–13, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Skinner:2014:UPA

[740] Gideon Skinner and Jayesh Navin Shah. Understanding public attitudes to science. Significance (Oxford, England), 11(3):14–18, July 2014. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dean:2014:TOT

[741] Christopher Dean and Tim Cole. The timing of our tooth growth is an evolutionary relic. Significance (Oxford, England), 11(3):19–23, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:DFCc

[742] Anonymous. Dr Fisher’s casebook: Why do we never learn? Significance (Oxford, England), 11(3):24, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2014:WNP

[743] Julian Champkin. “We need the public to become better BS detectors”: Sir Iain Chalmers. Significance (Oxford, England), 11(3):25–30, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Greenfield:2014:SCH

[744] Oliver Greenfield and Paul Allin. Statisticians can help save the world — but they only have 396 days. Significance (Oxford, England), 11(3): 31–34, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kuhn:2014:WAB

[745] Max Kuhn and Kjell Johnson. Who’s afraid of the big black box?: Statisticians’ vital role in big data and predictive modelling. Significance (Oxford, England), 11(3):35–37, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Amoros:2014:RL

[746] Jaume Amorós. Recapturing Laplace. Significance (Oxford, England), 11 (3):38–39, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Champkin:2014:WWE

[747] Julian Champkin. The world will end in 3150: John Craig and the Second Coming. Significance (Oxford, England), 11(3):40–42, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:Rc

[748] Anonymous. Reviews. Significance (Oxford, England), 11(3):43, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:ELCc

[749] Anonymous. Editorial/letters/crossward. Significance (Oxford, England), 11(3):44–45, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jolly:2014:CLS

[750] Tobias Jolly. Could a ‘safest’ option on sat navs save lives? Significance (Oxford, England), 11(3):46–48, July 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:Cb

[751] Anonymous. Content. Significance (Oxford, England), 11(4):1, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:NIEa

[752] Anonymous. News/interview/editorial. Significance (Oxford, England), 11(4):2–7, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:CM

[753] Anonymous. Cultural movements. Significance (Oxford, England), 11 (4):8–9, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Smith:2014:DPN

[754] Sam Smith. Data and privacy: Now you see me; New model for data sharing; Modern governance and staticians. Significance (Oxford, England), 11(4):10–17, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2014:KW

[755] Brian Tarran. Keeping watch. Significance (Oxford, England), 11(4): 18–21, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Auerbach:2014:DNY

[756] Jonathan Auerbach. Does New York City really have as many rats as people? Significance (Oxford, England), 11(4):22–27, October 2014. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reinhart:2014:HP

[757] Alex Reinhart. Huff and puff. Significance (Oxford, England), 11(4): 28–33, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kemeny:2014:SFF

[758] Richard Kemeny. The statistical foundations of the Fourth Estate. Significance (Oxford, England), 11(4):34–35, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Watkins:2014:TWO

[759] Phillip Watkins and David Hand. “Things which ought to be expected can seem quite extraordinary if you’ve got the wrong model”. Significance (Oxford, England), 11(4):36–39, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Milner:2014:HWD

[760] Kate Milner and Jonathan Rougier. How to weigh a donkey in the Kenyan countryside. Significance (Oxford, England), 11(4):40–43, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hartman:2014:SUP

[761] Katherine Hartman, Hermann Habermann, Brian Harris-Kojetin, Christa Jones, Tom Louis, and Andrew Gelman. Strength under pressure/a world without statistics. Significance (Oxford, England), 11(4): 44–47, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:DFCd

[762] Anonymous. Dr Fisher’s casebook: Tools of the trade. Significance (Oxford, England), 11(4):48, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:BR

[763] Anonymous. Book reviews. Significance (Oxford, England), 11(4):49, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:LCa

[764] Anonymous. Letters/ crossword. Significance (Oxford, England), 11(4): 50–51, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:BWa

[765] Anonymous. Best of the web. Significance (Oxford, England), 11(4): 52, October 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:Cc

[766] Anonymous. Contents. Significance (Oxford, England), 11(5):1, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:NIEb

[767] Anonymous. News/interview/editorial. Significance (Oxford, England), 11(5):2–7, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2014:PD

[768] Brian Tarran. Playing with data. Significance (Oxford, England), 11(5): 8–9, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Majumder:2014:FAE

[769] Maimuna (Maia) Majumder. The fight against Ebola. Significance (Oxford, England), 11(5):10–13, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harford:2014:BDB

[770] Tim Harford. Big data: a big mistake? Significance (Oxford, England), 11(5):14–19, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Redelmeier:2014:CBB

[771] Donald Redelmeier and Sharon May. Caution: baby on board. Significance (Oxford, England), 11(5):20–25, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Owen:2014:LE

[772] Nisha Owen. Life on the edge. Significance (Oxford, England), 11(5): 26–29, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hanley:2014:HDO

[773] James Hanley. How deep is the ocean? Significance (Oxford, England), 11(5):30–33, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Swires-Hennessy:2014:CTN

[774] Ed Swires-Hennessy and Allan Reese. Cutting through the numbers. Significance (Oxford, England), 11(5):34–37, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Zelterman:2014:GD

[775] Daniel Zelterman. A groaning demographic. Significance (Oxford, England), 11(5):38–43, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Clifford:2014:SCP

[776] David Clifford, Kim-Anh Lê Cao, and Bevan Emma Huang. A statistician’s Christmas party. Significance (Oxford, England), 11(5):44–47, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hilbe:2014:LUE

[777] Joseph M. Hilbe, Jamie Riggs, Benjamin D. Wandelt, Rafael S. de Souza, Emille E. O. Ishida, Jessi Cisewski, Vladimir Surdin, Madhura Killedar, Roberto Trotta, Bruce Bassett, Yabebal Fantaye, and Chris Impey. Life, the universe, and everything. Significance (Oxford, England), 11(5):48– 75, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McLeod:2014:WMI

[778] Paula McLeod. The world’s most isolated statistician. Significance (Oxford, England), 11(5):76–77, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Marriott:2014:FST

[779] Nigel Marriott. The future of statistical thinking. Significance (Oxford, England), 11(5):78–80, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bendell:2014:SI

[780] Tony Bendell. Stress to impress. Significance (Oxford, England), 11(5): 81, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:DFCe

[781] Anonymous. Dr. Fisher’s casebook. Significance (Oxford, England), 11 (5):82, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:FR

[782] Anonymous. Film review. Significance (Oxford, England), 11(5):83, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:RR

[783] Anonymous. Recommended reading. Significance (Oxford, England), 11 (5):84, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:LCb

[784] Anonymous. Letters/ crossword. Significance (Oxford, England), 11(5): 85–87, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2014:BWb

[785] Anonymous. Best of the web. Significance (Oxford, England), 11(5): 88, December 2014. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:Ca

[786] Anonymous. Contents. Significance (Oxford, England), 12(1):1, February 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bram:2015:NIE

[787] Uri Bram and Brian Tarran. News/interview/editorial. Significance (Oxford, England), 12(1):2–7, February 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2015:HML

[788] Brian Tarran and Zoubin Ghahramani. How machines learned to think statistically. Significance (Oxford, England), 12(1):8–15, February 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wainer:2015:WNW

[789] Howard Wainer and Richard Feinberg. For want of a nail: Why unnecessarily long tests may be impeding the progress of Western civilisation. Significance (Oxford, England), 12(1):16–21, February 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Weisberg:2015:WNR

[790] Herbert I. Weisberg. What next for randomised clinical trials? Significance (Oxford, England), 12(1):22–27, February 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gray:2015:GSS

[791] Chris Gray. Game, set and stats. Significance (Oxford, England), 12(1): 28–31, February 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tseloni:2015:SP

[792] Andromachi Tseloni and Rebecca Thompson. Securing the premises. Significance (Oxford, England), 12(1):32–35, February 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jeong:2015:MB

[793] Yujin Jeong and Joseph L. Gastwirth. A murky business. Significance (Oxford, England), 12(1):36–39, February 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Allin:2015:GMM

[794] Paul Allin and David Hand. Getting the measure of modern life. Significance (Oxford, England), 12(1):40–42, February 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:DFCa

[795] Anonymous. Dr Fisher’s casebook: Still learning. Significance (Oxford, England), 12(1):43, February 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:Ra

[796] Anonymous. Reviews. Significance (Oxford, England), 12(1):44–45, February 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:LCa

[797] Anonymous. Letters/ crossword. Significance (Oxford, England), 12(1): 46–47, February 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:BWa

[798] Anonymous. Best of the web. Significance (Oxford, England), 12(1): 48, February 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:Cb

[799] Anonymous. Contents. Significance (Oxford, England), 12(2):1, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2015:NIEa

[800] Brian Tarran and Michael Wininger. News/interview/editorial. Significance (Oxford, England), 12(2):2–7, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Evans:2015:PPC

[801] Geoffrey Evans and Jon Mellon. The political popularity contest. Significance (Oxford, England), 12(2):8–10, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Baio:2015:PO

[802] Gianluca Baio and Roberto Cerina. A predictable outcome. Significance (Oxford, England), 12(2):11–13, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Price:2015:DWU

[803] Megan Price, Anita Gohdes, and Patrick Ball. Documents of war: Understanding the Syrian conflict. Significance (Oxford, England), 12(2):

14–19, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Stride:2015:AHH

[804] Chris Stride and Ffion Thomas. The Apprentice: How to be hired, not fired. Significance (Oxford, England), 12(2):20–24, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Luo:2015:NRF

[805] Tian Luo and Philip B. Stark. Nine out of 10 restaurants fail? Check, please. Significance (Oxford, England), 12(2):25–29, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dagpunar:2015:DSP

[806] John Dagpunar. Deferring a state pension — is it worthwhile? Significance (Oxford, England), 12(2):30–35, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Peacock:2015:L

[807] Jack Peacock, Jeshua Maxey, and Vijay Tymms. Left out? Significance (Oxford, England), 12(2):36–39, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lucas:2015:CWL

[808] Joseph E. Lucas and Guillermo Sapiro. Cancer: What’s luck got to do with it? Significance (Oxford, England), 12(2):40–42, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:DFa

[809] Anonymous. Dr. Fisher. Significance (Oxford, England), 12(2):43, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:Rb

[810] Anonymous. Reviews. Significance (Oxford, England), 12(2):44–45, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:LCb

[811] Anonymous. Letters/ crossword. Significance (Oxford, England), 12(2): 46–47, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:BWb

[812] Anonymous. Best of the web. Significance (Oxford, England), 12(2): 48, April 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:Cc

[813] Anonymous. Contents. Significance (Oxford, England), 12(3):1, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2015:NIEb

[814] Brian Tarran and Rick Wicklin. News/interview/editorial. Significance (Oxford, England), 12(3):2–9, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hill:2015:FEU

[815] Timothy Martyn Hill. Forecast error: The UK general election. Significance (Oxford, England), 12(3):10–15, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bales:2015:MSU

[816] Kevin Bales, Olivia Hesketh, and Bernard Silverman. Modern slavery in the UK: How many victims? Significance (Oxford, England), 12(3): 16–21, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Woodard:2015:FHF

[817] Bradford Woodard. Fighting healthcare fraud with statistics. Significance (Oxford, England), 12(3):22–25, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Puts:2015:FEB

[818] Marco Puts, Piet Daas, and Ton de Waal. Finding errors in Big Data. Significance (Oxford, England), 12(3):26–29, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Peng:2015:RCS

[819] Roger Peng. The reproducibility crisis in science: a statistical counterattack. Significance (Oxford, England), 12(3):30–32, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gelman:2015:WTS

[820] Andrew Gelman. Working through some issues. Significance (Oxford, England), 12(3):33–35, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:PSC

[821] Anonymous. Psychology and statistics, continued …: Readers respond to the BASP ban on p-values. Significance (Oxford, England), 12(3): 35–36, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gore:2015:GOA

[822] Katrina Gore and Phil Stanley. Getting our ACT together. Significance (Oxford, England), 12(3):36–39, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Grajales:2015:PC

[823] Carlos A. Gómez Grajales. Not part of the crowd. Significance (Oxford, England), 12(3):40–42, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:DFb

[824] Anonymous. Dr. Fisher. Significance (Oxford, England), 12(3):43, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:Rc

[825] Anonymous. Reviews. Significance (Oxford, England), 12(3):44–45, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:LCc

[826] Anonymous. Letters/ crossword. Significance (Oxford, England), 12(3): 46–47, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:BWc

[827] Anonymous. Best of the web. Significance (Oxford, England), 12(3):48, June 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:Cd

[828] Anonymous. Contents. Significance (Oxford, England), 12(4):1, August 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2015:NIEc

[829] Brian Tarran and John Bohannon. News/interview/editorial. Significance (Oxford, England), 12(4):2–9, August 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Brent:2015:ST

[830] Eric S. Brent. The statistics of tragedy. Significance (Oxford, England), 12(4):10–15, August 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Skorupski:2015:BFC

[831] William P. Skorupski and Howard Wainer. The Bayesian flip: Correcting the prosecutor’s fallacy. Significance (Oxford, England), 12(4):16–20, August 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spiegelhalter:2015:SRS

[832] David Spiegelhalter. Sex-rated statistics. Significance (Oxford, England), 12(4):21–25, August 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Marriott:2015:VDR

[833] Nigel Marriott. A very different result. Significance (Oxford, England), 12(4):26–29, August 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Greenberg:2015:RCC

[834] Adam Greenberg. The red card cliché. Significance (Oxford, England), 12(4):30–33, August 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Simpson:2015:BSI

[835] Sean L. Simpson, Jonathan H. Burdette, and Paul J. Laurienti. The brain science interface. Significance (Oxford, England), 12(4):34–39, August 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Norton:2015:SPL

[836] H. James Norton and George Divine. Simpson’s paradox … and how to avoid it. Significance (Oxford, England), 12(4):40–43, August 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:DFCb

[837] Anonymous. Dr Fisher’s casebook: “More or less” is about right. Significance (Oxford, England), 12(4):44, August 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:Rd

[838] Anonymous. Reviews. Significance (Oxford, England), 12(4):45, August 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:LCd

[839] Anonymous. Letters/ crossword. Significance (Oxford, England), 12(4): 46–47, August 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:BWd

[840] Anonymous. Best of the web. Significance (Oxford, England), 12(4): 48, August 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:Ce

[841] Anonymous. Contents. Significance (Oxford, England), 12(5):1, October 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Christensen:2015:NIE

[842] Hannah Christensen and Brian Tarran. News/interview/editorial. Significance (Oxford, England), 12(5):2–7, October 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Skeffington:2015:WBO

[843] James Skeffington. Warren Buffett: Oracle or orang-utan? Significance (Oxford, England), 12(5):8–11, October 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kovalchik:2015:GSS

[844] Stephanie Kovalchik. Grand Slams are short-changing women’s tennis. Significance (Oxford, England), 12(5):12–17, October 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Goldstein:2015:JWC

[845] Harvey Goldstein. Jumping to the wrong conclusions. Significance (Oxford, England), 12(5):18–21, October 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Pullinger:2015:FGD

[846] John Pullinger, Claire Melamed, George Bentley, Julie Shortridge, Stefanie Falconi, Ben Zaitchik, Seth Guikema, Jacqueline Joudo Larsen, Monti Narayan Datta, Kevin Bales, Delphine Robineau, Matthew Tranter, Nili Karsai, and Helen Brusilovsky. Fuel for growth and development. Significance (Oxford, England), 12(5):22–43, October 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:DFCc

[847] Anonymous. Dr Fisher’s casebook: Stereotypically yours. Significance (Oxford, England), 12(5):44, October 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:Re

[848] Anonymous. Reviews. Significance (Oxford, England), 12(5):45, October 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:LCe

[849] Anonymous. Letters/ crossword. Significance (Oxford, England), 12(5): 46–47, October 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:BWe

[850] Anonymous. Best of the web. Significance (Oxford, England), 12(5): 48, October 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:Cf

[851] Anonymous. Contents. Significance (Oxford, England), 12(6):1, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Olson:2015:NIE

[852] Randy Olson and Brian Tarran. News/interview/editorial. Significance (Oxford, England), 12(6):2–11, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Krall:2015:VSD

[853] Jenna R. Krall and Roger D. Peng. The Volkswagen scandal: Deception, driving and deaths. Significance (Oxford, England), 12(6):12–15, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2015:IUS

[854] Brian Tarran. “The idea of using statistics to think about individuals is quite strange”. Significance (Oxford, England), 12(6):16–19, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dempsey:2015:RTF

[855] Walter Dempsey, Peng Liao, Pedja Klasnja, Inbal Nahum-Shani, and Susan A. Murphy. Randomised trials for the Fitbit generation. Significance (Oxford, England), 12(6):20–23, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Haynes:2015:FGC

[856] Kaylea Haynes. The Fitbit generation: From couch potato to marathon runner. Significance (Oxford, England), 12(6):24–25, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Majumder:2015:ELH

[857] Maimuna (Maia) Majumder. Ebola: Lessons from a human tragedy. Significance (Oxford, England), 12(6):26–29, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wininger:2015:SRS

[858] Michael Wininger. A stopping rule for scholastic American football? Significance (Oxford, England), 12(6):30–33, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Zelin:2015:MCE

[859] Andrew Zelin. Modelling calls and effects. Significance (Oxford, England), 12(6):34–39, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bland:2015:PP

[860] Martin Bland. Promiscuous publication. Significance (Oxford, England), 12(6):40–41, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Stigler:2015:PEN

[861] Stephen M. Stigler. Is probability easier now than in 1560? Significance (Oxford, England), 12(6):42–43, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:DFCd

[862] Anonymous. Dr Fisher’s casebook: the kind of pedantry up with which I will put. Significance (Oxford, England), 12(6):44, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:Rf

[863] Anonymous. Reviews. Significance (Oxford, England), 12(6):45, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:LCf

[864] Anonymous. Letters/ crossword. Significance (Oxford, England), 12(6): 46–47, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2015:BWf

[865] Anonymous. Best of the web. Significance (Oxford, England), 12(6): 48, December 2015. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:Ea

[866] Brian Tarran. Editorial. Significance (Oxford, England), 13(1):1, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Ca

[867] Anonymous. Contents. Significance (Oxford, England), 13(1):2–3, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kennedy:2016:NI

[868] Helen Kennedy, Andy Kirk, Julian Stander, Luciana Dalla Valle, and Mario Cortina Borja. News/ interview. Significance (Oxford, England), 13(1):4–15, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bain:2016:CSS

[869] Robert Bain. Citizen science and statistics: Playing a part. Significance (Oxford, England), 13(1):16–21, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wainer:2016:DTH

[870] Howard Wainer. Don’t try this at home. Significance (Oxford, England), 13(1):22–23, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Orzack:2016:ONI

[871] Steven Hecht Orzack. Old and new ideas about the human sex ratio. Significance (Oxford, England), 13(1):24–27, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mais:2016:MWE

[872] David Mais, Daryl Lloyd, and Jennifer Davies. Modelling weather effects on road casualty statistics. Significance (Oxford, England), 13(1):28– 31, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Somerville:2016:BAP

[873] Andrew Somerville. A Bayesian analysis of peer reviewing. Significance (Oxford, England), 13(1):32–37, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Crede:2016:GFP

[874] Carsten Crede. Getting a fix on price-fixing cartels. Significance (Oxford, England), 13(1):38–41, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dalal:2016:RR

[875] Siddartha Dalal. From risk to resiliency. Significance (Oxford, England), 13(1):42–43, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:DFa

[876] Anonymous. Dr. Fisher. Significance (Oxford, England), 13(1):44, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Ra

[877] Anonymous. Reviews. Significance (Oxford, England), 13(1):45, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:LCa

[878] Anonymous. Letters/ crossword. Significance (Oxford, England), 13(1): 46–47, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Althouse:2016:CW

[879] Andrew D. Althouse. In closing/on the web. Significance (Oxford, England), 13(1):48, February 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:Eb

[880] Brian Tarran. Editorial. Significance (Oxford, England), 13(2):1, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Cb

[881] Anonymous. Contents. Significance (Oxford, England), 13(2):2–3, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:Na

[882] Brian Tarran. News. Significance (Oxford, England), 13(2):4–5, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:EBV

[883] Brian Tarran. The economy: a Brexit vote winner? Significance (Oxford, England), 13(2):6–7, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Golbeck:2016:WUA

[884] Amanda Golbeck. Are women underrepresented in the American statistics profession? Significance (Oxford, England), 13(2):8–9, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Woo:2016:EU

[885] Gordon Woo. Expect the unexpected. Significance (Oxford, England), 13 (2):10–11, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Timaeus:2016:WWH

[886] Ian Timaeus. What would happen if UK residents stopped having babies? Significance (Oxford, England), 13(2):12–13, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Southorn:2016:GEP

[887] Graham Southorn. Great expectations: the past, present and future of prediction. Significance (Oxford, England), 13(2):14–19, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Meyer:2016:GWS

[888] Renate Meyer and Nelson Christensen. Gravitational waves: a statistical autopsy of a black hole merger. Significance (Oxford, England), 13(2): 20–25, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Norton:2016:FFS

[889] Jim Norton, William E. Anderson, and George Divine. Flawed forensics: Statistical failings of microscopic hair analysis. Significance (Oxford, England), 13(2):26–29, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kiefer:2016:SCR

[890] Daniel Kiefer, Jerome Sacks, and Donald Ylvisaker. Statistics, civil rights and the US Supreme Court: a cautionary tale. Significance (Oxford, England), 13(2):30–33, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Buvinic:2016:CGD

[891] Mayra Buvinic and Ruth Levine. Closing the gender data gap. Significance (Oxford, England), 13(2):34–37, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Smith:2016:CPC

[892] Kim Bradford Smith. Counting people in a conflict zone. Significance (Oxford, England), 13(2):38–41, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:HP

[893] Brian Tarran. Historical perspective. Significance (Oxford, England), 13 (2):42, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:DFb

[894] Anonymous. Dr. Fisher. Significance (Oxford, England), 13(2):43, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cirillo:2016:WCW

[895] Pasquale Cirillo and Nassim Nicholas Taleb. What are the chances of war? Significance (Oxford, England), 13(2):44–45, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:LCb

[896] Anonymous. Letters/ crossword. Significance (Oxford, England), 13(2): 46–47, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Titterington:2016:RPG

[897] Mike Titterington. Remembering Peter Gavin Hall. Significance (Oxford, England), 13(2):48, April 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:Ec

[898] Brian Tarran. Editorial. Significance (Oxford, England), 13(3):1, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Cc

[899] Anonymous. Contents. Significance (Oxford, England), 13(3):2–3, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:Nb

[900] Brian Tarran. News. Significance (Oxford, England), 13(3):4–5, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bakhshi:2016:HCW

[901] Hasan Bakhshi. How can we measure the modern digital economy? Significance (Oxford, England), 13(3):6–7, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hemenway:2016:NED

[902] Brett Hemenway. A new era for data sharing? Significance (Oxford, England), 13(3):8–9, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Taylor:2016:LBA

[903] Benjamin Taylor. London’s burning: Assessing the impact of the 2014 fire station closures. Significance (Oxford, England), 13(3):10–11, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Breckon:2016:EBE

[904] Jonathan Breckon. An evidence base for evidence-informed policy. Significance (Oxford, England), 13(3):12–13, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hassad:2016:ZVB

[905] Rossi A. Hassad. Zika virus and birth defects: correlation or causation? Significance (Oxford, England), 13(3):14–15, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Paulden:2016:SR

[906] Dr Tim Paulden. Smashing the racket. Significance (Oxford, England), 13(3):16–21, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bredtmann:2016:OMD

[907] Julia Bredtmann, Carsten J. Crede, and Sebastian Otten. Olympic medals: Does the past predict the future? Significance (Oxford, England), 13(3):22–25, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gandy:2016:SSS

[908] Rob Gandy. Second season syndrome. Significance (Oxford, England), 13 (3):26–29, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Matthews:2016:BD

[909] Robert Matthews. Beautiful, but dangerous. Significance (Oxford, England), 13(3):30–31, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Adamou:2016:DI

[910] Alexander Adamou and Ole Peters. Dynamics of inequality. Significance (Oxford, England), 13(3):32–35, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dorling:2016:DIR

[911] Danny Dorling. Dynamics of inequality: In response. Significance (Oxford, England), 13(3):36–37, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Goodman:2016:PPB

[912] William Goodman. The promises and pitfalls of Benford’s law. Significance (Oxford, England), 13(3):38–41, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:TSS

[913] Anonymous. Teaching statistics with a sci-fi twist. Significance (Oxford, England), 13(3):42, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:DFCa

[914] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 13 (3):43, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:LPa

[915] Anonymous. Letters/ puzzle. Significance (Oxford, England), 13(3): 44–46, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Rb

[916] Anonymous. Reviews. Significance (Oxford, England), 13(3):47, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Wa

[917] Anonymous. On the web. Significance (Oxford, England), 13(3):48, June 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:Ed

[918] Brian Tarran. Editorial. Significance (Oxford, England), 13(4):1, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Cd

[919] Anonymous. Contents. Significance (Oxford, England), 13(4):2–3, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:Nc

[920] Brian Tarran. News. Significance (Oxford, England), 13(4):4–5, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:CCG

[921] Brian Tarran. California commits $5m for gun violence research. Significance (Oxford, England), 13(4):6–7, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:HMM

[922] Brian Tarran. How much marijuana goes up in smoke? Significance (Oxford, England), 13(4):7, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Borja:2016:BSD

[923] Mario Cortina Borja, Julian Stander, and Luciana Dalla Valle. Brexit: surname diversity and voting patterns. Significance (Oxford, England), 13(4):8–9, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Collett:2016:WSD

[924] Laura Collett. Were self-driving cars really about to kill someone? Significance (Oxford, England), 13(4):10, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Toher:2016:SSW

[925] Deirdre Toher. Scrambled statistics: What are the chances of finding multi-yolk eggs? Significance (Oxford, England), 13(4):11, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Matthews:2016:WCW

[926] Robert Matthews. Why can we not predict earthquakes? Significance (Oxford, England), 13(4):12–13, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bain:2016:OBB

[927] Robert Bain. Are our brains Bayesian? Significance (Oxford, England), 13(4):14–19, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kobak:2016:SFE

[928] Dmitry Kobak, Sergey Shpilkin, and Maxim S. Pshenichnikov. Statistical fingerprints of electoral fraud? Significance (Oxford, England), 13(4): 20–23, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Skorupski:2016:BCS

[929] William P. Skorupski and Howard Wainer. Breast cancer screenings: Does the evidence support the recommendations? Significance (Oxford, England), 13(4):24–27, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mesquita:2016:CCD

[930] Michel d. S. Mesquita and Kjell Einar Erikstad. Cold case: the death of common guillemots in the Barents Sea. Significance (Oxford, England), 13(4):28–33, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Chase:2016:AGS

[931] Harrison W. Chase and Mark E. Glickman. The analytics of getting sacked: Coach firings in the National Football League. Significance (Oxford, England), 13(4):34–37, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Baayen:2016:DDW

[932] Corine Baayen. The drugs don’t work … or do they? Significance (Oxford, England), 13(4):38–41, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harkness:2016:MM

[933] Timandra Harkness. Making maps. Significance (Oxford, England), 13 (4):42–43, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:DFCb

[934] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 13 (4):44, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Rc

[935] Anonymous. Reviews. Significance (Oxford, England), 13(4):45, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:LPb

[936] Anonymous. Letters/ puzzle. Significance (Oxford, England), 13(4):46– 47, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Wb

[937] Anonymous. On the web. Significance (Oxford, England), 13(4):48, August 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:FP

[938] Brian Tarran. A failure of prediction? Significance (Oxford, England), 13(5):1, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Ce

[939] Anonymous. Contents. Significance (Oxford, England), 13(5):2–3, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:Nd

[940] Brian Tarran. News. Significance (Oxford, England), 13(5):4–5, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hilgard:2016:VGV

[941] Joseph Hilgard. Video game violence and aggression: a proven connection? Significance (Oxford, England), 13(5):6–7, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lau:2016:EFL

[942] F. Din-Houn Lau and Axel Gandy. Enhancing football league tables. Significance (Oxford, England), 13(5):8–9, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rogers:2016:LW

[943] Dr Jennifer Rogers. Luck of the withdrawal. Significance (Oxford, England), 13(5):10–11, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:CCP

[944] Brian Tarran. Climate change in pictures. Significance (Oxford, England), 13(5):12, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gorard:2016:DMO

[945] Stephen Gorard. Does missing one week of school lead to lower grades? Significance (Oxford, England), 13(5):13, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lum:2016:PS

[946] Kristian Lum and William Isaac. To predict and serve? Significance (Oxford, England), 13(5):14–19, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Didelot:2016:HST

[947] Xavier Didelot. Heroic sacrifice or tragic mistake? Revisiting the Eyam plague, 350 years on. Significance (Oxford, England), 13(5):20–25, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kashlak:2016:FAA

[948] Adam B. Kashlak. The frequency of “America” in America. Significance (Oxford, England), 13(5):26–29, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Barfar:2016:EW

[949] Arash Barfar and Balaji Padmanabhan. Election watch. Significance (Oxford, England), 13(5):30–33, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wallbank:2016:ST

[950] Caroline Wallbank. Safer travels. Significance (Oxford, England), 13(5): 34–37, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Winchester:2016:SPH

[951] Niven Winchester. Scoring points: How statistics helped change a rugby competition. Significance (Oxford, England), 13(5):38–41, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:BCB

[952] Brian Tarran. The big C, big data and statistics. Significance (Oxford, England), 13(5):42–43, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:DFCc

[953] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 13 (5):44, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Rd

[954] Anonymous. Reviews. Significance (Oxford, England), 13(5):45, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:LPc

[955] Anonymous. Letters/ puzzle. Significance (Oxford, England), 13(5):46– 47, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Wc

[956] Anonymous. On the web. Significance (Oxford, England), 13(5):48, October 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:WDW

[957] Brian Tarran. Where do we go from here? Significance (Oxford, England), 13(6):1, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Cf

[958] Anonymous. Contents. Significance (Oxford, England), 13(6):2–3, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:Ne

[959] Brian Tarran. News. Significance (Oxford, England), 13(6):4–5, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Finazzi:2016:HSN

[960] Francesco Finazzi. How a smartphone network detects earthquakes in real time. Significance (Oxford, England), 13(6):6–7, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Angelini:2016:YSN

[961] Viola Angelini, Marco Bertoni, and Luca Corazzini. Are you satisfied now? Significance (Oxford, England), 13(6):8–9, December 2016. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jones:2016:MSP

[962] Corey Jones, Kathleen Campbell Garwood, and Alicia Graziosi Strandberg. A major selection problem. Significance (Oxford, England), 13(6): 10–11, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Borja:2016:VBP

[963] Mario Cortina Borja. A variation of the birthday problem. Significance (Oxford, England), 13(6):12–13, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nguyen:2016:SG

[964] Thanh Nguyen. Serious games. Significance (Oxford, England), 13(6): 14–19, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Schwarz:2016:ASN

[965] Alan Schwarz. ADHD: the statistics of a “national disaster”. Significance (Oxford, England), 13(6):20–23, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Purdam:2016:HMD

[966] Dr Kingsley Purdam. How many dead people vote? Significance (Oxford, England), 13(6):24–27, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wilkerson:2016:HD

[967] Rachel Lynne Wilkerson. Hungry for data. Significance (Oxford, England), 13(6):28–31, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ryan:2016:BCN

[968] Louise M. Ryan, Matt P. Wand, and Alan A. Malecki. Bringing coals to Newcastle. Significance (Oxford, England), 13(6):32–37, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Saltelli:2016:YSY

[969] Andrea Saltelli. Young statistician, you shall live in adventurous times. Significance (Oxford, England), 13(6):38–41, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2016:MP

[970] Brian Tarran. Math panic. Significance (Oxford, England), 13(6):42– 43, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:DFCd

[971] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 13 (6):44, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:BR

[972] Anonymous. Book Reviews. Significance (Oxford, England), 13(6):45, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:LCc

[973] Anonymous. Letters & crossword. Significance (Oxford, England), 13(6): 46–47, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2016:Wd

[974] Anonymous. On the web. Significance (Oxford, England), 13(6):48, December 2016. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:BS

[975] Brian Tarran. Behind the scenes. Significance (Oxford, England), 14 (1):1, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Ca

[976] Anonymous. Contents. Significance (Oxford, England), 14(1):2–3, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:N

[977] Brian Tarran. News. Significance (Oxford, England), 14(1):4–5, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Borja:2017:WMP

[978] Mario Cortina Borja and Peter Martin. What is the most popular birthday in England and Wales? Significance (Oxford, England), 14(1):6–7, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Limpert:2017:LND

[979] Eckhard Limpert and Werner A. Stahel. The log-normal distribution. Significance (Oxford, England), 14(1):8–9, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:WCN

[980] Anonymous. 2017 writing competition now open for entries. Significance (Oxford, England), 14(1):9, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:DHB

[981] Brian Tarran. Do hungry birds of a feather flock together? Significance (Oxford, England), 14(1):10–11, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:HWW

[982] Brian Tarran. How wrong we are. Significance (Oxford, England), 14(1): 12–13, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bain:2017:HBS

[983] Robert Bain. Hollywood’s best supporting actors. Significance (Oxford, England), 14(1):14–17, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lindeman:2017:EBE

[984] Mark Lindeman. Evidence-based elections: Beyond the “rigging” debate. Significance (Oxford, England), 14(1):18–23, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Smerdon:2017:WWE

[985] Jason E. Smerdon. What was Earth’s climate like before we were measuring it? Significance (Oxford, England), 14(1):24–29, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Eklund:2017:HOS

[986] Anders Eklund and Thomas Nichols. How open science revealed false positives in brain imaging. Significance (Oxford, England), 14(1):30– 33, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

deMoira:2017:SS

[987] Anne Pinot de Moira and Kate Kelly. Strictly statistics. Significance (Oxford, England), 14(1):34–37, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Raper:2017:WGS

[988] Simon Raper. Why good science is good business. Significance (Oxford, England), 14(1):38–41, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:HMS

[989] Brian Tarran. How to measure statistical literacy? Significance (Oxford, England), 14(1):42–43, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:DFCa

[990] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 14 (1):44, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Ra

[991] Anonymous. In Reviews. Significance (Oxford, England), 14(1):45, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Rb

[992] Anonymous. In response. Significance (Oxford, England), 14(1):46–47, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Wa

[993] Anonymous. On the web. Significance (Oxford, England), 14(1):48, February 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:TH

[994] Brian Tarran. Tragedy and hope. Significance (Oxford, England), 14 (2):1, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Cb

[995] Anonymous. Contents. Significance (Oxford, England), 14(2):2–3, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:CCH

[996] Brian Tarran. Counting crowds is hard to do. Significance (Oxford, England), 14(2):4–5, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:HWS

[997] Brian Tarran. “Hans was a storyteller; he turned statistics into a performing art”. Significance (Oxford, England), 14(2):6–7, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Olhede:2017:FTA

[998] Sofia Olhede and Russell Rodrigues. Fairness and transparency in the age of the algorithm. Significance (Oxford, England), 14(2):8–9, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Selby:2017:CCE

[999] David Selby. Cloudy with a chance of engagement. Significance (Oxford, England), 14(2):10–11, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Eastoe:2017:EVD

[1000] Emma Eastoe. Extreme value distributions. Significance (Oxford, England), 14(2):12–13, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kashlak:2017:WTS

[1001] Adam B. Kashlak. Was Trump’s speech presidential? Significance (Oxford, England), 14(2):14–15, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Langkjaer-Bain:2017:MTF

[1002] Robert Langkjær-Bain. The murky tale of Flint’s deceptive water data. Significance (Oxford, England), 14(2):16–21, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cisewski:2017:SEA

[1003] Jessi Cisewski. In search of Earth analogues: Detecting exoplanets amid stellar noise. Significance (Oxford, England), 14(2):22–25, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Towers:2017:FET

[1004] Sherry Towers and Michael D. White. The “Ferguson effect”, or too many guns?: Exploring violent crime in Chicago. Significance (Oxford, England), 14(2):26–29, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tirlea:2017:TPS

[1005] Marius A. Tirlea, Shahin Tavakoli, Davide Pigoli, and John A. D. Aston. A ticklish problem: Studying accent variation from speech recordings. Significance (Oxford, England), 14(2):30–33, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Warner:2017:DPD

[1006] Jamie Warner, Olga Dobromyrova, Dana Shay, and Qing Wang. Deeper pockets, deeper pleasure? Significance (Oxford, England), 14(2):34–37, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Matthews:2017:APV

[1007] Robert Matthews, Ron Wasserstein, and David Spiegelhalter. The ASA’s p-value statement, one year on. Significance (Oxford, England), 14(2): 38–41, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Golbeck:2017:HMU

[1008] Amanda L. Golbeck. How many US doctorates go to non-US citizens? Significance (Oxford, England), 14(2):42–43, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:DFCb

[1009] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 14 (2):44, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Rc

[1010] Anonymous. In review. Significance (Oxford, England), 14(2):45, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Rd

[1011] Anonymous. In response. Significance (Oxford, England), 14(2):46–47, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Wb

[1012] Anonymous. On the web. Significance (Oxford, England), 14(2):48, April 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:LH

[1013] Brian Tarran. Lessons from history. Significance (Oxford, England), 14 (3):1, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Cc

[1014] Anonymous. Contents. Significance (Oxford, England), 14(3):2–3, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:USCa

[1015] Brian Tarran. UK statisticians call for end to pre-release access to official statistics. Significance (Oxford, England), 14(3):4–5, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Olhede:2017:CAM

[1016] Sofia Olhede and Russell Rodrigues. The computer ate my personality. Significance (Oxford, England), 14(3):6–7, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bland:2017:LEF

[1017] Martin Bland. Has life expectancy fallen? Significance (Oxford, England), 14(3):8–9, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Eckel:2017:UDP

[1018] Sandrah Eckel. The ups and downs of pregnancy. Significance (Oxford, England), 14(3):10–11, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:SIS

[1019] Brian Tarran. Should international students be classed as “long-term migrants”? Significance (Oxford, England), 14(3):12–13, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

MacKay:2017:DBW

[1020] Niall MacKay, Chris Price, and Jamie Wood. Dogger Bank: Weighing the fog of war. Significance (Oxford, England), 14(3):14–19, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fong:2017:BBD

[1021] Wai Mun Fong. Bursting the big data bubble. Significance (Oxford, England), 14(3):20–23, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:LWM

[1022] Brian Tarran. Lord Woolton: the man who used statistics (and more) to feed a nation at war. Significance (Oxford, England), 14(3):24–29, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hohle:2017:RUW

[1023] Michael Höhle. Rank uncertainty: Why the “most popular” baby names might not be the most popular. Significance (Oxford, England), 14(3): 30–33, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Salsburg:2017:HFI

[1024] David Salsburg. How far is it from the Earth to the Sun? Significance (Oxford, England), 14(3):34–37, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Pan:2017:HPS

[1025] Jason Pan, Kelly H. Zou, Ching-Ray Yu, Franklin W. Sun, and Martin O. Carlsson. How to plan and schedule conferences with the help of artificial intelligence. Significance (Oxford, England), 14(3):38–41, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:SSN

[1026] Brian Tarran. “The social sciences need to build new foundations”. Significance (Oxford, England), 14(3):42–43, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:DFCc

[1027] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 14 (3):44, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Re

[1028] Anonymous. In review. Significance (Oxford, England), 14(3):45, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Rf

[1029] Anonymous. In response. Significance (Oxford, England), 14(3):46–47, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Wc

[1030] Anonymous. On the web. Significance (Oxford, England), 14(3):48, June 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:STA

[1031] Brian Tarran. Something to talk about. Significance (Oxford, England), 14(4):1, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Cd

[1032] Anonymous. Contents. Significance (Oxford, England), 14(4):2–3, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:EWS

[1033] Brian Tarran. An early warning system for the A&E winter crisis? Significance (Oxford, England), 14(4):4–5, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:TCT

[1034] Brian Tarran. “It’s time to be concerned. It’s not time to panic”. Significance (Oxford, England), 14(4):6–7, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lusinchi:2017:CDT

[1035] Dominic Lusinchi. “Clinton defeats Trump” 2016 polls and the shadow of 1948. Significance (Oxford, England), 14(4):8–9, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Newman:2017:PLD

[1036] Mark Newman. Power-law distribution. Significance (Oxford, England), 14(4):10–11, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Toher:2017:HMS

[1037] Deirdre Toher. Help! Is my sample big enough? Significance (Oxford, England), 14(4):12–13, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Langkjaer-Bain:2017:TS

[1038] Robert Langkjær-Bain. Trials of a statistician. Significance (Oxford, England), 14(4):14–19, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Auerbach:2017:NYC

[1039] Jonathan Auerbach. Are New York City drivers more likely to get a ticket at the end of the month? Significance (Oxford, England), 14(4): 20–25, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Curtice:2017:SSA

[1040] John Curtice, Stephen Fisher, Jouni Kuha, and Jonathan Mellon. Surprise, surprise! (again) The 2017 British general election exit poll. Significance (Oxford, England), 14(4):26–29, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nouvellet:2017:AMU

[1041] Pierre Nouvellet. Age matters in the UK’s Brexit referendum. Significance (Oxford, England), 14(4):30–33, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Purdam:2017:MPP

[1042] Kingsley Purdam, Sam Royston, and Graham Whitham. Measuring the “poverty penalty” in the UK. Significance (Oxford, England), 14(4): 34–37, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gunter:2017:WOA

[1043] Bert Gunter and Christopher Tong. What are the odds!? The “airport fallacy” and statistical inference. Significance (Oxford, England), 14(4): 38–41, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2017:GIDa

[1044] R. Allan Reese. Graphical interpretations of data: an introduction. Significance (Oxford, England), 14(4):42–43, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:DFCd

[1045] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 14 (4):44, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Rg

[1046] Anonymous. In review. Significance (Oxford, England), 14(4):45, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Rh

[1047] Anonymous. In response. Significance (Oxford, England), 14(4):46–47, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Wd

[1048] Anonymous. On the web. Significance (Oxford, England), 14(4):48, August 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:PCR

[1049] Brian Tarran. Post-conference reflections. Significance (Oxford, England), 14(5):1, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Ce

[1050] Anonymous. Contents. Significance (Oxford, England), 14(5):2–3, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:RSP

[1051] Brian Tarran. On the road to success with ParalympicsGB. Significance (Oxford, England), 14(5):4–5, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:WFE

[1052] Brian Tarran. Why facts are not enough in the fight against fake news. Significance (Oxford, England), 14(5):6–7, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Treasure:2017:CMC

[1053] Peter Treasure. Cancer: Misfortune or carelessness? Significance (Oxford, England), 14(5):8–9, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Olhede:2017:WDC

[1054] Sofia Olhede and Russell Rodrigues. Why data is not a commodity. Significance (Oxford, England), 14(5):10–11, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rogers:2017:HRR

[1055] Dr Jennifer Rogers. How “random” is Ryanair’s seating allocation? Significance (Oxford, England), 14(5):12–13, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Andrews:2017:GMG

[1056] Rj. Andrews and Howard Wainer. The Great Migration: a graphics novel. Significance (Oxford, England), 14(5):14–19, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spence:2017:WBP

[1057] Ian Spence, Colin R. Fenn, and Scott Klein. Who is buried in Playfair’s grave? Significance (Oxford, England), 14(5):20–23, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McAdam:2017:POH

[1058] Ronan McAdam. Plastic in the ocean: How much is out there? Significance (Oxford, England), 14(5):24–27, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lin:2017:WMS

[1059] Kevin Lin. We, the millennials The statistical significance of political significance. Significance (Oxford, England), 14(5):28–33, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cole:2017:RLF

[1060] Tim Cole. The remarkable life of Frances Wood. Significance (Oxford, England), 14(5):34–37, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:WSB

[1061] Brian Tarran. “Winning the Super Bowl would be nice”. Significance (Oxford, England), 14(5):38–41, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2017:GIDb

[1062] R. Allan Reese. Graphical interpretations of data: Walking the line. Significance (Oxford, England), 14(5):42–43, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:DFCe

[1063] Anonymous. Dr Fisher’s casebook: Initially. Significance (Oxford, England), 14(5):44, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Shah:2017:PTP

[1064] Hetan Shah. Post-truth publications. Significance (Oxford, England), 14 (5):45, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Ri

[1065] Anonymous. In response. Significance (Oxford, England), 14(5):46–47, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:We

[1066] Anonymous. On the Web. Significance (Oxford, England), 14(5):48, October 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:WT

[1067] Brian Tarran. Winter tales. Significance (Oxford, England), 14(6):1, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Cf

[1068] Anonymous. Contents. Significance (Oxford, England), 14(6):2–3, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2017:USCb

[1069] Brian Tarran. US Secretary of Commerce makes case for extra Census funds. Significance (Oxford, England), 14(6):4–5, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Winter:2017:MSM

[1070] Hugo Winter. Making sense of multiple extreme weather events. Significance (Oxford, England), 14(6):6–7, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Olhede:2017:WAG

[1071] Sofia Olhede and Patrick Wolfe. When algorithms go wrong, who is liable? Significance (Oxford, England), 14(6):8–9, December 2017. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Han:2017:PED

[1072] Gang Han, Brandon Zhao, Kendall Pye, and Hongwei Zhao. The piecewise exponential distribution. Significance (Oxford, England), 14(6):10– 11, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Raper:2017:SM

[1073] Simon Raper. The shock of the mean. Significance (Oxford, England), 14(6):12–17, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Sleeman:2017:CCD

[1074] Richard Sleeman, Colin Aitken, and Amy Wilson. Cocaine on cash: Does quantity vary by location? Significance (Oxford, England), 14(6): 18–23, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lovell:2017:WE

[1075] Christopher Lovell. Where is everybody? Significance (Oxford, England), 14(6):24–27, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hill:2017:HPF

[1076] Edward Hill, Michael Tildesley, and Thomas House. How predictable are flu pandemics? Significance (Oxford, England), 14(6):28–33, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wilmott:2017:BMB

[1077] Paul Wilmott and David Orrell. Big money’s big mistake. Significance (Oxford, England), 14(6):34–37, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Golbeck:2017:HOW

[1078] Amanda Golbeck. How one woman used regression to influence the salaries of many. Significance (Oxford, England), 14(6):38–41, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2017:GIDc

[1079] R. Allan Reese. Graphical interpretations of data: Key decisions. Significance (Oxford, England), 14(6):42–43, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:DFCf

[1080] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 14 (6):44, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Rj

[1081] Anonymous. In review. Significance (Oxford, England), 14(6):45, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Rk

[1082] Anonymous. In response. Significance (Oxford, England), 14(6):46–47, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2017:Wf

[1083] Anonymous. On the web. Significance (Oxford, England), 14(6):48, December 2017. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:I

[1084] Brian Tarran. The intro. Significance (Oxford, England), 15(1):1, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Ca

[1085] Anonymous. Contents. Significance (Oxford, England), 15(1):2–3, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:JDW

[1086] Brian Tarran. Justice Department wants citizenship question for 2020 Census. Significance (Oxford, England), 15(1):4–5, February 2018. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Chawla:2018:MYL

[1087] Dalmeet Singh Chawla. Mind your language. Significance (Oxford, England), 15(1):6–7, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rivera:2018:EDT

[1088] Roberto Rivera and Wolfgang Rolke. Estimating the death toll of Hurricane Maria. Significance (Oxford, England), 15(1):8–9, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Langkjaer-Bain:2018:DDD

[1089] Robert Langkjær-Bain. Data, debt and disaster in Puerto Rico. Significance (Oxford, England), 15(1):10–11, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fenton:2018:LVT

[1090] Norman Fenton, Martin Neil, and Nick Thieme. Lawnmowers versus terrorists. Significance (Oxford, England), 15(1):12–13, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bailey:2018:WOG

[1091] David Bailey. Why OUTLIERS are good for science. Significance (Oxford, England), 15(1):14–19, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Langkjaer-Bain:2018:FWD

[1092] Robert Langkjær-Bain. Five ways data is transforming music. Significance (Oxford, England), 15(1):20–23, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cahill:2018:NSS

[1093] Niamh Cahill. No “slowdown” in the story of global warming. Significance (Oxford, England), 15(1):24–27, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Charpentier:2018:WAL

[1094] Arthur Charpentier and Baptiste Coulmont. We are not alone! (At least, most of us aren’t). Significance (Oxford, England), 15(1):28–33, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bland:2018:DBS

[1095] Martin Bland. Diabetes, blood sugar and red wine: a personal study. Significance (Oxford, England), 15(1):34–37, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Canale:2018:CSB

[1096] Antonio Canale, Daniele Durante, Lucia Paci, and Bruno Scarpa. Connecting statistical brains. Significance (Oxford, England), 15(1):38–40, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Giorgi:2018:CHE

[1097] Emanuele Giorgi. Career highlights Emanuele Giorgi on teaching in Africa. Significance (Oxford, England), 15(1):41, February 2018. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2018:GIDa

[1098] R. Allan Reese. Graphical interpretations of data Truth and persuasion. Significance (Oxford, England), 15(1):42–43, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:DFCa

[1099] Anonymous. Dr Fisher’s casebook Lost in translation. Significance (Oxford, England), 15(1):44, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Ra

[1100] Anonymous. In Review. Significance (Oxford, England), 15(1):45, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Rb

[1101] Anonymous. In response. Significance (Oxford, England), 15(1):46–47, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Wa

[1102] Anonymous. On the web. Significance (Oxford, England), 15(1):48, February 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:DD

[1103] Brian Tarran. Decisions, decisions. Significance (Oxford, England), 15 (2):1, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Cb

[1104] Anonymous. Contents. Significance (Oxford, England), 15(2):2–3, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:OSF

[1105] Brian Tarran. One step forward, two steps back for stats independence in Puerto Rico. Significance (Oxford, England), 15(2):4–5, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:TVU

[1106] Brian Tarran. Trust and value: The UK Statistics Authority at 10. Significance (Oxford, England), 15(2):6–7, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:ITG

[1107] Brian Tarran. Is it time for GDP to be dethroned? Significance (Oxford, England), 15(2):8–9, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Pierson:2018:FTK

[1108] Steve Pierson. Five things to know about Trump’s FY19 budget request. Significance (Oxford, England), 15(2):9, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kizilersu:2018:WD

[1109] Ayşe Kızılersü, Markus Kreer, and Anthony W. Thomas. The Weibull distribution. Significance (Oxford, England), 15(2):10–11, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Krall:2018:ASW

[1110] Jenna R. Krall. Ask a statistician: What is the best way to report emissions? Significance (Oxford, England), 15(2):12–13, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Guttorp:2018:HSB

[1111] Peter Guttorp and Thordis L. Thorarinsdottir. How to save Bergen from the sea? Decisions under uncertainty. Significance (Oxford, England), 15 (2):14–18, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Constantinou:2018:TKA

[1112] Anthony C. Constantinou and Norman Fenton. Things to know about Bayesian networks: Decisions under uncertainty, part 2. Significance (Oxford, England), 15(2):19–23, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wolski:2018:HSC

[1113] Piotr Wolski. How severe is Cape Town’s “Day Zero” drought? Significance (Oxford, England), 15(2):24–27, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hayes:2018:RTF

[1114] Richard Hayes. Randomising towns to fight HIV. Significance (Oxford, England), 15(2):28–31, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

ONeill:2018:RRI

[1115] Rob O’Neill, Jeff Ralph, and Paul Smith. The rise and rise of inflation. Significance (Oxford, England), 15(2):32–37, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Zaharatos:2018:TWD

[1116] Brian Zaharatos and Michael Glawson. Trigger warning Disturbing data sets in the classroom. Significance (Oxford, England), 15(2):38–40, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:CSL

[1117] Brian Tarran. A cure for “statophobia” at the Library of Alexandria. Significance (Oxford, England), 15(2):41, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2018:GIDb

[1118] R. Allan Reese. Graphical interpretations of data: Know what I mean? Significance (Oxford, England), 15(2):42–43, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:DFCb

[1119] Anonymous. Dr Fisher’s casebook: Simply transparent. Significance (Oxford, England), 15(2):44, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Rc

[1120] Anonymous. In Review. Significance (Oxford, England), 15(2):45, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Rd

[1121] Anonymous. In response. Significance (Oxford, England), 15(2):46–47, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Wb

[1122] Anonymous. On the web. Significance (Oxford, England), 15(2):48, April 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:FS

[1123] Brian Tarran. Fields of statistics. Significance (Oxford, England), 15 (3):1, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Cc

[1124] Anonymous. Contents. Significance (Oxford, England), 15(3):2–3, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:WCW

[1125] Brian Tarran. What can we learn from the Facebook–Cambridge Analytica scandal? Significance (Oxford, England), 15(3):4–5, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Olhede:2018:AS

[1126] Sofia Olhede and Patrick Wolfe. The AI spring of 2018. Significance (Oxford, England), 15(3):6–7, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kobak:2018:PPR

[1127] Dmitry Kobak, Sergey Shpilkin, and Maxim S. Pshenichnikov. Putin’s peaks: Russian election data revisited. Significance (Oxford, England), 15(3):8–9, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kaur:2018:ZTC

[1128] Dipleen Kaur and Jenna R. Krall. “Zombie trucks”: Cleaning up the working dead. Significance (Oxford, England), 15(3):10–11, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Grajales:2018:ASD

[1129] Carlos Grajales. Ask a statistician: Does voter turnout affect election outcomes? Significance (Oxford, England), 15(3):12–13, June 2018. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Langkjaer-Bain:2018:WSM

[1130] Robert Langkjær-Bain. Where the seeds of modern statistics were sown. Significance (Oxford, England), 15(3):14–19, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mardia:2018:DSY

[1131] Kanti V. Mardia, Fred L. Bookstein, Balvinder S. Khambay, and John T. Kent. Deformations and smile: 100 years of D’Arcy Thompson’s On Growth and Form. Significance (Oxford, England), 15(3):20–25, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bornn:2018:SAU

[1132] Luke Bornn, Dan Cervone, and Javier Fernandez. Soccer analytics: Unravelling the complexity of “the beautiful game”. Significance (Oxford, England), 15(3):26–29, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Worcester:2018:BBT

[1133] Sir Robert Worcester, Roger Mortimore, Paul Baines, and Mark Gill. Brexit Britain, two years on. Significance (Oxford, England), 15(3): 30–33, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Salahub:2018:AHE

[1134] Christopher Salahub and Wayne Oldford. About “her emails”. Significance (Oxford, England), 15(3):34–37, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Schramm:2018:SPS

[1135] Harrison Schramm, Tracey Pérez Koehlmoos, and Scott Nestler. Statistics in pursuit of social justice. Significance (Oxford, England), 15(3): 38–40, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Thomas:2018:FTI

[1136] Hannah Thomas. Five things I’ve learned… How to communicate statistics. Significance (Oxford, England), 15(3):41, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2018:GIDc

[1137] R. Allan Reese. Graphical interpretations of data: Upright types. Significance (Oxford, England), 15(3):42–43, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:DFCc

[1138] Anonymous. Dr Fisher’s casebook: What do you expect? Significance (Oxford, England), 15(3):44, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Re

[1139] Anonymous. In review. Significance (Oxford, England), 15(3):45, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Rf

[1140] Anonymous. In response. Significance (Oxford, England), 15(3):46–47, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Wc

[1141] Anonymous. On the web. Significance (Oxford, England), 15(3):48, June 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:BNG

[1142] Brian Tarran. Breaking news, and ground-breaking events. Significance (Oxford, England), 15(4):1, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Cd

[1143] Anonymous. Contents. Significance (Oxford, England), 15(4):2–3, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:AGL

[1144] Brian Tarran. Andreas Georgiou’s legal troubles show no sign of ending. Significance (Oxford, England), 15(4):4–5, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Pearl:2018:MD

[1145] Judea Pearl and Dana Mackenzie. Mind over data. Significance (Oxford, England), 15(4):6–7, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hand:2018:WTY

[1146] David J. Hand. Who told you that? Significance (Oxford, England), 15 (4):8–9, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fanshawe:2018:MD

[1147] Tom Fanshawe. The matching distribution. Significance (Oxford, England), 15(4):10–11, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mastrodomenico:2018:TMT

[1148] Rob Mastrodomenico. Thanos, the Malthusian Titan. Significance (Oxford, England), 15(4):12–13, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Thieme:2018:RG

[1149] Nick Thieme. R generation. Significance (Oxford, England), 15(4):14–19, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (elec-

tronic). URL https://rss.onlinelibrary.wiley.com/doi/full/10.
1111/j.1740-9713.2018.01169.x.

Charlesworth:2018:CV

[1150] Brian Charlesworth and Anthony W. F. Edwards. A century of variance. Significance (Oxford, England), 15(4):20–25, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Raper:2018:TPB

[1151] Simon Raper. Turning points: Bernoulli’s golden theorem. Significance (Oxford, England), 15(4):26–29, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wininger:2018:PHS

[1152] Mike Wininger and Reva Johnson. Prosthetic hand signals. Significance (Oxford, England), 15(4):30–35, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kovalchik:2018:WTS

[1153] Stephanie Kovalchik. Why the tennis “serve clock” may be a waste of time. Significance (Oxford, England), 15(4):36–39, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Stark:2018:CCS

[1154] Philip B. Stark and Andrea Saltelli. Cargo-cult statistics and scientific crisis. Significance (Oxford, England), 15(4):40–43, August 2018. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:DFCd

[1155] Anonymous. Dr Fisher’s casebook: Life on the farm. Significance (Oxford, England), 15(4):44, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Rg

[1156] Anonymous. In review. Significance (Oxford, England), 15(4):45, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Rh

[1157] Anonymous. In response. Significance (Oxford, England), 15(4):46–47, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:WPCa

[1158] Anonymous. Wiley Prize Crossword: Double Scale by Goujeers. Significance (Oxford, England), 15(4):47, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:WCC

[1159] Anonymous. On the web: Chasing criminals through the family tree. Significance (Oxford, England), 15(4):48, August 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:JJS

[1160] Brian Tarran. Judge, jury and statistician. Significance (Oxford, England), 15(5):1, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Ce

[1161] Anonymous. Contents. Significance (Oxford, England), 15(5):2–3, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:ADS

[1162] Brian Tarran. Algorithms “don’t see the world the same way we do”. Significance (Oxford, England), 15(5):4–5, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Duffy:2018:FFF

[1163] Bobby Duffy. False “facts” and faulty thinking. Significance (Oxford, England), 15(5):6–7, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Scarrott:2018:ZMS

[1164] Diana Scarrott. Zachary Macaulay: the statistician who fought to end slavery in British colonies. Significance (Oxford, England), 15(5):8–9, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Geraci:2018:NLD

[1165] Marco Geraci and Mario Cortina Borja. Notebook: The Laplace distribution. Significance (Oxford, England), 15(5):10–11, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:CAM

[1166] Brian Tarran. The case of Anar Meshimbayeva. Significance (Oxford, England), 15(5):12–13, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mwareya:2018:TPM

[1167] Ray Mwareya. Team Pachedu makes its mark in Zimbabwe. Significance (Oxford, England), 15(5):13, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Thieme:2018:SC

[1168] Nick Thieme. Statistics in court. Significance (Oxford, England), 15(5): 14–17, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jacobsen:2018:DD

[1169] Jonny Jacobsen. Defending the data. Significance (Oxford, England), 15 (5):18–23, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mannshardt:2018:AQU

[1170] Elizabeth Mannshardt and Liz Naess. Air quality in the USA. Significance (Oxford, England), 15(5):24–27, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Smith:2018:CSS

[1171] Letisha Smith. Cooking up statistics: the science and the art. Significance (Oxford, England), 15(5):28–33, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:GRS

[1172] Brian Tarran. Gerrymandering, redistricting and statistics: a Q&A with Vince Barabba. Significance (Oxford, England), 15(5):34–39, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gelman:2018:ESP

[1173] Andrew Gelman. Ethics in statistical practice and communication: Five recommendations. Significance (Oxford, England), 15(5):40–43, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:DFCe

[1174] Anonymous. Dr Fisher’s casebook: For the sake of argument. Significance (Oxford, England), 15(5):44, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Ri

[1175] Anonymous. In review. Significance (Oxford, England), 15(5):45, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Rj

[1176] Anonymous. In response. Significance (Oxford, England), 15(5):46–47, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:WPCb

[1177] Anonymous. Wiley Prize Crossword: Squeezed in by Sam Buttrey. Significance (Oxford, England), 15(5):47, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:WHG

[1178] Anonymous. On the web: Hollywood’s gender inequality in numbers. Significance (Oxford, England), 15(5):48, October 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:SPA

[1179] Brian Tarran. A starting point for action. Significance (Oxford, England), 15(6):1, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Cf

[1180] Anonymous. Contents. Significance (Oxford, England), 15(6):2–3, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:TSL

[1181] Brian Tarran. Tanzania’s stats law amendments ‘threaten fact-checking’. Significance (Oxford, England), 15(6):4–5, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lewis:2018:CA

[1182] Michael Anthony Lewis. Civic arithmetic. Significance (Oxford, England), 15(6):6–7, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2018:PR

[1183] Brian Tarran. Playing with risk. Significance (Oxford, England), 15(6): 7, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Purdam:2018:NPM

[1184] Kingsley Purdam, Elisabeth Garratt, and Graham Whitham. A new poverty measure aims for consensus. Significance (Oxford, England), 15 (6):8, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rodamar:2018:TOL

[1185] Jeffery Rodamar. There ought to be a law! Campbell versus Goodhart. Significance (Oxford, England), 15(6):9, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spagat:2018:DCD

[1186] Michael Spagat, Britta Jewell, and Nicholas Jewell. Dealing with confusion over death “counts”. Significance (Oxford, England), 15(6):10, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bland:2018:DRM

[1187] Martin Bland. Death rates and modern life expectancy. Significance (Oxford, England), 15(6):11, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Langkjaer-Bain:2018:DRW

[1188] Robert Langkjær-Bain. Data rights and wrongs. Significance (Oxford, England), 15(6):12–15, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Towe:2018:WEF

[1189] Ross Towe, Jonathan Tawn, and Rob Lamb. Why extreme floods are more common than you might think. Significance (Oxford, England), 15(6):16–21, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ralph:2018:HTC

[1190] Jeff Ralph. How times have changed for teenagers. Significance (Oxford, England), 15(6):22–25, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kendall:2018:DRS

[1191] Graham Kendall. Did a roulette system “break the bank”? Significance (Oxford, England), 15(6):26–29, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Schoenberg:2018:PGL

[1192] Frederic Paik Schoenberg. Is poker a game of luck or skill? Significance (Oxford, England), 15(6):30–33, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Sawyer:2018:SSS

[1193] Kim Sawyer. Synchronicity and the search for significance. Significance (Oxford, England), 15(6):34–37, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McShane:2018:WMB

[1194] Blakeley B. McShane and Ulf Böckenholt. Want to make behavioural research more replicable? Promote single paper meta-analysis. Significance (Oxford, England), 15(6):38–40, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:DFCf

[1195] Anonymous. Dr Fisher’s casebook: the cycle of confusion. Significance (Oxford, England), 15(6):41, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Grant:2018:CC

[1196] Robert Grant. Calculate and communicate. Significance (Oxford, England), 15(6):42–44, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Rk

[1197] Anonymous. In review. Significance (Oxford, England), 15(6):45, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:Rl

[1198] Anonymous. In response. Significance (Oxford, England), 15(6):46–47, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:WPCc

[1199] Anonymous. Wiley Prize Crossword: Reflexive by Goujeers. Significance (Oxford, England), 15(6):47, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2018:SRA

[1200] Anonymous. Still running against the odds. Significance (Oxford, England), 15(6):48, December 2018. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:RT

[1201] Brian Tarran. The return of the Titanic. Significance (Oxford, England), 16(1):1, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Ca

[1202] Anonymous. Contents. Significance (Oxford, England), 16(1):2–3, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:DJF

[1203] Brian Tarran. Democrats join fight to keep ERS in Washington, DC. Significance (Oxford, England), 16(1):4–5, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Olhede:2019:AIF

[1204] Sofia Olhede and Patrick J. Wolfe. Artificial intelligence and the future of work: Will our jobs be taken by machines? Significance (Oxford, England), 16(1):6–7, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cochran:2019:WB

[1205] James J. Cochran. What is the bootstrap? Significance (Oxford, England), 16(1):8–9, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dagpunar:2019:GD

[1206] John Dagpunar. The gamma distribution. Significance (Oxford, England), 16(1):10–11, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:WCW

[1207] Brian Tarran. What can the web tell us about jobs and gender? Significance (Oxford, England), 16(1):12, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Winter:2019:HDW

[1208] Hugo Winter and Kate Brown. How do we know how fast the wind may blow? Significance (Oxford, England), 16(1):13, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Friendly:2019:VTD

[1209] Michael Friendly, Jürgen Symanzik, and Ortac Onder. Visualising the Titanic disaster. Significance (Oxford, England), 16(1):14–19, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Raper:2019:TPF

[1210] Simon Raper. Turning points: Fisher’s random idea. Significance (Oxford, England), 16(1):20–23, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kashlak:2019:WNO

[1211] Adam B. Kashlak. A wonderful night for Oscar speeches. Significance (Oxford, England), 16(1):24–27, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:CYS

[1212] Brian Tarran. The crisis in Yemen: Short of food, short of data? Significance (Oxford, England), 16(1):28, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Oakford:2019:DBD

[1213] Samuel Oakford. Deaths before data. Significance (Oxford, England), 16(1):29–31, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Johnston:2019:WGM

[1214] Ron Johnston, Kelvyn Jones, and David Manley. Why geography matters. Significance (Oxford, England), 16(1):32–37, February 2019. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bird:2019:GDU

[1215] Sheila M. Bird. Gender diversity in UK public appointments: a personal investigation. Significance (Oxford, England), 16(1):38–40, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:DFCa

[1216] Anonymous. Dr Fisher’s casebook: a walk on the Wilde side. Significance (Oxford, England), 16(1):41, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2019:GIDa

[1217] R. Allan Reese. Graphical interpretations of data: East is east. … Significance (Oxford, England), 16(1):42–44, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Ra

[1218] Anonymous. In review. Significance (Oxford, England), 16(1):45, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Rb

[1219] Anonymous. In response. Significance (Oxford, England), 16(1):46–47, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:WPCa

[1220] Anonymous. Wiley Prize Crossword: Down to Earth by Sam Buttrey. Significance (Oxford, England), 16(1):47, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:MML

[1221] Anonymous. March Madness, and the limits of basketball knowledge. Significance (Oxford, England), 16(1):48, February 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:EU

[1222] Brian Tarran. Evidence of uncertainty. Significance (Oxford, England), 16(2):1, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Cb

[1223] Anonymous. Contents. Significance (Oxford, England), 16(2):2–3, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:ESS

[1224] Brian Tarran. Is this the end of “statistical significance”? Significance (Oxford, England), 16(2):4–5, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jones:2019:SAN

[1225] Chris Jones and Arthur Pewsey. The sinh-arcsinh normal distribution. Significance (Oxford, England), 16(2):6–7, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Langkjaer-Bain:2019:WWT

[1226] Robert Langkjær-Bain. The world we think we know. Significance (Oxford, England), 16(2):8–9, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lohr:2019:EAO

[1227] Sharon Lohr. Edith Abbott and the origins of Chicago’s crime statistics. Significance (Oxford, England), 16(2):10–11, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:LFE

[1228] Brian Tarran. The limits of forensic evidence. Significance (Oxford, England), 16(2):12–15, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kafadar:2019:NOM

[1229] Karen Kafadar. The need for objective measures in forensic evidence. Significance (Oxford, England), 16(2):16–20, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Stern:2019:RVF

[1230] Hal S. Stern, Maria Cuellar, and David Kaye. Reliability and validity of forensic science evidence. Significance (Oxford, England), 16(2): 21–24, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mejia:2019:WDM

[1231] Robin Mejia, Maria Cuellar, Dana Delger, and Bill Eddy. What does a match mean? A framework for understanding forensic comparisons. Significance (Oxford, England), 16(2):25–28, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Carriquiry:2019:MLF

[1232] Alicia Carriquiry, Heike Hofmann, Xiao Hui Tai, and Susan VanderPlas. Machine learning in forensic applications. Significance (Oxford, England), 16(2):29–35, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jackson:2019:SAW

[1233] Graham Jackson and Alex Biedermann. “Source” or “activity”: What is the level of issue in a criminal trial? Significance (Oxford, England), 16 (2):36–39, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:QFQ

[1234] Brian Tarran. Quality in focus: A Q&A with the Forensic Science Regulator for England and Wales. Significance (Oxford, England), 16(2): 40–42, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:DFCb

[1235] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 16 (2):43–44, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2019:RGS

[1236] R. Allan Reese. R guides to statistics and graphics: [Getting Started with R: An Introduction for Biologists, Beckerman, Andrew, Childs, Dylan and Petchey, Owen; Oxford University Press; 256pp; February 2017; £22.99, $34.95; ISBN 978-0-19-878784-6. The New Statistics with R, Hector, Andy; Oxford University Press; 224pp; January 2015; £31.49, $49.95; ISBN 978-0-19-872906-8. R Graphics, third edition, Murrell, Paul; Chapman and Hall/CRC; 423pp; November 2018; £74.00, $99.95; ISBN 978-1-4987-8905-9]. Significance (Oxford, England), 16(2):45, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Rc

[1237] Anonymous. In response. Significance (Oxford, England), 16(2):46–47, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:WPCb

[1238] Anonymous. Wiley Prize Crossword: Cyclic Group by Goujeers. Significance (Oxford, England), 16(2):47, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:PCM

[1239] Anonymous. Preventing cancer: mere rhetoric or a promising plan? Significance (Oxford, England), 16(2):48, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:CC

[1240] Brian Tarran. Cause for concern. Significance (Oxford, England), 16 (3):1, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Cc

[1241] Anonymous. Contents. Significance (Oxford, England), 16(3):2–3, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:FSC

[1242] Brian Tarran. Forensic science in “crisis”, needs funding for research. Significance (Oxford, England), 16(3):4–5, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Andrews:2019:DS

[1243] Rj Andrews. Data shadows. Significance (Oxford, England), 16(3):6–7, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:PUA

[1244] Brian Tarran. Police use of algorithms “needs new regs”. Significance (Oxford, England), 16(3):7, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Roy:2019:RWV

[1245] Prannoy Roy and Dorab R. Sopariwala. On the rise: women voters in India. Significance (Oxford, England), 16(3):8–9, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McVicar:2019:DDU

[1246] Duncan McVicar, Julie Moschion, and Jan van Ours. Does drug use lead to homelessness for young, disadvantaged people? Significance (Oxford, England), 16(3):10–11, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Yan:2019:TGH

[1247] Yuan Yan and Marc G. Genton. The Tukey g-and-h distribution. Significance (Oxford, England), 16(3):12–13, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mastrodomenico:2019:ASD

[1248] Rob Mastrodomenico. Ask a statistician: Does Manchester United really have 100 million followers in China? Significance (Oxford, England), 16 (3):14–15, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Langkjaer-Bain:2019:TLF

[1249] Robert Langkjær-Bain. The troubling legacy of Francis Galton. Significance (Oxford, England), 16(3):16–21, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Raper:2019:TPP

[1250] Simon Raper. Turning points Pearson and the parameter. Significance (Oxford, England), 16(3):22–25, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Di-Luvi:2019:HDI

[1251] Gian Carlo Di-Luvi. How does it make you feel? Significance (Oxford, England), 16(3):26–29, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Duckworth:2019:CRC

[1252] Frank Duckworth, Tony Lewis, and Steven Stern. Cricket’s raining champion: Two decades of Duckworth–Lewis (and Stern). Significance (Oxford, England), 16(3):30–35, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Golbeck:2019:MSW

[1253] Amanda L. Golbeck. A “mover and shaker” for women in statistics. Significance (Oxford, England), 16(3):36–38, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Thornton:2019:FAL

[1254] Suzanne Thornton, Brittany Green, and Emma Benn. Friends and allies: LGBT+ inclusion in statistics and data science. Significance (Oxford, England), 16(3):39–41, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2019:GIDb

[1255] R. Allan Reese. Graphical interpretations of data: a hole lot of questions. Significance (Oxford, England), 16(3):42–43, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:DFCc

[1256] Anonymous. Dr Fisher’s casebook: Keeping it under wraps. Significance (Oxford, England), 16(3):44, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Rd

[1257] Anonymous. In Review. Significance (Oxford, England), 16(3):45, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Re

[1258] Anonymous. In response. Significance (Oxford, England), 16(3):46–47, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:WPCc

[1259] Anonymous. Wiley Prize Crossword: Nylon by Sam Buttrey. Significance (Oxford, England), 16(3):47, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:WRR

[1260] Anonymous. Water result for RSS Stat of the Year. Significance (Oxford, England), 16(3):48, June 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:TT

[1261] Brian Tarran. Tick tock. … Significance (Oxford, England), 16(4):1, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Cd

[1262] Anonymous. Contents. Significance (Oxford, England), 16(4):2–3, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:Na

[1263] Brian Tarran. News. Significance (Oxford, England), 16(4):4–5, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:TEM

[1264] Brian Tarran. Through the “eyes” of the machine. Significance (Oxford, England), 16(4):6–7, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lowe:2019:AIS

[1265] David Lowe. “Artificial intelligence”, or statistics? Significance (Oxford, England), 16(4):7, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Leemis:2019:TB

[1266] Lawrence M. Leemis and Raghu Pasupathy. The ties that bind. Significance (Oxford, England), 16(4):8–9, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cernat:2019:MES

[1267] Alexandru Cernat, Joseph W. Sakshaug, and Javier Castillo. Measurement error: Skin colour and the role of the interviewer. Significance (Oxford, England), 16(4):10–11, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harkness:2019:TR

[1268] Timandra Harkness. Take a risk. Significance (Oxford, England), 16(4): 12–13, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:WLW

[1269] Brian Tarran. THE S WORD … and what to do about it. Significance (Oxford, England), 16(4):14, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Sheldon:2019:WDI

[1270] Neil Sheldon. What does it all mean? Significance (Oxford, England), 16(4):15–17, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cready:2019:CG

[1271] William M. Cready. Complacency at the gates. Significance (Oxford, England), 16(4):18–19, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hahn:2019:SIS

[1272] Gerald J. Hahn, Necip Doganaksoy, and William Q. Meeker. Statistical intervals, not statistical significance. Significance (Oxford, England), 16 (4):20–22, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Chambers:2019:RRR

[1273] Chris Chambers. The registered reports revolution: Lessons in cultural reform. Significance (Oxford, England), 16(4):23–27, August 2019. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Python:2019:DFT

[1274] Andre Python, Janine B. Illian, Charlotte M. Jones-Todd, and Marta Blangiardo. The deadly facets of terrorism. Significance (Oxford, England), 16(4):28–31, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Molgaard:2019:MVS

[1275] Craig A. Molgaard. Military vital statistics: The Spanish flu and the First World War. Significance (Oxford, England), 16(4):32–37, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Marriott:2019:GPG

[1276] Nigel Marriott. Gender pay gaps: a median-sized problem. Significance (Oxford, England), 16(4):38–41, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2019:GIDc

[1277] R. Allan Reese. Graphical interpretations of data: Time and tide. … Significance (Oxford, England), 16(4):42–43, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:DFCd

[1278] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 16 (4):44, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Rf

[1279] Anonymous. In Review. Significance (Oxford, England), 16(4):45, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Rg

[1280] Anonymous. In response. Significance (Oxford, England), 16(4):46–47, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:P

[1281] Anonymous. Puzzle. Significance (Oxford, England), 16(4):47, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Ce

[1282] Anonymous. In Closing. Significance (Oxford, England), 16(4):48, August 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:HD

[1283] Brian Tarran. Heroes of democracy. Significance (Oxford, England), 16 (5):1, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Cf

[1284] Anonymous. Contents. Significance (Oxford, England), 16(5):2–3, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:Nb

[1285] Brian Tarran. News. Significance (Oxford, England), 16(5):4–5, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kazemi:2019:OUD

[1286] Arianna Kazemi, Connor Kennedy, Gabri Silverman, and Nicholas G. Reich. Opioids in the USA: Disparities in addiction and incarceration. Significance (Oxford, England), 16(5):6–7, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Marchant:2019:DBW

[1287] Paul Marchant. Do brighter, whiter street lights improve road safety? Significance (Oxford, England), 16(5):8–9, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nestler:2019:VGD

[1288] Scott Nestler and Andrew Hall. The variance gamma distribution. Significance (Oxford, England), 16(5):10–11, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Shaw:2019:FBA

[1289] Liam P. Shaw and Luke F. Shaw. The flying bomb and the actuary. Significance (Oxford, England), 16(5):12–17, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dalal:2019:HMR

[1290] Siddhartha Dalal, Berkay Adlim, and Michael Lesk. How to measure relative bias in media coverage? Significance (Oxford, England), 16(5): 18–23, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Woodward:2019:MB

[1291] Phil Woodward and Sam Woodward. Mining the BoardGameGeek. Significance (Oxford, England), 16(5):24–29, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Billard:2019:ASA

[1292] Lynne Billard. The American Statistical Association and the US Census: a shared history. Significance (Oxford, England), 16(5):30–34, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Auerbach:2019:WAD

[1293] Jonathan Auerbach, Quentin Brummet, John Czajka, George C. Hough, Eddie Hunsinger, and Joseph Salvo. Will administrative data save government surveys? Significance (Oxford, England), 16(5):35–39, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Carrigan:2019:SLC

[1294] Chris Carrigan. Saving lives, or costing them? The unintended consequences of data protection. Significance (Oxford, England), 16(5):40–41, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Raper:2019:MOB

[1295] Simon Raper. In my opinion… Bullshit jobs in statistics. Significance (Oxford, England), 16(5):42–43, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:DFCe

[1296] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 16 (5):44, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Rh

[1297] Anonymous. In Review. Significance (Oxford, England), 16(5):45, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Ri

[1298] Anonymous. In response. Significance (Oxford, England), 16(5):46–47, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:WPCd

[1299] Anonymous. Wiley Prize Crossword: Break Time by Sam Buttrey. Significance (Oxford, England), 16(5):47, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Cg

[1300] Anonymous. In Closing. Significance (Oxford, England), 16(5):48, October 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:TY

[1301] Brian Tarran. It’s that time of year. Significance (Oxford, England), 16 (6):1, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Ch

[1302] Anonymous. Contents. Significance (Oxford, England), 16(6):2–3, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:GCC

[1303] Brian Tarran. German commission calls for risk-based regulation of algorithmic systems. Significance (Oxford, England), 16(6):4–5, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nagaraja:2019:MSW

[1304] Chaitra H. Nagaraja. Measuring society: Who, what, when, where, why and how. Significance (Oxford, England), 16(6):6–7, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:ETT

[1305] Brian Tarran. Education trials and tribulations. Significance (Oxford, England), 16(6):8–9, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Stewart:2019:SAU

[1306] Ian Stewart. Six ages of uncertainty. Significance (Oxford, England), 16(6):10–11, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2019:US

[1307] Brian Tarran. An uncertain state. Significance (Oxford, England), 16(6): 11, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hodge:2019:PIC

[1308] David Hodge. Pick of the ice cream sticks. Significance (Oxford, England), 16(6):12–13, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fawcett:2019:SF

[1309] Lee Fawcett. A severe forecast. Significance (Oxford, England), 16(6): 14–19, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mateen:2019:AWC

[1310] Bilal Mateen and Raphael Sonabend. All I want for Christmas is… Rigorous validation of predictive models to prevent hasty generalisations. Significance (Oxford, England), 16(6):20–24, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Langkjaer-Bain:2019:CM

[1311] Robert Langkjær-Bain. Charts and minds. Significance (Oxford, England), 16(6):25–27, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mendoza:2019:SPQ

[1312] Manuel Mendoza. Statistics and politics: Quick counts in Mexico. Significance (Oxford, England), 16(6):28–31, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

deWet:2019:CSB

[1313] Tertius de Wet, Jef Teugels, and Pieta van Deventer. Campanometry: The statistics of bells. Significance (Oxford, England), 16(6):32–37, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Alba:2019:GMH

[1314] Sandra Alba, Franz Wong, and Yngve Bråten. Gender matters in household surveys. Significance (Oxford, England), 16(6):38–41, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bland:2019:TT

[1315] Martin Bland. Targeted teaching. Significance (Oxford, England), 16(6): 42–43, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:DFCf

[1316] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 16 (6):44, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Rj

[1317] Anonymous. In Review. Significance (Oxford, England), 16(6):45, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2019:Rk

[1318] Anonymous. In response. Significance (Oxford, England), 16(6):46–47, December 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Brown:2019:TSA

[1319] Celia Brown. “Thus spake an angel who forthwith explained the RCT… ”. Significance (Oxford, England), 16(6):48, December 2019. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:NYF

[1320] Brian Tarran. New year, familiar problems. Significance (Oxford, England), 17(1):1, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:Ca

[1321] Anonymous. Contents. Significance (Oxford, England), 17(1):2–3, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:Na

[1322] Brian Tarran. News. Significance (Oxford, England), 17(1):4–5, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Pilling:2020:NS

[1323] Mark Pilling. “Nudge” and statistics. Significance (Oxford, England), 17 (1):6–7, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dawkins:2020:WDC

[1324] Laura Dawkins. What drives commuter behaviour? Significance (Oxford, England), 17(1):8–9, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Held:2020:IRS

[1325] Leonhard Held and Simon Schwab. Improving the reproducibility of science. Significance (Oxford, England), 17(1):10–11, February 2020. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rubio:2020:FTP

[1326] Francisco J. Rubio and Mark F. J. Steel. The family of two-piece distributions. Significance (Oxford, England), 17(1):12–13, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wallace:2020:AIW

[1327] Michael Wallace. Analysis in an imperfect world. Significance (Oxford, England), 17(1):14–19, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Finkelstein:2020:WPB

[1328] Michael O. Finkelstein and Bruce Levin. Why plea bargains are a bad deal for some. Significance (Oxford, England), 17(1):20–25, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Oke:2020:CSS

[1329] Jason Oke and Tom Fanshawe. Cancer statistics: a survival guide. Significance (Oxford, England), 17(1):26–29, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Salvo:2020:CWI

[1330] Joseph J. Salvo, Annette Jacoby, and Arun Peter Lobo. Census 2020: Why increasing self-response is key to a good count. Significance (Oxford, England), 17(1):30–33, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Raper:2020:LBT

[1331] Simon Raper. Leo Breiman’s “Two Cultures”. Significance (Oxford, England), 17(1):34–37, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Goldstein:2020:LE

[1332] Harvey Goldstein. Living by the evidence. Significance (Oxford, England), 17(1):38–40, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harron:2020:PHG

[1333] Katie Harron, George Leckie, Bill Browne, James Carpenter, Fiona Steele, and Tim Cole. Professor Harvey Goldstein at 80. Significance (Oxford, England), 17(1):41, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:SLA

[1334] Brian Tarran. “Statistical literacy for all!”. Significance (Oxford, England), 17(1):42–43, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:DFCa

[1335] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 17 (1):44–45, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:LPa

[1336] Anonymous. Letters/ puzzle. Significance (Oxford, England), 17(1):46– 47, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:W

[1337] Anonymous. On The Web. Significance (Oxford, England), 17(1):48, February 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:FN

[1338] Brian Tarran. Finding Nightingale. Significance (Oxford, England), 17 (2):1, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:Cb

[1339] Anonymous. Contents. Significance (Oxford, England), 17(2):2–3, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:Nb

[1340] Brian Tarran. News. Significance (Oxford, England), 17(2):4–5, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Borja:2020:IBF

[1341] Mario Cortina Borja. The importance of being Florence. Significance (Oxford, England), 17(2):6–7, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Schwab:2020:DWC

[1342] Simon Schwab and Leonhard Held. Different worlds Confirmatory versus exploratory research. Significance (Oxford, England), 17(2):8–9, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McBride:2020:SBD

[1343] Rebekah McBride. The stories behind the data. Significance (Oxford, England), 17(2):10–11, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Alba:2020:TTA

[1344] Sandra Alba. Time to talk about trust. Significance (Oxford, England), 17(2):12–13, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fricker:2020:IDO

[1345] Ronald D. Fricker, Jr. and Steven E. Rigdon. Investigating a disease outbreak. Significance (Oxford, England), 17(2):14, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Andrews:2020:LTL

[1346] Rj Andrews. Let there be light: a tribute to Florence Nightingale. Significance (Oxford, England), 17(2):16–17, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McDonald:2020:RGO

[1347] Lynn McDonald. The real goods and the oversell. Significance (Oxford, England), 17(2):18–21, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lorenzo-Arribas:2020:FNN

[1348] Altea Lorenzo-Arribas and Pilar Cacheiro. Florence Nightingale’s network: Women, power, and scientific collaboration. Significance (Oxford, England), 17(2):22–25, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hedley:2020:FNV

[1349] Alison Hedley. Florence Nightingale and Victorian data visualisation. Significance (Oxford, England), 17(2):26–30, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:LO

[1350] Brian Tarran. A life in objects. Significance (Oxford, England), 17(2): 31–32, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Langkjaer-Bain:2020:MN

[1351] Robert Langkjær-Bain. Modern Nightingales. Significance (Oxford, England), 17(2):33–37, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rawlins:2020:NCA

[1352] Michael D. Rawlins and Andrew Dillon. NICE comes of age. Significance (Oxford, England), 17(2):38–41, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Golbeck:2020:SIC

[1353] Amanda Golbeck. Supporting an inclusive community: A Caucus for Women in Statistics. Significance (Oxford, England), 17(2):42–44, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:DFCb

[1354] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 17 (2):45, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:R

[1355] Anonymous. In response. Significance (Oxford, England), 17(2):46–47, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Singpurwalla:2020:SBS

[1356] Nozer D. Singpurwalla. “Sleeping beauties” in statistics. Significance (Oxford, England), 17(2):48, April 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:CPa

[1357] Brian Tarran. Change of plan. Significance (Oxford, England), 17(3):1, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:Cc

[1358] Anonymous. Contents. Significance (Oxford, England), 17(3):2–3, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:TCC

[1359] Brian Tarran. Tracking the Covid-19 conversation. Significance (Oxford, England), 17(3):4–5, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bo:2020:ICE

[1360] Erlend E. Bø, Elin Halvorsen, and Thor O. Thoresen. Investigating the “Carnegie effect”. Significance (Oxford, England), 17(3):6–7, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Baxter:2020:DLB

[1361] Michael Baxter. Have data leaks been plugged? Significance (Oxford, England), 17(3):8–9, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kabat:2020:TFM

[1362] Geoffrey C. Kabat. The two faces of meta-analysis. Significance (Oxford, England), 17(3):10–11, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:CPb

[1363] Brian Tarran. The Covid-19 pandemic. Significance (Oxford, England), 17(3):12, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ball:2020:HDE

[1364] Patrick Ball. How do epidemiologists know how many people will get Covid-19? Significance (Oxford, England), 17(3):12–14, June 2020. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cochran:2020:WWN

[1365] James J. Cochran. Why we need more coronavirus tests than we think we need. Significance (Oxford, England), 17(3):14–15, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bilder:2020:TSS

[1366] Christopher R. Bilder, Peter C. Iwen, Baha Abdalhamid, Joshua M. Tebbs, and Christopher S. McMahan. Tests in short supply? Try group testing. Significance (Oxford, England), 17(3):15–16, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Finazzi:2020:ICP

[1367] Francesco Finazzi and Alessandro Fassò. The impact of the Covid-19 pandemic on Italian mobility. Significance (Oxford, England), 17(3): 17–18, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cochran:2020:PEG

[1368] James J. Cochran. Pandemics and exponential growth. Significance (Oxford, England), 17(3):18–21, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fricker:2020:YFD

[1369] Ronald D. Fricker and Steven E. Rigdon. Yellow fever: Discovering the cause and designing effective experiments. Significance (Oxford, England), 17(3):22–27, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Friendly:2020:GGV

[1370] Michael Friendly and Howard Wainer. Galton’s gleam: Visual thinking and graphic discoveries. Significance (Oxford, England), 17(3):28–33, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Wheeler:2020:CSL

[1371] Dr Nicole Wheeler. Computer says …. Significance (Oxford, England), 17(3):34–37, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rubin:2020:TTS

[1372] Mark Rubin. That’s not a two-sided test! It’s two one-sided tests! Significance (Oxford, England), 17(3):38–41, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hand:2020:DD

[1373] David J. Hand. Dark data. Significance (Oxford, England), 17(3):42–44, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:DFCc

[1374] Anonymous. Dr Fisher’s casebook. Significance (Oxford, England), 17 (3):45, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:RP

[1375] Anonymous. In response/Puzzle. Significance (Oxford, England), 17(3): 46–47, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:AGU

[1376] Anonymous. An audio guide to uncertainty. Significance (Oxford, England), 17(3):48, June 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:Cd

[1377] Anonymous. Contents. Significance (Oxford, England), 17(4):1, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:ARR

[1378] Brian Tarran. Award “retired” over R. A. Fisher’s links to eugenics. Significance (Oxford, England), 17(4):2–3, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rivera:2020:RDP

[1379] Roberto Rivera and Janet Rosenbaum. Racial disparities in police stops in US cities. Significance (Oxford, England), 17(4):4–5, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Herbert:2020:SBP

[1380] Annie Herbert, Gareth Griffith, Gibran Hemani, and Luisa Zuccolo. The spectre of Berkson’s paradox: Collider bias in Covid-19 research. Significance (Oxford, England), 17(4):6–7, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:BDT

[1381] Brian Tarran. Battle deaths over time: What does the data say? Significance (Oxford, England), 17(4):7, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Schwab:2020:SAC

[1382] Simon Schwab and Leonhard Held. Science after Covid-19: Faster, better, stronger? Significance (Oxford, England), 17(4):8–9, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Neal:2020:CTL

[1383] Samuel R. Neal and Cindy Zheng. Covid-19 through the lens of the peerreviewed literature. Significance (Oxford, England), 17(4):10–11, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gunter:2020:GTF

[1384] Bert Gunter. Graphical tour de force, or mere artwork? Significance (Oxford, England), 17(4):12–13, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Python:2020:WWK

[1385] Andre Python. What we know — and don’t know — about terrorism. Significance (Oxford, England), 17(4):14, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:NLN

[1386] Brian Tarran. New look, new structure, new opportunities. Significance (Oxford, England), 17(4):15, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Langkjaer-Bain:2020:BI

[1387] Robert Langkjær-Bain. Not buying it! Significance (Oxford, England), 17(4):16–21, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harkness:2020:JGF

[1388] Timandra Harkness. John Graunt at 400: Fighting disease with numbers. Significance (Oxford, England), 17(4):22–25, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jay:2020:QDD

[1389] Matthew A. Jay and Mario Cortina Borja. Quomodo dicitur “data science” Latine? Significance (Oxford, England), 17(4):26–29, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

LanataBriones:2020:IAC

[1390] Cecilia T. Lanata Briones and Claudia Daniel. Inflation in Argentina: a controversial figure. Significance (Oxford, England), 17(4):30–35, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Efron:2020:CRR

[1391] Bradley Efron, Shun ichi Amari, Donald B. Rubin, Arni S. R. Srinivasa Rao, and David R. Cox. C. R. Rao’s century. Significance (Oxford, England), 17(4):36–38, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Plastino:2020:WBI

[1392] Angel Ricardo Plastino and Angelo Plastino. What’s the big idea? Cramér–Rao inequality and Rao distance. Significance (Oxford, England), 17(4):39, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Das:2020:EM

[1393] Shikta Das. Epidemiology and me. Significance (Oxford, England), 17 (4):40, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:HW

[1394] Anonymous. Herb Weisberg. Significance (Oxford, England), 17(4):41, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lane:2020:ACU

[1395] Julia Lane. After Covid-19, the US statistical system needs to change. Significance (Oxford, England), 17(4):42–43, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

King:2020:BED

[1396] Thomas King. Books: Effective Data Storytelling. Significance (Oxford, England), 17(4):43, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Steuer:2020:TDS

[1397] Detlef Steuer. Time for data science to professionalise. Significance (Oxford, England), 17(4):44–45, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:LPb

[1398] Anonymous. Letters/puzzle. Significance (Oxford, England), 17(4):46– 47, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:SSC

[1399] Anonymous. The secret statistician: Co-authors. Significance (Oxford, England), 17(4):48, August 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:Ta

[1400] Anonymous. TOC. Significance (Oxford, England), 17(5):1, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:Na

[1401] Anonymous. News. Significance (Oxford, England), 17(5):2–3, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Paulden:2020:CRM

[1402] Tim Paulden. A cutting re-mark. Significance (Oxford, England), 17(5): 4–5, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Forsberg:2020:PUP

[1403] Ole J. Forsberg. Polls and the US presidential election: real or fake? Significance (Oxford, England), 17(5):6–7, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kobak:2020:SPR

[1404] Dmitry Kobak, Sergey Shpilkin, and Maxim S. Pshenichnikov. Suspect peaks in Russia’s “referendum” results. Significance (Oxford, England), 17(5):8–9, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:DD

[1405] Anonymous. Digital detectives. Significance (Oxford, England), 17(5): 9, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jackson:2020:SLG

[1406] James Jackson. Statistics and the “little grey cells”. Significance (Oxford, England), 17(5):10–11, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Olhede:2020:BA

[1407] Sofia Olhede and Patrick J. Wolfe. Blame the algorithm? Significance (Oxford, England), 17(5):12, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:Ea

[1408] Brian Tarran. Editorial. Significance (Oxford, England), 17(5):13, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gustar:2020:LMF

[1409] Andrew Gustar. The laws of musical fame and obscurity. Significance (Oxford, England), 17(5):14–17, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Boulesteix:2020:RCM

[1410] Anne-Laure Boulesteix, Sabine Hoffmann, Alethea Charlton, and Heidi Seibold. A replication crisis in methodological research? Significance (Oxford, England), 17(5):18–21, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Iliev:2020:SSC

[1411] Iliyan R. Iliev, Xin Huang, and Yulia R. Gel. Speaking out or speaking in? Changes in political rhetoric over time. Significance (Oxford, England), 17(5):22–25, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Holmes:2020:SZ

[1412] Lisa Holmes, Katie Edwards, Andy Moss, Simon Tollington, Anna Fowler, David Hughes, and Maria Sudell. Statistics at the zoo. Significance (Oxford, England), 17(5):26–29, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Schuckers:2020:TSS

[1413] Michael Schuckers and Sarah Campbell. Telling statistical stories. Significance (Oxford, England), 17(5):30–33, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reese:2020:GID

[1414] R. Allan Reese. Graphical interpretations of data: Be alert! Significance (Oxford, England), 17(5):34–35, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harkness:2020:WDI

[1415] Timandra Harkness. What does it mean to mislead? Significance (Oxford, England), 17(5):36–38, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hoang:2020:WBI

[1416] Lê Nguyên Hoang. What’s the big idea? Bayes’ rule. Significance (Oxford, England), 17(5):39, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Leahy:2020:SYN

[1417] Joy Leahy. “As a statistician, you never know where you might end up!”. Significance (Oxford, England), 17(5):40–41, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:SHU

[1418] Brian Tarran. “Statistics help us see the world clearly if we’re willing to look”. Significance (Oxford, England), 17(5):42–43, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kolassa:2020:WAM

[1419] John Kolassa. When academia meets industry meets government. Significance (Oxford, England), 17(5):44–45, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:La

[1420] Anonymous. Letters. Significance (Oxford, England), 17(5):46–47, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:Pa

[1421] Anonymous. Puzzle. Significance (Oxford, England), 17(5):47, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:SSE

[1422] Anonymous. The secret statistician: Eugenics, then and now. Significance (Oxford, England), 17(5):48, October 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:Tb

[1423] Anonymous. TOC. Significance (Oxford, England), 17(6):1, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:Nb

[1424] Anonymous. News. Significance (Oxford, England), 17(6):2–3, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fricker:2020:CCA

[1425] Ronald D. Fricker, Jr. Covid-19 cases around the world. Significance (Oxford, England), 17(6):4–5, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Diniz:2020:BPW

[1426] Marcio Alves Diniz and David Richard Bellhouse. Bayes and Price: when did it start? Significance (Oxford, England), 17(6):6–7, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Russell:2020:DFC

[1427] Matthew Russell. Do forests have the capacity for 1 trillion extra trees? Significance (Oxford, England), 17(6):8–9, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Held:2020:RPR

[1428] Leonhard Held, Samuel Pawel, and Simon Schwab. Replication power and regression to the mean. Significance (Oxford, England), 17(6):10– 11, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Diggle:2020:DMU

[1429] Peter J. Diggle, Tim Gowers, Frank Kelly, and Neil Lawrence. Decisionmaking with uncertainty. Significance (Oxford, England), 17(6):12, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2020:Eb

[1430] Brian Tarran. Editorial. Significance (Oxford, England), 17(6):13, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Raper:2020:ED

[1431] Simon Raper. The end of determinism. Significance (Oxford, England), 17(6):14–17, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kaloskampis:2020:SDC

[1432] Ioannis Kaloskampis, Chaitanya Joshi, Catrin Cheung, David Pugh, and Louisa Nolan. Synthetic data in the civil service. Significance (Oxford, England), 17(6):18–23, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ibrahim:2020:PNN

[1433] Maria Ibrahim. Primum non nocere (first, do no harm). Significance (Oxford, England), 17(6):24–27, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Small:2020:NOS

[1434] Hugh Small. Nightingale’s overlooked Scutari statistics. Significance (Oxford, England), 17(6):28–33, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nolan:2020:RW

[1435] Deborah Nolan and Sara Stoudt. Reading to write. Significance (Oxford, England), 17(6):34–37, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Sawyer:2020:CQP

[1436] Kim Sawyer and Paul Halpern. Coincidence, quantum physics and statistics. Significance (Oxford, England), 17(6):38–40, December 2020. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lum:2020:CS

[1437] Kristian Lum. Career story. Significance (Oxford, England), 17(6):40, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Vinod:2020:WBI

[1438] H. D. Vinod. What’s the big idea? Ridge regression and regularisation. Significance (Oxford, England), 17(6):41, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Barnett:2020:HSY

[1439] Professor Adrian Barnett. How scrambling your data can help you think more clearly. Significance (Oxford, England), 17(6):42–44, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:RCB

[1440] Anonymous. Review: Calling Bullshit. Significance (Oxford, England), 17(6):45, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:Lb

[1441] Anonymous. Letters. Significance (Oxford, England), 17(6):46–47, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:Pb

[1442] Anonymous. Puzzle. Significance (Oxford, England), 17(6):47, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2020:SST

[1443] Anonymous. The secret statistician: a time of change. Significance (Oxford, England), 17(6):48, December 2020. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Ta

[1444] Anonymous. TOC. Significance (Oxford, England), 18(1):1, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Na

[1445] Anonymous. News. Significance (Oxford, England), 18(1):2–3, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Forsberg:2021:UEP

[1446] Ole J. Forsberg. US election polls: a quick postmortem. Significance (Oxford, England), 18(1):4–5, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Stubbs:2021:HLW

[1447] Joshua E. Stubbs and Toby C. T. Stubbs. How to lose weight well, according to How to Lose Weight Well. Significance (Oxford, England), 18(1):6–7, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Aldrich:2021:PS

[1448] John Aldrich. Passionate statisticians. Significance (Oxford, England), 18(1):8–9, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dunne-Willows:2021:ASS

[1449] Michael Dunne-Willows. Ask a statistician: a suspicious sequence? Significance (Oxford, England), 18(1):10, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2021:Ea

[1450] Brian Tarran. Editorial. Significance (Oxford, England), 18(1):11, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fricker:2021:COY

[1451] Ronald D. Fricker, Jr. Covid-19: One year on. … Significance (Oxford, England), 18(1):12–15, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kobak:2021:EMR

[1452] Dmitry Kobak. Excess mortality reveals Covid’s true toll in Russia. Significance (Oxford, England), 18(1):16–19, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Piegorsch:2021:TFH

[1453] Walter W. Piegorsch, Rachel R. McCaster, and Susan L. Cutter. From terrorism to flooding: How vulnerable is your city? Significance (Oxford, England), 18(1):20–25, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Zheng:2021:NMH

[1454] Yanyan Zheng, Marco Cavaglià, Ryan Quitzow-James, and Kentaro Mogushi. A needle in (many) haystacks: Using the false alarm rate to sift gravitational waves from noise. Significance (Oxford, England), 18(1): 26–31, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harkness:2021:LSM

[1455] Timandra Harkness. “‘The scientific method’ doesn’t leave you”. Significance (Oxford, England), 18(1):32–35, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Diebold:2021:WBI

[1456] Francis X. Diebold. What’s the big idea? “Big Data’’ and its origins. Significance (Oxford, England), 18(1):36–37, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Isaac:2021:CS

[1457] William Isaac. Career story. Significance (Oxford, England), 18(1):38– 39, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hoffman:2021:CVS

[1458] Katherine Hoffman. Covid-19: a view from the sidelines. Significance (Oxford, England), 18(1):40–43, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

vanDoorn:2021:SPC

[1459] Johnny van Doorn and Eric-Jan Wagenmakers. Strong public claims may not reflect researchers’ private convictions. Significance (Oxford, England), 18(1):44–45, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:La

[1460] Anonymous. Letters. Significance (Oxford, England), 18(1):46–47, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Pa

[1461] Anonymous. Puzzle. Significance (Oxford, England), 18(1):47, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:SSS

[1462] Anonymous. The secret statistician: Statistics for pleasure, if not profit. Significance (Oxford, England), 18(1):48, February 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Ca

[1463] Anonymous. Contents. Significance (Oxford, England), 18(2):1, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Nb

[1464] Anonymous. News. Significance (Oxford, England), 18(2):2–3, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kennedy-Shaffer:2021:LDV

[1465] Lee Kennedy-Shaffer. Land does not vote. Neither do percentages. Significance (Oxford, England), 18(2):4–5, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Schramm:2021:CHS

[1466] Harrison Schramm, Isaac Rubin, and Norah Schramm. Covid-19 and high school grades: an early case study. Significance (Oxford, England), 18(2):6–7, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Spearing:2021:SMS

[1467] Harry Spearing. Swimming meets statistics. Significance (Oxford, England), 18(2):8–9, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gregory:2021:TIS

[1468] Mary Gregory. Three ingredients for statistical leadership. Significance (Oxford, England), 18(2):10, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2021:Eb

[1469] Brian Tarran. Editorial. Significance (Oxford, England), 18(2):11, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harkness:2021:HDEa

[1470] Timandra Harkness. The history of the data economy. Significance (Oxford, England), 18(2):12–15, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Matthews:2021:PVS

[1471] Robert Matthews. The p-value statement, five years on. Significance (Oxford, England), 18(2):16–19, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Huebner:2021:SNW

[1472] Marianne Huebner, David E. Meltzer, and Aris Perperoglou. Strength in numbers: Women in Olympic-style weightlifting. Significance (Oxford, England), 18(2):20–25, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Zuccolotto:2021:AOB

[1473] Paola Zuccolotto, Marica Manisera, and Marco Sandri. Alley-oop! Basketball analytics in R. Significance (Oxford, England), 18(2):26–31, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

McConway:2021:SHS

[1474] Professor Kevin McConway and Professor Sir David Spiegelhalter. Sound human, steer clear of jargon, and be prepared. Significance (Oxford, England), 18(2):32–34, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Franklin:2021:CMC

[1475] Christine Franklin. As Covid makes clear, statistics education is a must. Significance (Oxford, England), 18(2):35, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Silge:2021:IAB

[1476] Julia Silge. “I’ve always been most interested in highly applied work and how people get stuff done in the real world”. Significance (Oxford, England), 18(2):36–37, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harron:2021:WBI

[1477] Katie Harron. What’s the big idea? Data linkage. Significance (Oxford, England), 18(2):38–39, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nutt:2021:RBC

[1478] David Nutt. A risky business? Comparing the harms of alcohol and other recreational drugs. Significance (Oxford, England), 18(2):40–42, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

King:2021:RNN

[1479] Thomas King. Review: News, Numbers, and Public Opinion in a Data-Driven World. Edited by An Nguyen; Bloomsbury; xvi + 288pp; December 2017; £28.99 (paperback); ISBN 978-1-5013-5400-7. Significance (Oxford, England), 18(2):43, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Allin:2021:BBB

[1480] Paul Allin and David J. Hand. Building back better needs better use of statistics. Significance (Oxford, England), 18(2):44–45, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Lb

[1481] Anonymous. Letters. Significance (Oxford, England), 18(2):46–47, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Pb

[1482] Anonymous. Puzzle. Significance (Oxford, England), 18(2):47, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:TMH

[1483] Anonymous. Timothy Martyn Hill: Political betting. Significance (Oxford, England), 18(2):48, April 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Cb

[1484] Anonymous. Contents. Significance (Oxford, England), 18(3):1, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Nc

[1485] Anonymous. News. Significance (Oxford, England), 18(3):2–3, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:NLA

[1486] Anonymous. Nan Laird awarded International Prize in Statistics. Significance (Oxford, England), 18(3):4, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Paulden:2021:SFS

[1487] Dr Tim Paulden. Soccer fans in shiny sticker shock. Significance (Oxford, England), 18(3):5, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Schwab:2021:SPS

[1488] Simon Schwab and Leonhard Held. Statistical programming: Small mistakes, big impacts. Significance (Oxford, England), 18(3):6–7, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kenett:2021:DFI

[1489] Professor Ron Kenett and Dr Shirley Coleman. Data and the Fourth Industrial Revolution. Significance (Oxford, England), 18(3):8–9, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kasza:2021:SMW

[1490] Jessica Kasza. SSA members, welcome! Significance (Oxford, England), 18(3):10, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2021:Ec

[1491] Brian Tarran. Editorial. Significance (Oxford, England), 18(3):11, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mitra:2021:SS

[1492] Dr Sinjini Mitra. A scanner smartly. Significance (Oxford, England), 18 (3):12–17, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2021:PIS

[1493] Brian Tarran. Prepare for impact: Space debris and statistics. Significance (Oxford, England), 18(3):18–23, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kurz:2021:UDP

[1494] Christoph Kurz. Understanding differential privacy. Significance (Oxford, England), 18(3):24–27, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gastwirth:2021:CFW

[1495] Joseph L. Gastwirth and Qing Shi. The changing face of wealth inequality in the USA. Significance (Oxford, England), 18(3):28–31, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Haughton:2021:BO

[1496] Jonathan Haughton. Beware of outliers! Significance (Oxford, England), 18(3):32–33, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Caulcutt:2021:HCW

[1497] Roland Caulcutt. How can we help decision-makers? Significance (Oxford, England), 18(3):34–35, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Golbeck:2021:RLT

[1498] Amanda L. Golbeck. The road less travelled… How Elizabeth L. Scott and F. N. David made an impact. Significance (Oxford, England), 18 (3):36–38, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Borja:2021:FOO

[1499] Mario Cortina Borja and Brian Tarran. A farewell to one of our founders: John Haigh (1941–2021). Significance (Oxford, England), 18(3):39, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:CS

[1500] Anonymous. Career story. Significance (Oxford, England), 18(3):40–41, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bird:2021:CTS

[1501] Sheila M. Bird. Covid tests in secondary schools: a statistical cause célèbre. Significance (Oxford, England), 18(3):42–45, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Lc

[1502] Anonymous. Letters. Significance (Oxford, England), 18(3):46–47, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Pd

[1503] Anonymous. Puzzle. Significance (Oxford, England), 18(3):47, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:JSK

[1504] Anonymous. Julian Stander, Katerina Tzioli and Mario Cortina Borja: 200 years of Greek independence… and PMs. Significance (Oxford, England), 18(3):48, June 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Cc

[1505] Anonymous. Contents. Significance (Oxford, England), 18(4):1, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Nd

[1506] Anonymous. News. Significance (Oxford, England), 18(4):2–3, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jackson:2021:NFN

[1507] Alan Jackson. No fans, no home advantage? Significance (Oxford, England), 18(4):4–5, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Alba:2021:BPB

[1508] Sandra Alba. Bats, parachutes and bridges: How can epidemiologists improve global health research practice? Significance (Oxford, England), 18(4):6–7, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cernat:2021:MNQ

[1509] Alexandru Cernat and Joseph W. Sakshaug. “My next question is, could I have some blood?”. Significance (Oxford, England), 18(4):8–9, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fletcher:2021:PPH

[1510] Michael Fletcher. Probability puzzle: Higher or lower? Significance (Oxford, England), 18(4):10, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2021:JD

[1511] Brian Tarran. The joy of data. Significance (Oxford, England), 18(4): 11, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bonnett:2021:DRB

[1512] Laura Bonnett. Driving is a risky business. Significance (Oxford, England), 18(4):12–15, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harkness:2021:HDEb

[1513] Timandra Harkness. The history of the data economy: Part II: Analytics arrives. Significance (Oxford, England), 18(4):16–19, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Stigler:2021:EPJ

[1514] Stephen Stigler. Early probability and John Law. Significance (Oxford, England), 18(4):20–23, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Leinwand:2021:WEP

[1515] Benjamin Leinwand, Puyao Ge, Vidyadhar Kulkarni, and Richard Smith. Winning an election, not a popularity contest. Significance (Oxford, England), 18(4):24–29, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lum:2021:HMS

[1516] Kristian Lum, Naim Kabir, and Joe Bak-Coleman. How to mislead with statistics. Significance (Oxford, England), 18(4):30–32, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2021:QAK

[1517] Brian Tarran. Questions answered: Katie Buchhorn. Significance (Oxford, England), 18(4):33, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Laird:2021:APE

[1518] Nan Laird and Edward C. Hirschland. From the Apollo programme to the EM algorithm and beyond. Significance (Oxford, England), 18(4): 34–39, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Horton:2021:WBI

[1519] Nicholas J. Horton. What’s the big idea? The EM algorithm. Significance (Oxford, England), 18(4):40–41, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Seibold:2021:SRY

[1520] Heidi Seibold, Alethea Charlton, Anne-Laure Boulesteix, and Sabine Hoffmann. Statisticians, roll up your sleeves! There’s a crisis to be solved. Significance (Oxford, England), 18(4):42–44, August 2021. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Shalabh:2021:B

[1521] Shalabh. Books. Significance (Oxford, England), 18(4):45, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Ld

[1522] Anonymous. Letters. Significance (Oxford, England), 18(4):46–47, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Pe

[1523] Anonymous. Puzzle. Significance (Oxford, England), 18(4):47, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:ISW

[1524] Anonymous. Ian Stewart: a world without mathematics. Significance (Oxford, England), 18(4):48, August 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Cd

[1525] Anonymous. Contents. Significance (Oxford, England), 18(5):1, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Ne

[1526] Anonymous. News. Significance (Oxford, England), 18(5):2–3, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:AFS

[1527] Anonymous. Alliance formed to set standards for data science. Significance (Oxford, England), 18(5):4–5, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:PC

[1528] Anonymous. Paper cuts. Significance (Oxford, England), 18(5):6–7, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2021:WSW

[1529] Brian Tarran. “What stood out was how much pressure the system had been under”. Significance (Oxford, England), 18(5):8–9, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fletcher:2021:PPCa

[1530] Michael Fletcher. Probability puzzle: a certain bet. Significance (Oxford, England), 18(5):10, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2021:WRN

[1531] Brian Tarran. A welcome return to normal. Significance (Oxford, England), 18(5):11, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jackson:2021:PWT

[1532] Conner Jackson. Pietro the weather tortoise and the pursuit of soggy bun prevention. Significance (Oxford, England), 18(5):12–15, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harkness:2021:HDEc

[1533] Timandra Harkness. The history of the data economy: Part III: the new kings and queens of data. Significance (Oxford, England), 18(5): 16–19, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hirschauer:2021:IUN

[1534] Norbert Hirschauer, Sven Grüner, Oliver Mußhoff, Claudia Becker, and Antje Jantsch. Inference using non-random samples? Stop right there! Significance (Oxford, England), 18(5):20–24, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Verma:2021:HCP

[1535] Rakesh M. Verma and David J. Marchette. How to catch a phish with statistics. Significance (Oxford, England), 18(5):26–29, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gandy:2021:WBM

[1536] Rob Gandy. What’s both massive and slender, and thin but convincing? Significance (Oxford, England), 18(5):30–31, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Foster:2021:RNT

[1537] Craig A. Foster. Race and the NFL: Teaching the chi-square goodness of fit test. Significance (Oxford, England), 18(5):32–33, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Richardson:2021:CLL

[1538] John T. E. Richardson. A closer look at the lady tasting tea. Significance (Oxford, England), 18(5):34–37, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Raviv:2021:WBI

[1539] Eran Raviv. What’s the big idea? Deep learning algorithms. Significance (Oxford, England), 18(5):38–39, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Basner:2021:WWB

[1540] Mathias Basner. Who was the best Friend? Significance (Oxford, England), 18(5):40–43, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:B

[1541] Anonymous. Books. Significance (Oxford, England), 18(5):45, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Le

[1542] Anonymous. Letters. Significance (Oxford, England), 18(5):46–47, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Pf

[1543] Anonymous. Puzzle. Significance (Oxford, England), 18(5):47, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:LVW

[1544] Anonymous. Leighton Vaughan Williams: The Doomsday Argument. Significance (Oxford, England), 18(5):48, October 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Tb

[1545] Anonymous. TOC. Significance (Oxford, England), 18(6):1, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Nf

[1546] Anonymous. News. Significance (Oxford, England), 18(6):2–3, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mealli:2021:ACQ

[1547] Fabrizia Mealli. Answering causal questions: Angrist, Imbens and the Nobel prize. Significance (Oxford, England), 18(6):4–5, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Pearce:2021:WRB

[1548] Michael Pearce and Adrian E. Raftery. Will this be a record-breaking century for human longevity? Significance (Oxford, England), 18(6): 6–7, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fletcher:2021:PPCb

[1549] Michael Fletcher. Probability puzzle: Classroom bingo. Significance (Oxford, England), 18(6):7, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Small:2021:FNG

[1550] Hugh Small. Florence Nightingale’s “greatest feat”. Significance (Oxford, England), 18(6):8–9, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2021:Ed

[1551] Brian Tarran. Editorial. Significance (Oxford, England), 18(6):11, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harkness:2021:HDEd

[1552] Timandra Harkness. The history of the data economy: Part IV: the future. Significance (Oxford, England), 18(6):12–15, December 2021. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

vanZwet:2021:AEE

[1553] Erik van Zwet, Simon Schwab, and Sander Greenland. Addressing exaggeration of effects from single RCTs. Significance (Oxford, England), 18(6):16–21, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bailey:2021:HBO

[1554] David H. Bailey and Marcos López de Prado. How “backtest overfitting” in finance leads to false discoveries. Significance (Oxford, England), 18 (6):22–25, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Unwin:2021:IDS

[1555] Antony Unwin and Kim Kleinman. The iris data set: In search of the source of virginica. Significance (Oxford, England), 18(6):26–29, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Brown:2021:WRK

[1556] Tracey Brown, Christine Franklin, and Leonor Sierra. What is risk knowhow? Significance (Oxford, England), 18(6):30–31, December 2021. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:RKHa

[1557] Anonymous. The risk know-how framework. Significance (Oxford, England), 18(6):32–33, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:RKHb

[1558] Anonymous. Risk know-how: the expert view. Significance (Oxford, England), 18(6):34–35, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Smykla:2021:TLC

[1559] Margaret L. Smykla. The tragedy — and legacy — of Chen Wen-chen. Significance (Oxford, England), 18(6):36–37, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Chao:2021:RCW

[1560] Anne Chao and Liza Levina. Remembering Chen Wen-chen. Significance (Oxford, England), 18(6):38, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Chao:2021:CWC

[1561] Anne Chao. Chen Wen-chen’s academic contributions. Significance (Oxford, England), 18(6):39, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Duan:2021:TDC

[1562] Naihua Duan and Chii-Ruey Hwang. Taiwan, democracy and Chen Wenchen. Significance (Oxford, England), 18(6):40–41, December 2021. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Boyd:2021:SSS

[1563] Professor Sir Ian L. Boyd. Scepticism, science and statistics. Significance (Oxford, England), 18(6):42–45, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Lf

[1564] Anonymous. Letters. Significance (Oxford, England), 18(6):46–47, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2021:Pg

[1565] Anonymous. Puzzle. Significance (Oxford, England), 18(6):47, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2021:QAM

[1566] Brian Tarran. Questions answered: Marcus du Sautoy. Significance (Oxford, England), 18(6):48, December 2021. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Ta

[1567] Anonymous. TOC. Significance (Oxford, England), 19(1):1, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Na

[1568] Anonymous. News. Significance (Oxford, England), 19(1):2–3, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bird:2022:MRD

[1569] Sheila M. Bird. Medical research, data sharing, and properly informed consent. Significance (Oxford, England), 19(1):4–5, February 2022. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ferguson:2022:MSE

[1570] Wallace Ferguson. Microscopium Statisticum and the etymology of statistics. Significance (Oxford, England), 19(1):6–7, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fletcher:2022:PPH

[1571] Michael Fletcher. Probability puzzle: How many balls in the bag? Significance (Oxford, England), 19(1):8, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2022:Ea

[1572] Brian Tarran. Editorial. Significance (Oxford, England), 19(1):9, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rougier:2022:RHG

[1573] Jonathan Rougier. The “reliable history” of geohazards. Significance (Oxford, England), 19(1):10–13, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Bowen:2022:ADP

[1574] Claire McKay Bowen. The art of data privacy. Significance (Oxford, England), 19(1):14–19, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Owora:2022:RCD

[1575] Arthur H. Owora, John Dawson, Gary Gadbury, Luis M. Mestre, Greg Pavela, Tapan Mehta, Colby J. Vorland, Pengcheng Xun, and David B. Allison. Randomisation can do many things — but it cannot “fail”. Significance (Oxford, England), 19(1):20–23, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fox:2022:DMC

[1576] James Alan Fox, Nathan E. Sanders, Emma E. Fridel, Grant Duwe, and Michael Rocque. Does media coverage of mass public shootings create a contagion effect? Significance (Oxford, England), 19(1):24–28, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Alba:2022:BLJ

[1577] Sandra Alba. Blurring the lines: Journalism and epidemiology in the time of corona. Significance (Oxford, England), 19(1):30–33, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gabriel:2022:MAH

[1578] Florence Gabriel. Maths anxiety — and how to overcome it. Significance (Oxford, England), 19(1):34–35, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2022:CRU

[1579] Brian Tarran. “Covid has reminded us that independent science is really important. And statistics is part of that” A conversation with Steve MacFeely. Significance (Oxford, England), 19(1):36–39, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Thornton:2022:TSB

[1580] Suzanne Thornton, Dooti Roy, Stephen Parry, Donna LaLonde, Wendy Martinez, Renee Ellis, and David Corliss. Towards statistical best practices for gender and sex data. Significance (Oxford, England), 19(1): 40–45, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:La

[1581] Anonymous. Letters. Significance (Oxford, England), 19(1):46–47, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Pa

[1582] Anonymous. Puzzle. Significance (Oxford, England), 19(1):47, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Buttrey:2022:QAS

[1583] Sam Buttrey. Questions answered: Sam Buttrey. Significance (Oxford, England), 19(1):48, February 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Tb

[1584] Anonymous. TOC. Significance (Oxford, England), 19(2):1, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Nb

[1585] Anonymous. News. Significance (Oxford, England), 19(2):2–3, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Heyard:2022:WSD

[1586] Rachel Heyard and Leonhard Held. When should data and code be made available? Significance (Oxford, England), 19(2):4–5, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fratczak:2022:CDV

[1587] Monika Fratczak. Can data visualisations mobilise people to act? Significance (Oxford, England), 19(2):6–7, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fletcher:2022:ONP

[1588] Michael Fletcher. Oh no! Puncture! Significance (Oxford, England), 19(2):8, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2022:Eb

[1589] Brian Tarran. Editorial. Significance (Oxford, England), 19(2):9, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kobak:2022:USA

[1590] Dmitry Kobak. Underdispersion: a statistical anomaly in reported Covid data. Significance (Oxford, England), 19(2):10–13, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Prestige:2022:CLU

[1591] Emily Prestige, Julian Stander, and Yinghui Wei. Covid lockdowns in the UK: Estimating their effects on transmission. Significance (Oxford, England), 19(2):14–18, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:CIM

[1592] Anonymous. Covid-19 in Italy: Modelling, communications, and collaborations. Significance (Oxford, England), 19(2):19–21, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Mitra:2022:PSR

[1593] Sinjini Mitra and Denise Stanley. Pandemic studies: Remote learning during Covid. Significance (Oxford, England), 19(2):22–24, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Harkness:2022:CC

[1594] Timandra Harkness. The Covid communicators. Significance (Oxford, England), 19(2):26–29, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Firth:2022:RSD

[1595] David Firth, Nancy Reid, Deborah G. Mayo, and Heather Battey. Remembering Sir David Cox, 1924–2022. Significance (Oxford, England), 19 (2):30–37, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tibshirani:2022:WCP

[1596] Robert Tibshirani. What is Cox’s proportional hazards model? Significance (Oxford, England), 19(2):38–39, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Borja:2022:SDC

[1597] Mario Cortina Borja, Julian Stander, and Giovanni Sebastiani. On Sir David Cox’s publications. Significance (Oxford, England), 19(2):39–40, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:R

[1598] Anonymous. References. Significance (Oxford, England), 19(2):41, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2022:WRA

[1599] Brian Tarran. “I want to be ready to answer questions we don’t yet know we need to know”: an interview with Alison Pritchard. Significance (Oxford, England), 19(2):42–44, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

King:2022:BR

[1600] Thomas King. Book Reviews. Significance (Oxford, England), 19(2): 45, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Lb

[1601] Anonymous. Letters. Significance (Oxford, England), 19(2):46–47, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Pb

[1602] Anonymous. Puzzle. Significance (Oxford, England), 19(2):47, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Leach:2022:OMT

[1603] Michael Leach. One more thing. … Significance (Oxford, England), 19 (2):48, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rubin:2022:TTS

[1604] Mark Rubin. That’s not a two-sided test! It’s two one-sided tests! Significance (Oxford, England), 19(2):50–53, April 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Tc

[1605] Anonymous. TOC. Significance (Oxford, England), 19(3):1, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Nc

[1606] Anonymous. News. Significance (Oxford, England), 19(3):2–3, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:GTC

[1607] Anonymous. Georgiou takes case to European Court of Human Rights. Significance (Oxford, England), 19(3):4, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2022:CDO

[1608] Brian Tarran. In conversation with a digital optimist. Significance (Oxford, England), 19(3):5, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Kwasny:2022:WW

[1609] Mary J. Kwasny. War of the Wordlers. Significance (Oxford, England), 19(3):6–7, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fletcher:2022:Pa

[1610] Michael Fletcher. Puzzle. Significance (Oxford, England), 19(3):8–9, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2022:E

[1611] Brian Tarran. Editorial. Significance (Oxford, England), 19(3):9, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Iacus:2022:DAJ

[1612] Stefano M. Iacus and Giuseppe Porro. Don’t ask, just listen … Using social networks to measure subjective well-being. Significance (Oxford, England), 19(3):10–15, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lamb:2022:WS

[1613] Dr Karen Lamb, Jessica Kasza, Dr Sophie Calabretto, Dr Rushani Wijesuriya, and Dr Linda McIver. Women of Significance? Significance (Oxford, England), 19(3):16–20, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Robinson:2022:CEE

[1614] Daniel Robinson and Howard Wainer. College entrance exams: a modern “shibboleth”? Significance (Oxford, England), 19(3):22–27, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Chen:2022:DTP

[1615] Sina Chen and John Körtner. Did “Trumpists” perform worse in US Senate election polls? Significance (Oxford, England), 19(3):28–31, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Raper:2022:CC

[1616] Simon Raper. A confusion of curves. Significance (Oxford, England), 19 (3):32–35, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Golbeck:2022:BTS

[1617] Amanda L. Golbeck. Big telescopes: the story behind a story of discrimination in science. Significance (Oxford, England), 19(3):36–39, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Tarran:2022:MIV

[1618] Brian Tarran. Making the invisible visible. Significance (Oxford, England), 19(3):40–41, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Hirschauer:2022:UQP

[1619] Norbert Hirschauer. Unanswered questions in the p-value debate. Significance (Oxford, England), 19(3):42–44, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:BR

[1620] Anonymous. Book reviews. Significance (Oxford, England), 19(3):44–45, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Lc

[1621] Anonymous. Letters. Significance (Oxford, England), 19(3):46–47, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Ca

[1622] Anonymous. Crossword. Significance (Oxford, England), 19(3):47, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:QAA

[1623] Anonymous. Questions answered: Anna Britten. Significance (Oxford, England), 19(3):48, June 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Td

[1624] Anonymous. TOC. Significance (Oxford, England), 19(4):1, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Nd

[1625] Anonymous. News. Significance (Oxford, England), 19(4):2–3, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Barnett:2022:ULT

[1626] Adrian Barnett and Elizabeth Gadd. University league tables have no legs to stand on. Significance (Oxford, England), 19(4):4–7, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gebregziabher:2022:SCW

[1627] Mulugeta Gebregziabher and James J. Cochran. Statistics in a civil war: the case of the Tigray health system. Significance (Oxford, England), 19 (4):8–9, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cernat:2022:DSC

[1628] Alexandru Cernat and Florian Keusch. Do surveys change behaviours? Significance (Oxford, England), 19(4):10–11, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Britten:2022:WSO

[1629] Anna Britten. Women of Significance: Our pledge. Significance (Oxford, England), 19(4):12–13, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fletcher:2022:Pb

[1630] Michael Fletcher. Puzzle. Significance (Oxford, England), 19(4):14–15, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Britten:2022:Ea

[1631] Anna Britten. Editorial. Significance (Oxford, England), 19(4):15, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Walker:2022:MMA

[1632] Prof. James T. Walker, Dr John S. Lee, and Stephen Broadberry. Measuring the Middle Ages. Significance (Oxford, England), 19(4):16–23, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nunez:2022:CC

[1633] Isaac Núñez. Canine confounders. Significance (Oxford, England), 19(4): 24–27, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anderberg:2022:RSD

[1634] Dan Anderberg, Helmut Rainer, and Fabian Siuda. The real scale of domestic violence during Covid-19. Significance (Oxford, England), 19 (4):28–31, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Smith:2022:FMF

[1635] Gary Smith. Full moons and forking paths. Significance (Oxford, England), 19(4):32–35, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Britten:2022:SPA

[1636] Anna Britten. “Statisticians are particularly attuned to the beauty of uncertainty”. Significance (Oxford, England), 19(4):36–40, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Rothman:2022:MPP

[1637] James Rothman. My prostate predictions. Significance (Oxford, England), 19(4):42–45, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Ld

[1638] Anonymous. Letters. Significance (Oxford, England), 19(4):46–47, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Cb

[1639] Anonymous. Crossword. Significance (Oxford, England), 19(4):47, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:QARa

[1640] Anonymous. Questions answered: Raghib Ali. Significance (Oxford, England), 19(4):48, August 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Te

[1641] Anonymous. TOC. Significance (Oxford, England), 19(5):1, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Ne

[1642] Anonymous. News. Significance (Oxford, England), 19(5):2–3, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fletcher:2022:Pc

[1643] Michael Fletcher. Puzzle. Significance (Oxford, England), 19(5):4–5, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lorenzo-Arribas:2022:E

[1644] Altea Lorenzo-Arribas and Sandra Alba. Editorial. Significance (Oxford, England), 19(5):5, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cooke:2022:FSB

[1645] Lucy Cooke. The female of the species: Baboon mothers and a courageous career move. Significance (Oxford, England), 19(5):6–10, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dobbs:2022:PWW

[1646] Joy Dobbs. A place where women thrived: UK Social Survey Official Statistics. Significance (Oxford, England), 19(5):12–15, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

PhD:2022:BNV

[1647] Fulvia Mecatti PhD, Paola Vicard PhD, Flaminia Musella PhD, and Lorenzo Giammei. Bayesian networks versus gender bias. Significance (Oxford, England), 19(5):16–20, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Columbus:2022:GED

[1648] Alyssa Columbus. Gender equity in digital health: AI as a double-edged sword. Significance (Oxford, England), 19(5):21–24, October 2022. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Griffith:2022:TAM

[1649] Emily Griffith and Julia Sharp. “They are not alone”: Mentoring women graduate students. Significance (Oxford, England), 19(5):25–27, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lorenzo-Arribas:2022:BH

[1650] Altea Lorenzo-Arribas and Sandra Alba. Bot out of hell. Significance (Oxford, England), 19(5):28–29, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lorenzo-Arribas:2022:WSD

[1651] Altea Lorenzo-Arribas. “We are still not done; we need to keep this organisation alive”. Significance (Oxford, England), 19(5):30–34, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Taylor:2022:SMM

[1652] Ursula Winifred Taylor and John Aldrich. Statistician, missionary, mother. Significance (Oxford, England), 19(5):35–37, October 2022. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Gandy:2022:VCB

[1653] Rob Gandy, Altea Lorenzo-Arribas, and Richard Startup. Violet Cane: a breaker of codes and moulds. Significance (Oxford, England), 19(5): 38–41, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

PhD:2022:BNE

[1654] Lee Kennedy-Shaffer PhD. Baseball’s natural experiment. Significance (Oxford, England), 19(5):42–45, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Le

[1655] Anonymous. Letters. Significance (Oxford, England), 19(5):46–47, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Cc

[1656] Anonymous. Crossword. Significance (Oxford, England), 19(5):47, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:QAM

[1657] Anonymous. Questions answered: Michelle McGhee. Significance (Oxford, England), 19(5):48, October 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Tf

[1658] Anonymous. TOC. Significance (Oxford, England), 19(6):1, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Nf

[1659] Anonymous. News. Significance (Oxford, England), 19(6):2–3, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Cockle:2022:RWH

[1660] Dr Sam Cockle. Revolutionising the way health is measured. Significance (Oxford, England), 19(6):4–7, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Fletcher:2022:WHO

[1661] Michael Fletcher. What’s happened to our chocolate bar? Significance (Oxford, England), 19(6):8–9, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Britten:2022:Eb

[1662] Anna Britten. Editorial. Significance (Oxford, England), 19(6):9, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Jackson:2022:PSE

[1663] James Jackson. Penalty shootouts: Are estimates on the spot? Significance (Oxford, England), 19(6):10–13, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Dowsett:2022:SRC

[1664] Matthew Dowsett. Seeing red: Challenging football’s “red card myth”. Significance (Oxford, England), 19(6):14–17, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Price:2022:BGU

[1665] Megan Price. Beautiful game, ugly truth? Significance (Oxford, England), 19(6):18–21, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Sanders:2022:CML

[1666] Michael Sanders and Vicky Clayton. Can machine learning save children at risk? Significance (Oxford, England), 19(6):22–27, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Lenzi:2022:HSK

[1667] Amanda Lenzi and Mihai Anitescu. How statistics keeps the lights on. Significance (Oxford, England), 19(6):28–30, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

PhD:2022:IMD

[1668] Scott Nestler PhD, Cap, Pstat, and Tom Keelin PhD. Introducing the metalog distributions. Significance (Oxford, England), 19(6):31–33, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Ryan:2022:S

[1669] Louise Ryan, John Henstridge, and Jessica Kasza. The SSA is 60! Significance (Oxford, England), 19(6):34–40, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

PhD:2022:WDA

[1670] Vincent Charles PhD, Ali Emrouznejad PhD, Tatiana Gherman PhD, and James Cochran. Why data analytics is an art. Significance (Oxford, England), 19(6):42–45, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Lf

[1671] Anonymous. Letters. Significance (Oxford, England), 19(6):46–47, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:Cd

[1672] Anonymous. Crossword. Significance (Oxford, England), 19(6):47, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2022:QARb

[1673] Anonymous. Questions answered: Rob Mastrodomenico. Significance (Oxford, England), 19(6):48, December 2022. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Anonymous:2023:Na

[1674] Anonymous. News. Significance (Oxford, England), 20(1):1–3, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

URL https://academic.oup.com/jrssig/article/20/1/1/7034187.

Rodriguez:2023:CCC

[1675] Jordi Prats Rodríguez. Chess cheats in check? Significance (Oxford, England), 20(1):4–5, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). URL https://academic.oup.com/

jrssig/article/20/1/4/7034180.

Sullivan:2023:BAS

[1676] Alice Sullivan. A big ask: Sex and data collection. Significance (Oxford, England), 20(1):6–9, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). URL https://academic.oup.com/

jrssig/article/20/1/6/7034186.

Tarran:2023:AG

[1677] Brian Tarran. Andreas Georgiou. Significance (Oxford, England), 20 (1):10–11, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-

9713 (electronic). URL https://academic.oup.com/jrssig/article/
20/1/10/7034176.

Fletcher:2023:Pa

[1678] Michael Fletcher. Puzzle. Significance (Oxford, England), 20(1):12, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (elec-

tronic). URL https://academic.oup.com/jrssig/article/20/1/12/
7034178.

Britten:2023:Ea

[1679] Anna Britten. Editorial. Significance (Oxford, England), 20(1):13, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (elec-

tronic). URL https://academic.oup.com/jrssig/article/20/1/13/
7034185.

Cressie:2023:ECB

[1680] Noel Cressie, Andrew Zammit-Mangion, Josh Jacobson, and Michael Bertolacci. Earth’s CO_{2} battle: a view from space. Significance (Oxford, England), 20(1):14–19, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). URL https://academic.oup.com/

jrssig/article/20/1/14/7034191.

SandraAlbaMsc:2023:MTS

[1681] Phd Sandra Alba, Msc and Drph Amera Khan. Music, tuberculosis and stats: Lessons from a Ugandan pop star. Significance (Oxford, England), 20(1):20–22, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-

9713 (electronic). URL https://academic.oup.com/jrssig/article/
20/1/20/7034182.

Lenzi:2023:HCS

[1682] Amanda Lenzi and Mihai Anitescu. How can statistics help to prevent blackouts? Significance (Oxford, England), 20(1):24–27, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). URL

https://academic.oup.com/jrssig/article/20/1/24/7034188.

Andrade:2023:DIP

[1683] Chris Andrade, Jonathan Auerbach, Icaro Bacelar, Hane Lee, and Angela Tan… Does it pay to park in front of a fire hydrant? Significance (Oxford, England), 20(1):28–30, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). URL https://academic.oup.

com/jrssig/article/20/1/28/7034189.

Hanley:2023:WLN

[1684] James Hanley and Michael Cronin. Which lottery numbers do players play? Significance (Oxford, England), 20(1):31–33, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). URL

https://academic.oup.com/jrssig/article/20/1/31/7034183.

Otto:2023:WN

[1685] Philipp Otto. What’s in a name? Significance (Oxford, England), 20 (1):34–37, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-

9713 (electronic). URL https://academic.oup.com/jrssig/article/
20/1/34/7034190.

Green:2023:WYD

[1686] Nathan Green. Why your data viz needs alt text. Significance (Oxford, England), 20(1):38–39, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). URL https://academic.oup.com/

jrssig/article/20/1/38/7034177.

Bennett:2023:IIG

[1687] Hayley Bennett. Interview: Ian Gordon. Significance (Oxford, England), 20(1):40–41, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-

9713 (electronic). URL https://academic.oup.com/jrssig/article/
20/1/40/7034179.

Bennett:2023:ICM

[1688] Hayley Bennett. Interview: Claire McKay Bowen. Significance (Oxford, England), 20(1):42–43, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). URL https://academic.oup.com/

jrssig/article/20/1/42/7034181.

Anonymous:2023:WCL

[1689] Anonymous. 2023 writing competition launch. Significance (Oxford, England), 20(1):44–45, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). URL https://academic.oup.com/

jrssig/article/20/1/44/7034175.

Anonymous:2023:La

[1690] Anonymous. Letters. Significance (Oxford, England), 20(1):46, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). URL

https://academic.oup.com/jrssig/article/20/1/46/7034174.

Anonymous:2023:Ca

[1691] Anonymous. Crossword. Significance (Oxford, England), 20(1):47, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (elec-

tronic). URL https://academic.oup.com/jrssig/article/20/1/47/
7034184.

Anonymous:2023:QAB

[1692] Anonymous. Questions answered: Bonnie Ghosh-Dastidar. Significance (Oxford, England), 20(1):48, February 2023. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic). URL https://academic.oup.com/

jrssig/article/20/1/48/7034192.

Anonymous:2023:Nb

[1693] Anonymous. News. Significance (Oxford, England), 20(2):2–3, April

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/2/2/7095718.

Tarran:2023:RTF

[1694] Brian Tarran. Real talk? The future of ChatGPT. Significance (Oxford, England), 20(2):4–5, April 2023. CODEN ???? ISSN ???? URL https:

//academic.oup.com/jrssig/article/20/2/4/7095709.

Gustar:2023:EVG

[1695] Andrew Gustar. Eurovision voting: a game of two halves. Significance (Oxford, England), 20(2):6–10, April 2023. CODEN ???? ISSN ????

URL https://academic.oup.com/jrssig/article/20/2/6/7095712.

Clarkson:2023:FPS

[1696] Dan Clarkson. Food poverty: a scale unimagined. Significance (Oxford, England), 20(2):11–13, April 2023. CODEN ???? ISSN ???? URL

https://academic.oup.com/jrssig/article/20/2/11/7095721.

Anonymous:2023:P

[1697] Anonymous. Puzzle. Significance (Oxford, England), 20(2):14, April

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/2/14/7095717.

Britten:2023:Eb

[1698] Anna Britten. Editorial. Significance (Oxford, England), 20(2):15, April

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/2/15/7095704.

Salahub:2023:RBJ

[1699] Chris Salahub. Racial bias in jury selection. Significance (Oxford, England), 20(2):16–20, April 2023. CODEN ???? ISSN ???? URL https:/

/academic.oup.com/jrssig/article/20/2/16/7095719.

Schmid:2023:NCW

[1700] Monika S. Schmid. No comprendo: Will we ever understand how foreign language skills are lost? Significance (Oxford, England), 20(2):22–25, April 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/2/22/7095707.

Stapleton:2023:BBC

[1701] Matt Stapleton. The beauty of binomial coverage curves. Significance (Oxford, England), 20(2):26–27, April 2023. CODEN ???? ISSN

???? URL https://academic.oup.com/jrssig/article/20/2/26/
7095714.

Doganaksoy:2023:PRH

[1702] Necip Doganaksoy, William Q Meeker, and Gerald J Hahn. Product reliability: How statistics fits in. Significance (Oxford, England), 20(2): 28–32, April 2023. CODEN ???? ISSN ???? URL https://academic.

oup.com/jrssig/article/20/2/28/7095711.

Lo:2023:CSU

[1703] Sherman Lo. Calling the shots: using a statistical model to rank archers. Significance (Oxford, England), 20(2):33–35, April 2023. CODEN ????

ISSN ???? URL https://academic.oup.com/jrssig/article/20/2/
33/7095716.

Cowan:2023:PPP

[1704] Mary Cowan, Shayda Kashef, and Ed Humpherson. Putting the “public” into public good. Significance (Oxford, England), 20(2):36–39, April

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/2/36/7095706.

Anonymous:2023:NYZ

[1705] Anonymous. Nonhlanhla Yende-Zuma: “We are slowly and carefully changing our thinking”. Significance (Oxford, England), 20(2):40–42, April 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/2/40/7095708.

Rogers:2023:NGH

[1706] Nick Rogers. Numbers game: How I fell in love with statistics. Significance (Oxford, England), 20(2):43–44, April 2023. CODEN ????

ISSN ???? URL https://academic.oup.com/jrssig/article/20/2/
43/7095710.

Anonymous:2023:BRa

[1707] Anonymous. Book reviews. Significance (Oxford, England), 20(2):45, April 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/2/45/7095720.

Anonymous:2023:Lb

[1708] Anonymous. Letters. Significance (Oxford, England), 20(2):46, April

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/2/46/7095705.

Anonymous:2023:Cb

[1709] Anonymous. Crossword. Significance (Oxford, England), 20(2):47, April

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/2/47/7095713.

Wan:2023:QA

[1710] Alan Wan. Questions answered. Significance (Oxford, England), 20(2): 48, April 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/2/48/7095715.

Anonymous:2023:Nc

[1711] Anonymous. News. Significance (Oxford, England), 20(3):2–3, June 2023.

CODEN ???? ISSN ???? URL https://academic.oup.com/jrssig/
article/20/3/2/7190564.

Fletcher:2023:Pb

[1712] Michael Fletcher. Puzzle. Significance (Oxford, England), 20(3):4, June

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/3/4/7190560.

Britten:2023:Ec

[1713] Anna Britten. Editorial. Significance (Oxford, England), 20(3):5, June

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/3/5/7190562.

Anonymous:2023:I

[1714] Anonymous. 100th issue. Significance (Oxford, England), 20(3):6–12, June 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/3/6/7190575.

Matiussi:2023:DB

[1715] Adam Matiussi and Joseph Shaw. Data at the ballet. Significance (Oxford, England), 20(3):14–17, June 2023. CODEN ???? ISSN ???? URL

https://academic.oup.com/jrssig/article/20/3/14/7190573.

Rice:2023:PRS

[1716] Caity Rice, Peter Winter, Alan Rowson, and Jeroen Peters. Pooling resources: Statistics, swimming and why winners don’t always finish first. Significance (Oxford, England), 20(3):18–23, June 2023. CODEN

???? ISSN ???? URL https://academic.oup.com/jrssig/article/
20/3/18/7190572.

Shatz:2023:LSN

[1717] Itamar Shatz. The law of small numbers: When statistics and psychology go to war. Significance (Oxford, England), 20(3):24–25, June 2023.

CODEN ???? ISSN ???? URL https://academic.oup.com/jrssig/
article/20/3/24/7190574.

Ditlevsen:2023:WNL

[1718] Susanne Ditlevsen and Rolf Poulsen. When numbers lie: a good bad example. Significance (Oxford, England), 20(3):26–29, June 2023. CODEN

???? ISSN ???? URL https://academic.oup.com/jrssig/article/
20/3/26/7190570.

Marchant:2023:BLR

[1719] Paul Marchant and Paul Norman. In the best light? Road safety and public spending. Significance (Oxford, England), 20(3):30–31, June 2023.

CODEN ???? ISSN ???? URL https://academic.oup.com/jrssig/
article/20/3/30/7190566.

Christofides:2023:DD

[1720] Sarah Christofides. Dragons and data. Significance (Oxford, England), 20(3):31–33, June 2023. CODEN ???? ISSN ???? URL https://

academic.oup.com/jrssig/article/20/3/31/7190569.

Bennett:2023:TSJ

[1721] Hayley Bennett. “I took strength just by looking at the sky, stars, trees, flowers”: Interview with Anar Meshimbayeva. Significance (Oxford, England), 20(3):34–37, June 2023. CODEN ???? ISSN ???? URL https:/

/academic.oup.com/jrssig/article/20/3/34/7190559.

Anonymous:2023:WWT

[1722] Anonymous. What we talk about when we talk about data. Significance (Oxford, England), 20(3):38–41, June 2023. CODEN ???? ISSN

???? URL https://academic.oup.com/jrssig/article/20/3/38/
7190561.

Waynforth:2023:HNN

[1723] David Waynforth. How “normal” is the normal distribution? Significance (Oxford, England), 20(3):42–44, June 2023. CODEN ????

ISSN ???? URL https://academic.oup.com/jrssig/article/20/3/
42/7190563.

Anonymous:2023:BRb

[1724] Anonymous. Book reviews. Significance (Oxford, England), 20(3):45, June 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/3/45/7190565.

Anonymous:2023:Lc

[1725] Anonymous. Letters. Significance (Oxford, England), 20(3):46, June

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/3/46/7190568.

Anonymous:2023:Cc

[1726] Anonymous. Crossword. Significance (Oxford, England), 20(3):47, June

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/3/47/7190571.

Ryan:2023:QA

[1727] Mary Ryan. Questions answered. Significance (Oxford, England), 20(3): 48, June 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/3/48/7190567.

Britten:2023:Na

[1728] Anna Britten. News. Significance (Oxford, England), 20(4):2–3, August

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/4/2/7233197.

Prats:2023:HHH

[1729] Jordi Prats. Hot, hotter, hottest? Significance (Oxford, England), 20(4): 4–5, August 2023. CODEN ???? ISSN ???? URL https://academic.

oup.com/jrssig/article/20/4/4/7233195.

Fletcher:2023:Pc

[1730] Michael Fletcher. Puzzle. Significance (Oxford, England), 20(4):6, August 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/4/6/7233193.

Britten:2023:Ed

[1731] Anna Britten. Editorial. Significance (Oxford, England), 20(4):7, August

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/4/7/7233191.

Gao:2023:VDS

[1732] Fengnan Gao and Richard D. Gill. A very Dutch scandal. Significance (Oxford, England), 20(4):8–12, August 2023. CODEN ???? ISSN ????

URL https://academic.oup.com/jrssig/article/20/4/8/7233204.

Ritskes-Hoitinga:2023:TTL

[1733] Merel Ritskes-Hoitinga, Justine Watkins, and Pleun Jornick. Testing times: Let’s leave the animals out of it. Significance (Oxford, England), 20(4):13–17, August 2023. CODEN ???? ISSN ???? URL https://

academic.oup.com/jrssig/article/20/4/13/7233199.

Riva:2023:SLL

[1734] Chris Dalla Riva. Sweet little lies: Deceptive statistics in the album charts. Significance (Oxford, England), 20(4):18–21, August 2023. CO-

DEN ???? ISSN ???? URL https://academic.oup.com/jrssig/
article/20/4/18/7233200.

Danny:2023:WPP

[1735] Dorling Danny. World population projections: Just little bits of history repeating? Significance (Oxford, England), 20(4):22–27, August 2023.

CODEN ???? ISSN ???? URL https://academic.oup.com/jrssig/
article/20/4/22/7233203.

Stoudt:2023:CTM

[1736] Sara Stoudt. Can TV make you a better stats communicator? Significance (Oxford, England), 20(4):28–31, August 2023. CODEN ????

ISSN ???? URL https://academic.oup.com/jrssig/article/20/4/
28/7233202.

Anonymous:2023:PSQ

[1737] Anonymous. In praise of “stupid” questions. Significance (Oxford, England), 20(4):32–33, August 2023. CODEN ???? ISSN ???? URL https:

//academic.oup.com/jrssig/article/20/4/32/7233190.

Lorenzo-Arribas:2023:MFE

[1738] Altea Lorenzo-Arribas. Millicent Fawcett and Eleanor Rathbone: The Royal Statistical Society suffragists. Significance (Oxford, England), 20 (4):34–39, August 2023. CODEN ???? ISSN ???? URL https://

academic.oup.com/jrssig/article/20/4/34/7233201.

Gruninger:2023:SFS

[1739] Servan Grüninger. Stefano Franscini: The statistician who built a nation. Significance (Oxford, England), 20(4):40–43, August 2023. CODEN ????

ISSN ???? URL https://academic.oup.com/jrssig/article/20/4/
40/7233192.

Anonymous:2023:BRc

[1740] Anonymous. Book reviews. Significance (Oxford, England), 20(4):45, August 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/4/45/7233194.

Anonymous:2023:Ld

[1741] Anonymous. Letters. Significance (Oxford, England), 20(4):46–47, August 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/4/46/7233196.

Anonymous:2023:Cd

[1742] Anonymous. Crossword. Significance (Oxford, England), 20(4):47, August 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/4/47/7233198.

Nuijten:2023:QMN

[1743] Michèle Nuijten. Q&A Michèle Nuijten. Significance (Oxford, England), 20(4):48, August 2023. CODEN ???? ISSN ???? URL https://

academic.oup.com/jrssig/article/20/4/48/7233189.

Britten:2023:Nb

[1744] Anna Britten. News. Significance (Oxford, England), 20(5):2–3, October

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/5/2/7296075.

Fletcher:2023:Pd

[1745] Michael Fletcher. Puzzle. Significance (Oxford, England), 20(5):4, October 2023. CODEN ???? ISSN ???? URL https://academic.oup.com/

jrssig/article/20/5/4/7296070.

Britten:2023:Ee

[1746] Anna Britten. Editorial. Significance (Oxford, England), 20(5):5, October 2023. CODEN ???? ISSN ???? URL https://academic.oup.com/

jrssig/article/20/5/5/7296066.

Hill:2023:FEW

[1747] Timothy Martyn Hill. Fatal error: Who is winning in the war between Russia and Ukraine? Significance (Oxford, England), 20(5):6–11, October 2023. CODEN ???? ISSN ???? URL https://academic.oup.com/

jrssig/article/20/5/6/7296077.

Prats:2023:FP

[1748] Jordi Prats. Fit for purpose. Significance (Oxford, England), 20(5):12–19, October 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/5/12/7296078.

Bhaduri:2023:OCD

[1749] Moinak Bhaduri. Oxygen, crampons, data: Can statistics help us climb Everest? Significance (Oxford, England), 20(5):20–25, October 2023.

CODEN ???? ISSN ???? URL https://academic.oup.com/jrssig/
article/20/5/20/7296079.

Gascon:2023:DFD

[1750] Gregg M Gascon, Katherine K Wallman, Sridhar Ravula, Joseph Cappelleri, Julia Lee, Kristian Lum, Chaitra H Nagaraja, Sudeshna Paul, Dooti Roy, and Edward Wu. DSLA 2023: Facing the dangers of data illiteracy. Significance (Oxford, England), 20(5):26–29, October 2023.

CODEN ???? ISSN ???? URL https://academic.oup.com/jrssig/
article/20/5/26/7296072.

Brunton-Smith:2023:MCP

[1751] Ian Brunton-Smith, Alexandru Cernat, David Buil-Gil, and Jose Pina-Sànchez. Measuring crime in place: Distinguishing between area victimisation and area offences. Significance (Oxford, England), 20(5):30–33, October 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/5/30/7296073.

Mallinckrodt:2023:BYC

[1752] Craig Mallinckrodt. Building your career in statistics. Significance (Oxford, England), 20(5):34–37, October 2023. CODEN ???? ISSN

???? URL https://academic.oup.com/jrssig/article/20/5/34/
7296074.

Runge:2023:SWC

[1753] Johnny Runge. “This should be a wake-up call”: How to improve people’s faith in economic statistics. Significance (Oxford, England), 20(5):38–39, October 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/5/38/7296068.

Britten:2023:MPS

[1754] Anna Britten. “My professor said, ‘you can have this great career in mathematics if you go into statistics”’: Interview with Robert Santos. Significance (Oxford, England), 20(5):40–43, October 2023. CODEN ????

ISSN ???? URL https://academic.oup.com/jrssig/article/20/5/
40/7296065.

Anonymous:2023:BRd

[1755] Anonymous. Book review. Significance (Oxford, England), 20(5):45, October 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/5/45/7296067.

Anonymous:2023:Le

[1756] Anonymous. Letters. Significance (Oxford, England), 20(5):46, October

2023. CODEN ???? ISSN ???? URL https://academic.oup.com/
jrssig/article/20/5/46/7296069.

Buttrey:2023:C

[1757] Sam Buttrey. Crossword. Significance (Oxford, England), 20(5):47, October 2023. CODEN ???? ISSN ???? URL https://academic.oup.

com/jrssig/article/20/5/47/7296071.

Anonymous:2023:QGE

[1758] Anonymous. Q&A Gwen Eadie. Significance (Oxford, England), 20(5): 48, October 2023. CODEN ???? ISSN ???? URL https://academic.

oup.com/jrssig/article/20/5/48/7296076.

Dykes:2020:EDS

[1759] Brent Dykes. Effective Data Storytelling: How to Drive Change with Data, Narrative and Visuals. John Wiley, New York, NY, USA, 2020. ISBN 1-119-61571-2, 1-119-61572-0 (e-book). xiv + 322 pp. LCCN HF5718 .D95 2020.